### v1

# Multiple Invertible and Partial-Equivariant Function for Latent Vector Transformation to Enhance Disentanglement in VAEs

Abstract
Disentanglement learning is a core issue for understanding and re-using trained information in Variational AutoEncoder (VAE), and effective inductive bias has been reported as a key factor. However, the actual implementation of such bias is still vague. In this paper, we propose a novel method, called Multiple Invertible and partial-equivariant transformation (MIPE-transformation), to inject inductive bias by 1) guaranteeing the invertibility of latent-to-latent vector transformation while preserving a certain portion of equivariance of input-to-latent vector transformation, called Invertible and partial-equivariant transformation (IPE-transformation), 2) extending the form of prior and posterior in VAE frameworks to an unrestricted form through a learnable conversion to an approximated exponential family, called Exponential Family conversion (EF-conversion), and 3) integrating multiple units of IPE-transformation and EF-conversion, and their training. In experiments on 3D Cars, 3D Shapes, and dSprites datasets, MIPE-transformation improves the disentanglement performance of state-of-the-art VAEs.
1 Introduction
Disentanglement learning to learn more interpretable representations is broadly useful in artificial intelligence fields such as classification Singla et al. (2021), zero-shot learning Tenenbaum (2018), and domain adaptation Li et al. (2019); Zou et al. (2020). The disentangled representation is defined as a change in a single dimension, which corresponds to unique semantic information. Several works have been conducted based on this framework.
A major model for enhancing the disentanglement learning is Variational AutoEncoder (VAE) Kingma & Welling (2013). Based on VAE, unsupervised disentangled representation learning has been elaborated Higgins et al. (2017); Chen et al. (2018); Kim & Mnih (2018); Jeong & Song (2019); Li et al. (2020) through the factorizable variations and control of uncorrelatedness of each dimension of representations. Moreover, VAE models that handle the shape of prior as a Gaussian mixture Dilokthanakul et al. (2016b) or von Mises-Fisher Davidson et al. (2018) were also developed, but the disentanglement is still incomplete. As a critical point, there is a report that unsupervised disentanglement learning is impossible without inductive bias Locatello et al. (2019).
Recently, such inductive bias has been introduced in various perspectives on transformation of latent vector space. Intel-VAE Miao et al. (2022) proposed the benefit of invertible transformation of the space to another latent space to provide better data representation, which includes hierarchical representations. Group theory based bias also shows significant improvement on disentanglement Zhu et al. (2021); Yang et al. (2022), whose definition follows Higgins et al. (2018a), which is based on the group theory. The works show that equivariant transformation between input and latent vector space has a key role of disentanglement.
Inspired by the above works, we propose a Multiple Invertible and partial-equivariant transformation (MIPE-transformation) method, which is simply insertable to VAEs. The partial-equivariance is defined by Romero & Lohit (2022) and we follow its definition: $f(g\cdot x)=g\cdot f(x)$ $\forall x\in\mathcal{X},\forall g\in G^{\prime},G^{\prime}\subset G$, where $G^{\prime}$ is a subset of group $G$ ($f$ is partially equivariant to group $G$). First, we assume that an encoder is partial-equivariant and we call it an encoder equivariance condition. The method adopts the matrix exponential to hold the invertible property of latent-to-latent (L2L) vector transformation. Then, we constrain the L2L transformation to a symmetric matrix exponential to be partial-equivariant to a subgroup between latent and transformed latent space. Because it extends the encoder to be partial-equivariant to a subgroup between input and transformed latent space. The IPE-transformation generates an uncertain form of latent vector distributions, so we provide a training procedure to force them to be close to an exponential family, called exponential family conversion (EF-conversion). This conversion enables the uncertain distribution to work in the typical training framework of VAEs. Then, we mathematically show that the multiple uses of IPE-transformation work as $\beta$ parameters Higgins et al. (2017) controlled for enhancing disentanglement learning. Also, we propose the implicit semantic mask to induce a semantic mask in the latent vector space, different to Yang et al. (2020). In experiments with quantitative and qualitative analysis, MIPE-transformation shows significant improvement in disentangled representation learning in 3D Cars, 3D Shapes, and dSprites tasks. Our main contributions are summarized as follows.
  1. 1.
    We propose to use a symmetric matrix exponential as a latent-to-latent vector transformation function for inducing inductive bias based on invertible and equivariant properties with mathematical analysis.
  2. 2.
    We provide a training procedure and losses for VAEs to learn unknown latent vector distribution as an approximated exponential family.
  3. 3.
    We propose the novel MIPE-transformation architecture to integrate multiple IPE-transformation and EF-conversion, which is widely applicable to state-of-the-art VAEs.
  4. 4.
    We empirically analyze the properties of MIPE-transformation and validate its effectiveness in disentanglement learning on benchmarks.
Table 1: Terms and Notations
${\bm{z}}$ Latent vector from encoder $\psi(\cdot)$ Invertible function
$\hat{{\bm{z}}}_{m}$ Transformed latent vector by $\psi_{m}(\cdot)$ $\hat{\bm{\epsilon}}_{m}$ Transformed prior samples by $\psi_{m}(\cdot)$
$\bm{\theta}_{\hat{{\bm{z}}}_{m}}$ Natural Parameter of posterior $\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}$ Natural Parameter of prior
$T$ Sufficient Statistics $A$ Log-Normalizer
$\bm{\nu}$ Evidence $D_{\mathrm{KL}}(\cdot||\cdot)$ Kullback-Leibler divergence
$f_{{\bm{x}}}(\cdot)$ Power Density Function $M_{n}(\mathbb{R})$ A set of $n\times n$ real matrix
$GL_{n}(\mathbb{R})$ General Linear Group $Sym_{n}(\mathbb{R})$ A set of $n\times n$ symmetric real matrix
$E_{M}$$\{\mathbf{e}^{M}|M\in M_{n}(\mathbb{R})\}$ $E_{S}$$\{\mathbf{e}^{S}|S\in Sym_{n}(\mathbb{R})\}$
$G_{S}$ $G_{S}:(\mathbf{e}^{S},\ast)$ $G_{I}$ Group of input space for symmetries
$G_{L}$ Group of latent space for symmetries $J$ $G_{S}\cap G_{L}$
$\psi_{M}(\cdot)$ $\psi_{M}(\cdot)\in M_{n}(\mathbb{R})$ $\psi_{E_{M}}(\cdot)$ $\psi_{E_{M}}(\cdot)\in E_{M}$
$\psi_{E_{S}}(\cdot)$ $\psi_{E_{S}}(\cdot)\in E_{S}$ $\mathbf{0}$ zero vector
$\mathbf{0}_{n,n}$ n by n zero matrix $\mathcal{X}$ Input space
$\mathcal{Z}$ Latent vector space $\hat{\mathcal{Z}}$ Transformed latent vector space
$\Xi$ $G_{I}\times G_{L}\rightarrow G_{L}$ $\Gamma$ $G_{L}\times G_{T}\rightarrow G_{T}$
$\Xi^{J}$ $G_{I}^{J}\times G_{L}^{J}\rightarrow G_{L}^{J}$ $\Gamma^{J}$ $G_{L}^{J}\times G_{T}^{J}\rightarrow G_{T}^{J}$
2 Related Work
Recently, various works have focused on the unsupervised disentanglement learning. Previous works are based on Bengio et al. (2013) definition. One of the branches is InfoGAN Chen et al. (2016) based works such as IB-GAN Jeon et al. (2021) implement extra regularizer to improve informativeness Eastwood & Williams (2018). The other branch is based on the VAE. $\beta$-VAE Higgins et al. (2017) penalizes Kullback-Leibler divergence (KL divergence) with weighted hyper-parameters. Factor VAE Kim & Mnih (2018) and $\beta$-TCVAE Chen et al. (2018) are trained with total correlation (TC) to make independent dimensions on a latent vector with discriminator and divided components of KL divergence term. Differently, we consider the recent disentanglement definition based on group theory Higgins et al. (2018a).
Following the definitions of disentangled representation learning by group theory, several works have emphasized equivariant and improved disentangled representation learning. Commutative Lie Group VAE (CLG-VAE) Zhu et al. (2021) proposed direct mapping of the latent vector into Lie algebra to obtain group structure (inductive bias) with constraints: commutative and hessian loss. Furthermore, Groupified VAE Yang et al. (2022) utilizes Spatial Broadcast Decoder Watters et al. (2019) to implement an equivariant function to the cyclic group with guaranteeing commutativity and invertibility of group actions. Topographic VAE Keller & Welling (2021) combines Student’s-t distributions and variational inference. It enforces rotated latent vectors to be equivariant. On the other hand, we apply unrestricted prior and posterior for disentanglement learning.
There are several inductive biases to learning unsupervised disentanglement, such as group theory based and sequential order. In this section, we briefly discuss sequential order inductive bias even though its method is considered in different domains such as text and video frames. To individualize the static (time-invariant) and dynamic (time-variant), Li & Mandt (2018); Bai et al. (2021) proposed the latent variables one ($f$) is only dependent on the given times series datasets $x_{1:T}$, and the other ($\mathbf{z}_{1:T})$ is dependent on the $x_{1:T}$ and $f$. Moreover Bai et al. (2021) propose the novel ELBO with maximizing mutual information between the input and the latent vectors. These works empirically show that sequential order which includes separated latent vectors improves unsupervised disentanglement learning with diverse qualitative analysis. Differently in group theory based approaches, the proposed methods consider equivariant function between input and latent vector space.
Other VAE approaches implement other prior from Gaussian distribution to transformed Gaussian distribution, Gaussian mixture distribution Dilokthanakul et al. (2016a) or von Mises-Fisher distribution Davidson et al. (2018). InteL-VAE Miao et al. (2022) shows that transformed Gaussian distribution by the invertible function trains hierarchical representation with manual function. We show more clear relation of invertibility to disentanglement and improve VAEs to use its unrestricted form of prior.
Invertible and equivariant Deep Neural Networks have been investigated with normalizing flows. As proven by Xiao & Liu (2020), utilized matrix exponential on Neural networks is invertible, but it only provides mathematical foundations of the tra nsformation. Matrix exponential is utilized to implement an invertible and equivariant function to improve the generative flow compare to linear function Hoogeboom et al. (2020). To specify the exponential familyt, other works contribute uncertainty of exponential family distribution with Bayesian update Charpentier et al. (2020; 2022b). In addition, Ranganath et al. (2014) hierarchically controls the natural parameter across the layers and determines the exponential family distribution with the moment of sufficient statistic. In our work, we show how to use it for disentanglement learning.
Figure 1: The overall architecture of our proposed MIPET-VAE. The invertible and partial-equivariant function $\psi(\cdot)$ for L2L transformation consists of a symmetric matrix exponential to be 1) invertible and 2) partial-equivariant. Then 3) EF conversion module converges the distribution of unrestricted $\hat{{\bm{z}}}$ to be EF with $\mathcal{L}_{el}$ loss. Also, it applies KL divergence loss ($\mathcal{L}_{kl}$) between the transformed posterior and prior, which are expressed by the power density function of EF. In the last, EF conversion reduces the computational error ($\mathcal{L}_{cali}$) between approximated and true KL divergence. 4) The reddish color represents the integration parts. The blue figures represent each property. The details of the gray box are in Figure 2.
3 Method
The overview of a VAE equipped with MIPE-transformation is shown in Figure 1. The MIPE-transformation has three main components: 1) IPE-transformation Unit to transform latent vectors with invertible and partial-equivariant properties, 2) EF-conversion Unit to extend VAEs to learn the exponential family distribution of latent vectors, and 3) integrated training and generation process for multiple uses of IPE-transformation and EF-conversion.
3.1 Invertible and Partial-Equivariant Function for L2L Transformation
3.1.1 Invertible Property by Using Matrix Exponential
To guarantee the invertible property of IPE-transformation, we use a function $\psi(\cdot)=\mathbf{e}^{{\bm{M}}}\ast\cdot$ for the transformation, where ${\bm{M}}$ is in $n\times n$ real number matrix set $M_{n}(\mathbb{R})$ Xiao & Liu (2020). The operator $\ast$ is matrix multiplication, and $\mathbf{e}^{{\bm{M}}}=\sum_{k}^{\infty}\frac{{\bm{M}}^{k}}{k!}$. Our motivation is to use the benefits of injecting explicit inductive bias for disentanglement Locatello et al. (2019); Miao et al. (2022). InteL-VAE effectively extracts hierarchical representation, which includes low-level features (affect to a specific factor) and high-level features (affect to complex factors) with an invertible transformation function Miao et al. (2022).
Figure 2: $G_{I}$, and $G_{L}$ are obtained through encoder $q_{\phi}$ (encoder equivariance condition). The left side figure shows the relation between each space and symmetries. If $\psi(\cdot)$ is equivariant function over all $G_{L}$, and $G_{T}$, then there exist $\Gamma$, where $\Gamma:G_{L}\rightarrow G_{T}$, and $\Xi\circ\Gamma:G_{I}\rightarrow G_{T}$. However, unrestricted $\psi(\cdot)$ has no guarantee to be partial- or full-equivariant. The red arrows represent our method: L2L transformation guarantees $\Gamma^{J}:G_{L}^{J}\rightarrow G_{T}^{J}$, and $\Xi^{J}\circ\Gamma^{J}:G_{I}^{J}\rightarrow G_{T}^{J}$, given the encoder equivariance condition $\Xi:G_{I}\rightarrow G_{L}$.
3.1.2 Why Should L2L Transformation Be Equivariant?
Let’s consider equivariant function between the input and transformed latent vector space, directly used for a decoder in the VAE frameworks. All L2L transformations do not extend the encoder equivariance condition to the relation between input and transformed latent space. This problem is more precisely shown in Figure 2, which illustrates partial equivariance condition over the input space $\mathcal{X}$, latent vector space $\mathcal{Z}$, and its transformed latent vector space $\hat{\mathcal{Z}}$ with a corresponding group of symmetries $G_{I}$, $G_{L}$, and $G_{T}$, respectively. In the VAEs literature, it has not been reported to restrict L2L transformation to guarantee equivariant function between two spaces, so we propose a solution to guarantee at least a part of symmetries to be equivariant.
3.1.3 Equivariance Property with Symmetric Matrix Exponential
To enhance the equivariance of L2L transformation, we set ${\bm{M}}$ of $\psi(\cdot)$ to a symmetric matrix. We show that 1) a group with the constraint guarantees equivariance of $\psi(\cdot)$ over the specific group, 2) $\psi(\cdot)$ being equivariant over subset of symmetries between the input space and transformed latent vector space, and 3) the constraint increases the probability of $\psi(\cdot)$ to be in the group (equal to be equivariant over the subset of symmetries).
We particularly call the transformations as symmetries Higgins et al. (2022) to distinguish them from IPE- and I2L-transformations. For the generality of our method, we consider an arbitrary VAE model that has no restriction on creating intersections to any set as Figure 2.
In the next, we show that matrix exponential with symmetric matrix partially preserves encoder equivariance condition better than other matrices.
Proposition 3.1.
Any $\psi(\cdot)\in G_{S}$, notated as $\psi_{G_{S}}(\cdot)$, is equivariant to group $G_{S}$.
Proof. The group $G_{S}$ is closed to matrix multiplication, and its element is always a symmetric matrix by definition. Then, any two elements in $G_{S}$ are commutative because if matrix multiplication of two symmetric matrices is symmetric then both are commutative. As a result, $\psi_{G_{S}}(\cdot)$ and group elements of $G_{S}$ are commutative ($G_{S}$ is an abelian group). Because of the commutativity, $\psi_{G_{S}}(g_{s}\circ{\bm{z}})=\mathbf{e}^{{\bm{S}}}g_{s}{\bm{z}}=g_{s}\mathbf{e}^{{\bm{S}}}{\bm{z}}=g_{s}\circ\psi_{G_{S}}({\bm{z}})$ for $g_{s}\in G_{S}$ if the group action $\circ$ is set to matrix multiplication, where $\psi_{G_{S}}\in G_{S}$. This equation satisfies the general definition of an equivariant function that a function $f(\cdot)$ is equivariant if $f(g\circ{\bm{z}})=g\circ f({\bm{z}})$ for all $g$ in a group $G$ by matching $f$, $g$, and $G$ to $\psi_{G_{S}}$, $g_{s}$, and $G_{S}$, respectively. $\blacksquare$
Proposition 3.2.
If $q_{\phi}$ is equivariant over defined on group of symmetries $G_{I}^{a}$ and $G_{L}^{a}$, then $\psi_{G_{S}}(q_{\phi}(\cdot))$ is equivariant to symmetries in $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$ by the equivariance of $q_{\phi}$.
Proof. The function $\psi_{G_{S}}(\cdot)$ is an equivariant function over group elements in $G_{S}\cap G_{L}$ by Proposition 3.1. Then, the composite function, $\psi_{G_{S}}(\cdot)$ and $q_{\phi}$, is an equivariant function of $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$. Let $g_{L}^{a}$ be a group element in $G_{S}\cap G_{L}$, and $g_{I}^{a}$ is a group element in $G_{I}$ corresponding to $G_{S}\cap G_{L}$. More details are in Appendix C.1.
$\blacksquare$
Figure 3: Equivariant map: $\mathcal{X}$, $\mathcal{Z}$, and $\hat{\mathcal{Z}}$ are input space, latent vector space, and transformed latent vector space by L2L transfomration function $\psi(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. respectively. ${\bm{x}}\in\mathcal{X}$, ${\bm{z}}\in\mathcal{Z}$, and $\hat{{\bm{z}}}\in\hat{\mathcal{Z}}$.
Then,
Proposition 3.3.
$Pr(\psi_{E_{S}}(\cdot)\in G_{S})>Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})$.
Proof. All $\mathbf{e}^{{\bm{S}}}\in E_{S}$ are in $E_{M}$ since $Sym_{n}(\mathbb{R})\subset M_{n}(\mathbb{R})$. However, $E_{M}\not\subset E_{S}$ because $\mathbf{e}^{{\bm{S}}}$ is always symmetric, but $\mathbf{e}^{{\bm{M}}}$ can be an asymmetric matrix. Therefore $E_{M}\not\subset E_{S}$. Therefore, the probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{S})}$ is greater than $Pr(\psi_{E_{M}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{M})}$. In the same way, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(M_{n}(\mathbb{R}))}$ because $E_{M}\subset M_{n}(\mathbb{R})$ and non-invertible functions are only in $M_{n}(\mathbb{R})$. $\blacksquare$
Therefore, $\psi_{E_{S}}$ clearly increases the probability of preserving a certain type of equivariance compared to unrestricted $\psi$ functions.
The conditional probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})$, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})$, and $Pr(\psi_{M}(\cdot)\in G_{S})$ is changed by the distribution of the observation of $\psi(\cdot)$, which depends on the model parameters. However, the inequality $Pr(\psi_{E_{S}}(\cdot)\in G_{S})>Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})$ is not changed regardless of the distribution of observation of $\psi(\cdot)$. We empirically validate the impact of equivariance with the uncertain $P(\cdot)$ to disentanglement in Section D.3.
3.1.4 Relation Between $\psi(\cdot)$ and Disentanglement
In addition to invertible and partial-equivariant properties, our IPE-transformation also guarantees zero Hessian matrix, which enhances disentanglement without any additional loss of Peebles et al. (2020). Hessian matrix of the transformation $\nabla_{\bm{z}}^{2}\psi({\bm{z}})=\nabla_{\bm{z}}(\nabla_{\bm{z}}\mathbf{e}^{{\bm{M}}}{\bm{z}})=0$ because of the irrelevance of ${\bm{M}}$ to ${\bm{z}}$. By this property, $\psi(\cdot)$ leads that independently factorizes each dimension Peebles et al. (2020), and it injects group theory based inductive bias simultaneously. This is because the group decomposition of ${\bm{z}}$ space $G=G_{1}\times G_{2}\times\cdots\times G_{k}$ corresponds to group decomposition of the transformed latent vector $\hat{{\bm{z}}}$ space $G^{\prime}=G_{1}^{\prime}\times G_{2}^{\prime}\times\cdots\times G_{k}^{\prime}$ such that each $G_{i}^{\prime}$ is fixed by the action of all the $G_{j}$ for $j\neq i$ Yang et al. (2022); Higgins et al. (2018b). This correspondence of decomposition is expected to transfer the independence between dimensions of ${\bm{z}}$ to the space of $\hat{{\bm{z}}}$ Higgins et al. (2018a).
3.2 Exponential Family Conversion for Unknown Prior
In VAE frameworks, the Gaussian normal distribution is applied as a prior. However, a prior from data is usually unknown and may not follow the Gaussian distribution Miao et al. (2022). As a solution, we present a training procedure for VAEs to build an exponential family distribution from a latent variable of an arbitrary distribution. Then, we introduce training losses obtained from the unit IPE-transformation function and EF-conversion.
3.2.1 Elements of Exponential Family Distribution Settings
First, we set sufficient statistics $T(\cdot)$, log-normalizer $A(\cdot)$, and carrier or base measure $B(\cdot)$ are deterministic functions by maximizing conjugate prior for parameter $\mathcal{\xi}$. To determine the natural parameter of posterior and prior $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$, and $\hat{\bm{\epsilon}}_{m}$, we use a natural parameter generator (NPG) designed by multi-layer perceptron Charpentier et al. (2022a). As introduced in Bishop (2006); Charpentier et al. (2022a), we assume exponential family always admits a conjugate prior:
$q(\bm{\theta}|\mathcal{\xi},\bm{\nu})=\text{exp}(\bm{\nu}\bm{\theta}^{\intercal}\mathcal{\xi}-\bm{\nu}A(\bm{\theta})+B^{\prime}(\mathcal{\xi},\bm{\nu})),$ (1)
where $B^{\prime}(\cdot)$ is a normalize coefficient and $\nu$ is evidence, and it is expressed by prior natural parameter $\mathcal{\xi}$. However, generated natural parameter $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$ is not guaranteed as the appropriate parameter of the exponential family corresponds to conjugate prior. To satisfy this condition, we assume observation is a set of independent identically distributed, then Eq. 13 is modified: $p(\mathbf{X}|\bm{\theta})=\prod_{n=1}^{N}h(\mathbf{x}_{n})\text{exp}(\bm{\theta}^{\intercal}\sum_{n=1}^{N}T(\mathbf{x}_{n})-A(\bm{\theta}))$ Bishop (2006), where observation $\mathbf{X}=\{\mathbf{x}_{1},\cdots\mathbf{x}_{N}\}$. In the next, we multiply the modified formation by the prior Eq. 1 to obtain the posterior distribution Bishop (2006) as Eq. 2.
3.2.2 Distribution Approximation As an Exponential Family
The procedure represents a posterior distribution in the exponential family by adopting the following form:
$p(\bm{\theta}|\mathbf{X},\mathcal{\xi},\bm{\nu})\propto\text{exp}(\bm{\theta}^{\intercal}(\sum_{n=1}^{N}T(\mathbf{x}_{n})+\bm{\nu}\mathcal{\xi})-A(\bm{\theta})),$ (2)
where sufficient statistics $T(\cdot)$ and log-normalizer, $A(\cdot)$ are known functions, samples $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\}$ from distribution, and natural parameter of posterior $\bm{\theta}$ and of prior $\mathcal{\xi}$ Bishop (2006). The functions $T(\cdot)$, and $A(\cdot)$ are deterministic functions to maximize posterior distribution. The evidence is implemented as learnable parameters $\bm{\nu}\in\mathbb{R}^{n\times n}$. The natural parameter is generated by a multi-layer perceptron as Charpentier et al. (2022a). This general form approximating an exponential family distribution with learnable parameters can extend VAEs to use a wider distribution for latent variables by simply matching $\mathbf{X}$ to generated latent variables. After IPE-transformation, we can apply the form by using the $\hat{{\bm{z}}}_{m}$, $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$, and $\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}$ for $\mathbf{X}$, $\bm{\theta}$, and $\mathcal{\xi}$, respectively.
3.2.3 EF Similarity Loss
We added a loss to converge the unrestricted distributions of $\hat{{\bm{z}}}$ to the power density function of the exponential family by constraining the posterior maximization as:
$\begin{split}\text{maximize}&~{}\log p(\bm{\theta}_{\hat{{\bm{z}}}_{m}}|\hat{{\bm{z}}}_{m},\bm{\theta}_{\hat{\bm{\epsilon}}_{m}},\bm{\nu}_{m})~{}~{}~{}\text{s.t.}~{}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\geq 0\end{split}$ (3)
$\begin{split}\Rightarrow\mathcal{L}_{s}(\hat{{\bm{z}}}_{m},\hat{\bm{\epsilon}}_{m})&=\log p(\bm{\theta}_{\hat{{\bm{z}}}_{m}}|\hat{{\bm{z}}}_{m},\bm{\theta}_{\hat{\bm{\epsilon}}_{m}},\bm{\nu}_{m})+\bm{\lambda}_{m}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{\mathbf{x}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\end{split}$ (4)
$\begin{split}\Rightarrow\mathcal{L}_{el}\coloneqq||\nabla_{\hat{{\bm{z}}}_{m},\hat{\bm{\epsilon}}_{m},\bm{\lambda}_{m}}\mathcal{L}_{s}||_{2}^{2}=0.\end{split}$ (5)
The notation $\bm{\theta}_{k}$ is a generated natural parameter by a given $k\in\{\hat{{\bm{z}}},\hat{\bm{\theta}}\}$, and $f_{{\bm{x}}}({\bm{x}}|\bm{\theta})$ is a power density function of the exponential family. Moreover, $\bm{\lambda}_{m}$ is a trainable parameter for optimizing the Lagrange multiplier, and $D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))$ is a KL divergence of the exponential family.
3.2.4 KL Divergence for Evidence of Lower Bound
The KL divergence of Gaussian distribution Kingma & Welling (2013) is computed using mean and variance, which are the parameters of a Gaussian distribution. To introduce a loss as the KL divergence of Gaussian distribution, we compute KL divergence of the exponential family in Eq. 2 using the learnable parameter $T(\cdot)$ and $A(\cdot)$ with given natural parameter $\bm{\theta}_{\hat{{\bm{z}}}}$ and $\bm{\theta}_{\hat{\bm{\epsilon}}}$, expressed as:
$\begin{split}\mathcal{L}_{kl}&\coloneqq D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\\ &=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{{\bm{z}}}}}A(\bm{\theta}_{\hat{{\bm{z}}}})-\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{\bm{\epsilon}}}}A(\bm{\theta}_{\hat{\bm{\epsilon}}}).\end{split}$ (6)
3.2.5 KL Divergence Calibration Loss
To reduce the error between the approximation and true matrix for the matrix exponential Bader et al. (2019), we add a loss to minimize the difference of their KL divergence measured by mean squared error (MSE) as:
$\begin{split}\mathcal{L}_{cali}=\text{MSE}(&D_{\mathrm{KL}}(q_{\phi}({\bm{z}}|{\bm{x}})||p_{\theta}({\bm{z}})),D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))),\end{split}$ (7)
which is the KL divergence calibration loss ($\mathcal{L}_{cali}$).
3.2.6 Implicit Semantic Mask
We propose an implicit semantic mask to improve disentanglement learning. We apply mask matrix $\mathcal{M}$ which consists of 0 or 1 element to log-normalizer to prevent less effective weight flow as:
$\mathcal{M}_{ij}=\begin{cases}1\text{ if }|\mathcal{W}_{ij}|\geq\mathbf{\mu}_{|\mathcal{W}_{ij}|}-\lambda\mathbf{\sigma}_{|\mathcal{W}_{ij}|}\\ 0\text{ otherwise}\end{cases},$ (8)
where $\mathcal{W}$ is the weight of log-normalizer, $\lambda$ is a hyper-parameter, $\mathbf{\mu}_{|\mathcal{W}_{ij}|}$, and $\mathbf{\sigma}_{|\mathcal{W}_{ij}|}$ are the mean, and standard deviation of weight respectively. Previous work Yang et al. (2020) utilizes a semantic mask in input space directly, but we inject the semantic mask implicitly on the latent space.
3.3 Integration for Multiple IPE-Transformation and EF-Conversion
We mathematically extend IPE-transformation to MIPE-transformation, which is the equivalent process of $\beta$-VAE to enhance disentanglement. Each IPE-transformation function operates independently, then the reconstruction error for objective function is defined as:
$\begin{split}\mathcal{L}_{rec}&\coloneqq\frac{1}{k}\sum_{i=1}^{k}\Bigg{[}\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg{]}=\frac{1}{k}\sum_{i=1}^{k}E_{q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}})),\end{split}$ (9)
where $\hat{{\bm{z}}}_{i}=\psi_{i}({\bm{z}})$. Therefore, we define ELBO as:
$\begin{split}\mathcal{L}^{\prime}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})&=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi,\psi_{i}}({\bm{z}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}}))-\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})||p_{\psi_{i}}({\bm{z}})).\end{split}$ (10)
However, following Eq. 10, k samples are generated, and each sample is disentangled for different factors. We implement the output as the average of the sum of the k samples to obtain a single sample with a superposition effect from k samples. Moreover, the KL divergence term in Eq. 10 represents that increasing number of MIPE-transformation is equal to increasing $\beta$ hyper-parameter in $\beta$-VAE Higgins et al. (2017).
The VAEs equipped with MIPE-transformation (MIPET-VAEs) can be trained with the following loss:
$\begin{split}&\mathcal{L}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})=\mathcal{L}_{rec}-\mathcal{L}_{kl}-\mathcal{L}_{el}-\mathcal{L}_{cali}.\end{split}$ (11)
Table 2: Performance (mean $\pm$ std) of four metrics on dSprites, 3D Shapes, and 3D Cars. The $\alpha=1$ and $\gamma=1$ of $\beta$-TCVAE as Chen et al. (2018).
dSprites Disentanglement Metric
FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
original MIPET original MIPET original MIPET original MIPET
$\beta$-VAE 69.15($\pm$5.88) 74.19($\pm$5.62) 9.49($\pm$8.30) 19.72($\pm$11.37) 2.43($\pm$2.07) 5.08($\pm$2.90) 18.57($\pm$12.41) 28.81($\pm$10.19)
$\beta$-TCVAE 78.50($\pm$7.93) 79.87($\pm$5.80) 26.00($\pm$9.06) 35.04($\pm$4.07) 7.31($\pm$0.61) 7.70($\pm$1.63) 41.80($\pm$8.55) 47.83($\pm$5.01)
CLG-VAE 79.06($\pm$6.83) 81.80($\pm$3.17) 23.40($\pm$7.89) 36.34($\pm$5.55) 7.37($\pm$0.96) 8.03($\pm$0.83) 37.68($\pm$7.83) 44.73($\pm$5.11)
Control-VAE 62.36($\pm$8.62) 67.71($\pm$6.41) 4.36($\pm$2.86) 7.34($\pm$4.10) 2.11($\pm$1.88) 1.93($\pm$1.63) 10.40($\pm$3.42) 15.18($\pm$4.61)
3D Shapes Disentanglement Metric
FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
original MIPET original MIPET original MIPET original MIPET
$\beta$-VAE 71.76($\pm$12.26) 75.19($\pm$8.16) 37.33($\pm$22.34) 47.37($\pm$10.13) 7.48($\pm$4.12) 9.20($\pm$2.44) 52.07($\pm$17.92) 54.95($\pm$8.99)
$\beta$-TCVAE 76.62($\pm$10.23) 80.59($\pm$8.57) 52.93($\pm$20.5) 54.49($\pm$9.44) 10.64($\pm$5.93) 11.58($\pm$3.32) 65.32($\pm$11.37) 66.22($\pm$7.32)
CLG-VAE 77.04($\pm$8.22) 80.17($\pm$8.43) 49.74($\pm$8.18) 53.87($\pm$7.41) 9.20($\pm$2.44) 12.83($\pm$3.01) 57.70($\pm$8.60) 60.74($\pm$7.77)
Control-VAE 71.05($\pm$14.35) 71.89($\pm$8.33) 24.88($\pm$13.68) 32.28($\pm$10.74) 6.60($\pm$3.59) 7.14($\pm$2.09) 40.08($\pm$13.45) 43.06($\pm$8.68)
3D Cars Disentanglement Metric
FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
original MIPET original MIPET original MIPET original MIPET
$\beta$-VAE 89.48($\pm$5.22) 88.95($\pm$5.94) 6.90($\pm$2.70) 7.27($\pm$1.99) 1.30($\pm$0.48) 1.88($\pm$1.12) 19.85($\pm$4.87) 18.90($\pm$4.49)
$\beta$-TCVAE 95.84($\pm$3.40) 96.43($\pm$2.42) 11.87($\pm$2.90) 10.80($\pm$1.22) 1.55($\pm$0.38) 1.88($\pm$1.12) 27.91($\pm$4.31) 26.08($\pm$2.47)
CLG-VAE 86.11($\pm$7.12) 91.06($\pm$5.09) 6.19($\pm$2.42) 8.51($\pm$2.11) 2.06($\pm$0.60) 1.99($\pm$0.93) 16.91($\pm$4.01) 18.31($\pm$2.83)
Control-VAE 88.76($\pm$7.66) 89.10($\pm$6.90) 4.68($\pm$2.67) 5.08($\pm$2.68) 1.16($\pm$0.74) 1.45($\pm$0.86) 14.70($\pm$3.84) 15.22($\pm$4.15)
Table 3: p-value of t-test for original vs MIPET results of Table 2, which are averaged over models (bold: positive and significant, italic: positive but insignificant, normal: lower performance).
p-value VAEs CLG-VAE $\beta$-TCVAEs
FVM MIG SAP DCI FVM MIG SAP DCI FVM MIG SAP DCI
dSprites 0.000 0.000 0.000 0.000 0.030 0.000 0.005 0.000 0.281 0.000 0.170 0.009
3D Shapes 0.080 0.007 0.016 0.191 0.085 0.029 0.000 0.088 0.111 0.383 0.277 0.390
3D Cars 0.659 0.250 0.003 0.583 0.003 0.000 0.630 0.071 0.278 0.923 0.119 0.933
4 Experiment Settings
4.0.1 Models
As baseline models, we select VAE, $\beta$-VAE, $\beta$-TCVAE, and CLG-VAE. These models are compared to their extension to adopt MIPET, abbreviated by adding the MIPET prefix. We apply the proposed method to $\beta$-TCVAE only with the EF similarity loss term because $\beta$-TCVAE penalizes the divided KL divergence terms. We set the same encoder and decoder architecture in each model to exclude the overlapped effects. Also, we follow the same model architecture which are introduced in previous works Kim & Mnih (2018) and model details are in Table 8-8.
4.0.2 Datasets
We compare well-known VAEs to CHIC-VAEs: VAE, $\beta$-VAE, $\beta$-TCVAE, and CLG-VAE on the following data sets with 1) dSprites Matthey et al. (2017) which consists of 737,280 binary $64\times 64$ images of dSprites with five independent ground truth factors(number of values), $i.e.$ shape(3), orientation(40), scale(6), x-position(32), and y-position(32). 2) 3D Shapes Burgess & Kim (2018) which consists of 480,000 RGB $64\times 64\times 3$ images of 3D Shapes with six independent ground truth factors: shape(4) orientation(15), scale(8), wall color(10), floor color(10), and object color(10). 3) 3D Cars Reed et al. (2015) which consists of 17,568 RGB $64\times 64\times 3$ images of 3D Shapes with three independent ground truth factors: car models(183), azimuth directions(24), and elevations(4).
4.0.3 Training
We set 256 mini-batch size in the datasets (dSprites, 3D Shapes, and 3D Cars), Adam optimizer with learning rate $4\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and epochs from $\{30,67,200\}$ as a common setting for all the comparative methods. For the comparison, we follow training and inference on the whole dataset. We train each model for 30, 67, and 200 epochs on the dSprites, 3D Shapes, and 3D Cars, respectively, as introduced in Kim & Mnih (2018); Ren et al. (2022). We tune $\beta$ from $\{1,2,4,10\}$ and $\{4,6\}$ for $\beta$-VAE and $\beta$-TCVAE, respectively. We set the dimension size of the latent vectors from $\{6,10\}$ for 10 on dSprites and 3D Cars datasets and 6 for 3D Shapes, but we set 10 for CLG-VAE because it sets 10 dimensions size on 3D Shapes in Zhu et al. (2021). Regarding the CLG-VAE, we fix $\lambda_{\text{decomp}}$, $\lambda_{\text{hessian}}$, and forward group features as 40, 20, and 0.2, respectively. Because the hyper-parameters showed the best result in Zhu et al. (2021). We set group reconstruction from $\{0.2,0.5,0.7\}$. For Control-VAE, we set the maximum KL divergence value from $\{10,12,\ldots,20\}$. In addition, we set masking ratio $\lambda$ from $\{0.0,0.5,\ldots 2.0,\infty\}$. To check the impact of MIPE-transformation, we do not consider the Groupified VAE because the latter is implemented with an extended decoder (different capacity).
4.0.4 Evaluation
We conduct experiments on NVIDIA A100, RTX 2080 Ti, and RTX 3090. We set 100 samples to evaluate global empirical variance in each dimension and run it a total of 800 times to estimate the FVM score introduced in Kim & Mnih (2018). For the MIG Chen et al. (2018), SAP Kumar et al. (2018), and DCI Eastwood & Williams (2018), we follow default values introduced in Michlo (2021), training and evaluation 100 and 50 times with 100 mini-batches, respectively. We evaluate four disentanglement metrics for a less biased understanding of the actual states of disentanglement.
Table 4: Impact of the number of MIPE-transformation function on the $\beta$-TCVAE and $\beta$-VAE with dSprites, 3D Shapes, and 3D Cars datasets in terms of the four metrics. The blue and red box plots represent each model’s single and multiple IPE-transformation cases, respectively. (A-$n$: MIPET-$\beta$-TCVAE (4), B-$n$: MIPET-$\beta$-TCVAE (6), C-$n$: MIPET-$\beta$-VAE, $n$: the number of MIPE-transformation).
dataset Metrics
FVM IMG SAP DCI
dSprites [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
3D Shapes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
3D Cars [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 5: Impact of the mask (mean±$\pm$±std.) and its ratio $\lambda$ in Eq. 8 on 3D Cars. ($\infty$: no masking case, gray box: the best setting over all metrics, bold text: the best in each metric.) Each model runs with ten random seeds.
ratio $\beta$-VAE (1) CLG-VAE (0.5)
$\lambda$ FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$ FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DIC $\uparrow$
0.0 90.46($\pm$6.50) 4.84($\pm$2.32) 1.29($\pm$0.81) 16.76($\pm$4.68) 90.06($\pm$4.44) 9.28($\pm$2.09) 1.82($\pm$0.82) 19.12($\pm$3.41)
0.5 91.35($\pm$5.52) 5.37($\pm$2.74) 1.17($\pm$0.67) 16.65($\pm$3.76) 88.69($\pm$4.78) 6.90($\pm$1.96) 1.85($\pm$0.67) 17.52($\pm$3.16)
1.0 91.78($\pm$6.20) 4.99($\pm$2.27) 1.36($\pm$0.81) 16.50($\pm$2.53) 83.60($\pm$11.48) 8.12($\pm$3.66) 2.37($\pm$1.50) 17.07($\pm$3.89)
1.5 90.04($\pm$5.88) 7.22($\pm$2.87) 1.36($\pm$0.48) 18.23($\pm$2.84) 84.76($\pm$6.86) 7.70($\pm$2.11) 2.05($\pm$0.73) 17.06($\pm$2.77)
2.0 87.79($\pm$8.88) 4.75($\pm$2.49) 1.01($\pm$0.99) 16.64($\pm$3.75) 85.78($\pm$4.18) 7.83($\pm$1.79) 1.91($\pm$0.96) 17.26($\pm$2.07)
$\infty$ 89.43($\pm$11.72) 3.74($\pm$2.32) 0.77($\pm$0.39) 15.45($\pm$4.59) 82.96($\pm$11.84) 8.07($\pm$2.52) 2.32($\pm$1.02) 17.46($\pm$4.07)
5 Results and Discussion
5.1 Quantitative Analysis
5.1.1 Disentanglement Metrics
We set the number of IPE-transformation functions to be equal to balancing hyper-parameter $\beta$ on $\beta$-VAE because of Eq. 11. The number of IPE-transform functions of $\beta$-TCVAE is 3. However, in the case of CLG-VAE, we set it to 1 because its approach is based on the group theory, not directly controlling a KL divergence term such as $\beta$-VAE. We average each model performance value with 40, 20, 60, and 30 cases in VAEs, $\beta$-TCVAEs, Control-VAE and CLG-VAEs, respectively.
As shown in Table 2, MIPET-VAEs disentanglement performance is broadly improved with four metrics on each dataset. In particular, most FVM results significantly affect the model performance and stability on all datasets. Therefore, our proposed method obtains a specific dimension that corresponds to a specific single factor. These results imply that applied to MIPE-transformation functions on VAEs elaborate disentangled representation learning.
We additionally estimate the p-value of each metrics over models in Table 3. Previous work shows the average case of each models Yang et al. (2022).
[…]
Table 6: Ablation study for the equivariant property (w/o E), and EF-conversion (w/o EF). Each metric is averaged over 40 and 20 settings of $\beta$-VAE and $\beta$-TCVAE, respectively.
3D Shapes $\beta$-VAE $\beta$-TCVAE
MIPET MIPET (w/o E) MIPET (w/o EF) MIPET MIPET (w/o E) MIPET (w/o EF)
FVM 75.19($\pm$8.16) 74.91($\pm$10.46) 22.27($\pm$1.29) 80.59($\pm$8.57) 77.90($\pm$8.66) 66.38($\pm$7.57)
MIG 47.37($\pm$10.13) 47.45($\pm$8.98) 0.28($\pm$0.09) 54.49($\pm$9.44) 51.37($\pm$11.54) 36.08($\pm$17.42)
SAP 9.20($\pm$2.44) 9.43($\pm$2.59) 0.26($\pm$0.07) 11.58($\pm$3.32) 10.23($\pm$3.13) 7.13($\pm$3.09)
DCI 54.95($\pm$8.99) 54.23($\pm$9.05) 0.10($\pm$0.02) 66.22($\pm$7.32) 61.18($\pm$8.87) 56.85($\pm$11.72)
5.1.2 Sensitivity to the Number of IPE-transformation and EF-conversion
We investigate the impact of the MIPE-transformation function. As presented in Table. 4, MIPE-transformation is better than IPE-transformation for disentanglement learning on each dataset. Indeed, MIPET-$\beta$-VAEs results more generally and clearly show the impact of the MIPE-transfomation function. Our derivation in Section C.1 clearly explains MIPE-transformation impact. This result shows the impact of the multiple uses of IPE-transformation and EF-conversion.
5.1.3 Impact of Implicit Semantic Mask
We set masking hyper-parameter $\lambda$ from $\{0.0,0.5,\cdots,2.0,\infty\}$, and each model has different $\lambda$ for best case. In Table 5, VAE and CLG-VAE with masked log-normalizer show better and well-balanced results than the models without masking, which implies improvement of disentanglement.
5.1.4 Ablation Study
We conduct an ablation study to evaluate the separate impact of equivariant property and the EF-conversion. We have already presented the impact of the multiple uses of IPE-transform and EF-conversion in the previous paragraph. We evaluate the impact of the other properties by setting MIPE-transformation 1) without equivariant (w/o E), which is implemented as an asymmetric matrix, and 2) without EF-conversion (w/o EF). To exclude group theory interference with other methods, we select $\beta$-VAE and $\beta$-TCVAE. As the results are shown in Table 6, most of the results show that MIPET-VAEs performance is better than other cases. In particular, MIPET (w/o EF) results are lower than MIPET (w/o E) results and are clearly shown in all cases.
Figure 4: Qualitative results on dSprites. The left-side grids are input images and their variants by changing activations of each dimension of latent vectors. The first row shows input images. The right-side table shows matching pre-defined factors of the dataset (red: MIPET, blue: no MIPET).
Figure 5: Qualitative analysis result of $\beta$-VAE (4.0) with 3D Cars.
5.2 Qualitative Analysis
[…]
In Figure 4, $\beta$-TCVAE struggles with y-position and rotation, as shown on the 6th row, and with scale and shape represented on the 7th row. On the contrary, MIPET-$\beta$-TCVAE separates y-position and rotation factor (10th, and 7th rows), also the activated dimensions of MIPET-$\beta$-TCVAE are not overlapped with each factor. Applied our method on $\beta$-TCVAE shows better disentangled representation on dSprites dataset. These results also show that our proposed method improves disentangled representation learning.
The qualitative analysis with 3D Cars dataset, as show in Figure 5, the left side is the $\beta$-TCVAE result, and it struggles with body, and azimuth factors shown in the 7th row. However, MIPET-$\beta$-TCVAE separates azimuth (6th row) and body (1st row). In particular, MIPET-$\beta$-TCVAE learns color factor (3rd row) which does not exist on $\beta$-TCVAE.
6 Conclusion
In this paper, we address the problem of injecting inductive bias for learning unsupervised disentangled representations. To build the bias in VAE frameworks, we propose MIPE-transformation composed of 1) IPE-transformation for the benefits of invertibility and partial-equivariant for disentanglement, 2) a training loss and module to adapt unrestricted prior and posterior to an approximated exponential family, and 3) integration of multiple units of IPE-transformation function and EF-conversion for more expressive bias. The method is easily equipped on state-of-the-art VAEs for disentanglement learning and shows significant improvement on dSprites, 3D Shapes, and 3D Cars datasets. We expect that our method can be applied to more VAEs, and extended to downstream applications. Our work is limited to holding partial equivariance of I2L transformation, so more direct methods to induce it can be integrated in the future.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (2022R1A2C2012054).
Appendix A Appendix
Appendix B Preliminaries
B.1 Group Theory
Binary operation: Binary operation on a set $S$ is a function that $\ast:S\times S\rightarrow S$, where $\times$ is a cartesian product.
Group: A group is a set $G$ together with binary operation $\ast$, that combines any two elements $g_{a}$ and $g_{b}$ in $G$, such that the following properties:
  • closure: $g_{a},g_{b}\in G\Rightarrow g_{a}\ast g_{b}\in G$.
  • Associativity: $\forall g_{a},g_{b},g_{c}\in G$, s.t. $(g_{a}\ast g_{b})\ast g_{c}=g_{a}\ast(g_{b}\ast g_{c})$.
  • Identity element: There exists an element $e\in G$, s.t. $\forall g\in G,e\ast g=g\ast e=g$.
  • Inverse element: $\forall g\in G,\exists g^{-1}\in G$: $g\ast g^{-1}=g^{-1}\ast g=e$.
Group action: Let $(G,\ast)$ be a group and set $X$, binary operation $\cdot:G\times X\rightarrow X$, such that following properties:
  • Identity: $e\cdot x=x$, where $e\in G,x\in X$.
  • Compatibility: $\forall g_{a},g_{b}\in G,x\in X,(g_{a}\ast g_{b})\cdot x=g_{a}\cdot(g_{b}\cdot x)$.
Equivariant map: Let $G$ be a group and $X_{1},X_{2}$ be two sets with corresponding group action of $G$ in each sets: $T^{X_{1}}_{g},T^{X_{2}}_{g}$, where $g\in G$. Then a function $f:X_{1}\rightarrow X_{2}$ is equivariant if $f(T^{X_{1}}_{g}\cdot X_{1})=T^{X_{2}}_{g}\cdot f(X_{1})$.
Partial Equivariance:Romero & Lohit (2022): Let subset of $G$ be $\Upsilon\subset G$, then $f$ is a partially equivariant map to $G$:
$f(T^{X_{1}}_{\upsilon}\cdot X_{1})=T^{X_{2}}_{\upsilon}\cdot f(X_{1}),~{}\text{where}~{}\forall\upsilon\in\Upsilon.$ (12)
Homomorphsim: Let $(G,\cdot),(H,\circ)$ be two groups. If mapping function $h:G\rightarrow H$, s.t. $h(g_{i}\cdot g_{j})=f(g_{i})\circ f(g_{j})$, then $f$ is called homomorphism.
B.2 Exponential Family
Power density function of the exponential family (PDF) generalized formulation:
$\begin{split}f_{{\bm{x}}}({\bm{x}}|\bm{\theta})&=h({\bm{x}})\text{exp}(\bm{\theta}^{\intercal}T({\bm{x}})-A(\bm{\theta}))\\ &=\text{exp}(\bm{\theta}^{\intercal}T({\bm{x}})-A(\bm{\theta})+B({\bm{x}})),\end{split}$ (13)
where sufficient statistics $T(\cdot)$, log-normalizer $A(\cdot)$, and carrier or base measure $B(\cdot)$ are known functions, samples ${\bm{x}}$ from distribution, and natural parameter $\bm{\theta}$.
Appendix C Proof
Proposition C.1.
If $q_{\phi}$ is equivariant over defined on group of symmetries $G_{I}^{a}$ and $G_{L}^{a}$, then $\psi_{G_{S}}(q_{\phi}(\cdot))$ is equivariant to symmetries in $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$ by the equivariance of $q_{\phi}$.
Proof. The function $\psi_{G_{S}}(\cdot)$ is an equivariant function over group elements in $G_{S}\cap G_{L}$ by Proposition 3.1. Then, the composite function, $\psi_{G_{S}}(\cdot)$ and $q_{\phi}$, is an equivariant function of $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$. Let $g_{L}^{a}$ be a group element in $G_{S}\cap G_{L}$, and $g_{I}^{a}$ is a group element in $G_{I}$ corresponding to $G_{S}\cap G_{L}$, and $g_{T}^{a}$ is a group element where corresponding to $G_{S}\cap G_{L}$ on the latent vector space transformed from the original latent vector space. Then, group element $g_{T}^{a}$ is equal to $g_{L}^{a}$:
$\hat{{\bm{z}}}_{1}=\psi_{G_{S}}({\bm{z}}_{1}),\text{and}$ (14)
[…]
$\begin{split}&\text{then}~{}g_{L}^{a}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{a}\psi_{G_{S}}({\bm{z}}_{1})~{}(\because\hat{{\bm{z}}}_{2}=g_{T}^{a}\hat{{\bm{z}}}_{1})\\ &\Rightarrow(g_{L}^{a}-g_{T}^{a})\psi_{G_{S}}({\bm{z}}_{1})=\mathbf{0},\end{split}$ (16)
where $\mathbf{0}$ is a zero vector. Eq. 16 is defined when $\forall{\bm{z}}\in\mathcal{Z}$ by the equivariance definition. In other words, Eq. 16 is satisfied only if the kernel (linear algebra) of $g_{L}^{a}-g_{T}^{a}$, notated as ker$(g_{L}^{a}-g_{T}^{a})$, includes the basis of $\mathbb{R}^{n}$ vector space. If the standard basis of $\mathbb{R}^{n}$ vector space is in ker$(g_{L}^{a}-g_{T}^{a})$, then $(g_{L}^{a}-g_{T}^{a})=\mathbf{0}_{n,n}$, where $\mathbf{0}_{n,n}$ is an n by n zero matrix. Other bases of $\mathbb{R}^{n}$ vector space are expressed by the standard basis. Therefore $g_{L}^{a}-g_{T}^{a}=\mathbf{0}_{n,n}$.
Then, $\psi_{G_{S}}(g_{L}^{a}{\bm{z}}_{1})=g_{L}^{a}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{a}\psi_{G_{S}}({\bm{z}}_{1})$. The encoder is an equivariant function over input space $\mathcal{X}$ as $q_{\phi}(g_{I}^{a}{\bm{x}}_{1})=g_{L}^{a}q_{\phi}({\bm{x}}_{1})$. Mixing two equivarience property, we can derive another equivariance relation $g_{T}^{a}\psi_{G_{S}}(q_{\phi}({\bm{x}}_{1}))=\psi_{G_{S}}(q_{\phi}(g_{I}^{a}{\bm{x}}_{1}))$ This result implies that the equivariance between input space and a latent space is preserved for $G_{S}\cap G_{L}$ if the latent vector ${\bm{z}}$ is transformed by $\psi_{G_{S}}.$$\blacksquare$
We show that $\psi_{G_{S}}$ preserves equivariance between $G_{L}^{a}$ and $G_{I}^{a}$. If there exists equivariant function between input and latent vector space, there should be a group $G_{L}$ for a latent space and its corresponding group $G_{I}$ in an input space by definition of equivariance ($q_{\phi}(g_{I}x)=g_{L}q_{\phi}(x)$).
In other words, $\psi_{G_{S}}(\cdot)$ guarantees to preserve the equivariance of I2L-transformation to certain symmetries in $G_{S}\cap G_{L}$ after IPE-transformation as shown in Figure 2.
Let $P(B)$ be the probability of $\psi(\cdot)\in B$ for a subset $B\subset M_{n}(\mathbb{R})$ after VAE training, and $Pr(\psi_{B}\in B^{\prime})$ be the conditional probability of $\psi(\cdot)\in B^{\prime}$ given $\psi(\cdot)\in B$.
Then,
Proposition C.2.
$Pr(\psi_{E_{S}}(\cdot)\in G_{S})>Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})$.
Proof. All $\mathbf{e}^{{\bm{S}}}\in E_{S}$ are in $E_{M}$ since $Sym_{n}(\mathbb{R})\subset M_{n}(\mathbb{R})$. However, $E_{M}\not\subset E_{S}$ because $\mathbf{e}^{{\bm{S}}}$ is always symmetric, but $\mathbf{e}^{{\bm{M}}}$ can be an asymmetric matrix. All elements of $E_{S}$ are symmetric because of the matrix exponential property that $\mathbf{e}^{{\bm{M}}^{\intercal}}=(\mathbf{e}^{{\bm{M}}})^{\intercal}$. If ${\bm{M}}$ is a symmetric matrix then $\mathbf{e}^{{\bm{M}}^{\intercal}}=\mathbf{e}^{{\bm{M}}}=(\mathbf{e}^{{\bm{M}}})^{\intercal}$. Therefore, if ${\bm{M}}$ is symmetric then the exponential of ${\bm{M}}$ is also symmetric. We show a counter example to $E_{M}\subset E_{S}$. When ${\bm{M}}=\begin{bmatrix}1&1\\ 0&1\end{bmatrix}$,
$\begin{split}\mathbf{e}^{{\bm{M}}}&=\sum_{k=0}^{\infty}\frac{1}{k!}{\bm{M}}^{k}\\ &=I+\begin{bmatrix}1&1\\ 0&1\end{bmatrix}+\frac{1}{2!}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}^{2}+\cdots+\frac{1}{(n-1)!}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}^{(n-1)}+\cdots\\ &=I+\begin{bmatrix}\sum_{n=0}^{\infty}\frac{1}{n!}&1+\sum_{n=0}^{\infty}\frac{1}{(n-1)!}\\ 0&\sum_{n=0}^{\infty}\frac{1}{n!}\end{bmatrix}\\ &=\begin{bmatrix}1+e&1+e\\ 0&1+e\end{bmatrix}.\end{split}$ (17)
The matrix $\mathbf{e}^{{\bm{M}}}$ is asymmetric and not in $E_{S}$. Therefore $E_{M}\not\subset E_{S}$. Therefore, the probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{S})}$ is greater than $Pr(\psi_{E_{M}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{M})}$. In the same way, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(M_{n}(\mathbb{R}))}$ because $E_{M}\subset M_{n}(\mathbb{R})$ and non-invertible functions are only in $M_{n}(\mathbb{R})$. $\blacksquare$
C.0.1 KL Divergence for Evidence of Lower Bound
The KL divergence of Gaussian distribution Kingma & Welling (2013) is computed using mean and variance, which are the parameters of a Gaussian distribution. To introduce a loss as the KL divergence of Gaussian distribution, we compute KL divergence of the exponential family in Eq. 2 using the learnable parameter $T(\cdot)$ and $A(\cdot)$ with given natural parameter $\bm{\theta}_{\hat{{\bm{z}}}}$ and $\bm{\theta}_{\hat{\bm{\epsilon}}}$, expressed as:
$\begin{split}\mathcal{L}_{kl}&\coloneqq D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\\ &=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{{\bm{z}}}}}A(\bm{\theta}_{\hat{{\bm{z}}}})-\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{\bm{\epsilon}}}}A(\bm{\theta}_{\hat{\bm{\epsilon}}}).\end{split}$ (18)
Because $D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))$ is followed as:
$\begin{split}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\mathsf{d}{\bm{x}}\\ &-\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\mathsf{d}{\bm{x}}.\end{split}$ (19)
We designed sufficient statistics as matrix multiplication (multi-layer perceptron). Then,
$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\mathsf{d}{\bm{x}}&=\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\cdot\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}[\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\mathbf{T}({\bm{x}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+B({\bm{x}})]\text{d}{\bm{x}}\\ &=-A(\bm{\theta}_{\hat{{\bm{z}}}})\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})[\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\mathbf{T}({\bm{x}})+B({\bm{x}})]\text{d}{\bm{x}}\\ &=-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}B({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}},\end{split}$ (20)
and
$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\mathsf{d}{\bm{x}}&=-A(\bm{\theta}_{\hat{\bm{\epsilon}}})\\ &+\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}B({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}.\end{split}$ (21)
$\begin{split}\therefore D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})\\ &+\bm{\theta}_{\hat{{\bm{z}}}}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &-\bm{\theta}_{\hat{\bm{\epsilon}}}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}.\end{split}$ (22)
The mean of the sufficient statistic is followed as:
$\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta})\text{d}{\bm{x}}=\frac{\partial A^{\ast}(\bm{\theta})}{\partial\bm{\theta}}\approx\frac{\partial A(\bm{\theta})}{\partial\bm{\theta}}~{}~{}~{}~{}~{}\because A^{\ast}(\bm{\theta})=\bm{\theta}^{\intercal}A^{\ast},$ (23)
where $A^{\ast}(\cdot)$ is a true log-partition function of the exponential family (ideal case of $A(\cdot)$). However, estimating $A^{\ast}$ is difficult, and there is no direct method without random samplings, such as mini-batch weighted sampling or mini-batch stratified sampling Chen et al. (2018). Then, we approximate $A^{\ast}$ to $A$, and train $A$ to be close to $A^{\ast}$. Consequently, we obtain KL divergence of the exponential family as:
$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}(\mathbf{x}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})&=-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{{\bm{z}}}})}{\partial\bm{\theta}_{\hat{{\bm{z}}}}}\\ &+\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})B({\bm{x}})\text{d}{\bm{x}},\end{split}$ (24)
$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}(\mathbf{x}|\bm{\theta}_{\hat{\bm{\epsilon}}})&=-Z(\bm{\theta}_{\hat{\bm{\epsilon}}})+\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{\bm{\epsilon}}})}{\partial\bm{\theta}_{\hat{\bm{\epsilon}}}}\\ &+\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})B({\bm{x}})\text{d}{\bm{x}}.\end{split}$ (25)
Therefore, the final Kullback-Leibler divergence of exponential family is followed as:
$\begin{split}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{{\bm{z}}}})}{\partial\bm{\theta}_{\hat{{\bm{z}}}}}-\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{\bm{\epsilon}}})}{\partial\bm{\theta}_{\hat{\bm{\epsilon}}}}.\end{split}$ (26)
C.1 Integration for Multiple IPE-Transformation and EF-Conversion
We mathematically extend IPE-transformation to MIPE-transformation, which is the equivalent process of $\beta$-VAE to enhance disentanglement. Each IPE-transformation function operates independently, then the reconstruction error for objective function is defined as:
$\begin{split}\mathcal{L}_{rec}&\coloneqq\frac{1}{k}\sum_{i=1}^{k}\Bigg{[}\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg{]}\\ &=\frac{1}{k}\sum_{i=1}^{k}E_{q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}})),\end{split}$ (27)
where $\hat{{\bm{z}}}_{i}=\psi_{i}({\bm{z}})$. Becuase the log likelihood of $p({\bm{x}})$ can be derived as follows:
$\log p_{\theta}({\bm{x}})$ $=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}})\mathsf{d}\hat{{\bm{z}}}^{\prime}$ (28)
$=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\mathsf{d}\hat{{\bm{z}}}^{\prime}$ (29)
$\begin{split}&=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\cdot\\ &~{}~{}~{}~{}\Big{[}\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}-\log\frac{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\Big{]}\mathsf{d}\hat{{\bm{z}}}^{\prime}\end{split}$ (30)
$\geqq\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\mathsf{d}\hat{{\bm{z}}}^{\prime}$ (31)
$\begin{split}&=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\cdot\\ &~{}~{}~{}~{}\Big{[}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})+\log\frac{p(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\Big{]}\mathsf{d}\hat{{\bm{z}}}^{\prime},\end{split}$ (32)
[…]
$\begin{split}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})&=\frac{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k}|{\bm{x}})p_{\theta}({\bm{x}})}{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})}\\ &=\frac{p_{\theta}({\bm{x}})\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})}{\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i})}(\because(\hat{{\bm{z}}}_{i}\mathrel{\perp\!\!\!\perp}\hat{{\bm{z}}}_{j}s|{\bm{x}}))\\ &=\prod_{i=1}^{k}\frac{p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})p_{\theta}({\bm{x}}^{\frac{1}{k}})}{p_{\theta}(\hat{{\bm{z}}}_{i})}\\ &=p_{\theta}({\bm{x}})^{-(k-1)}\prod_{i=1}^{k}\frac{p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})p_{\theta}({\bm{x}})}{p_{\theta}(\hat{{\bm{z}}}_{i})}\\ &=p_{\theta}({\bm{x}})^{-(k-1)}\prod_{i=1}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i}),\end{split}$ (33)
where $\hat{{\bm{z}}}_{j}s=\cap_{j=1,j\neq i}^{k}\hat{{\bm{z}}}_{j}$. Therefore,
$\small\begin{split}&\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\Big{[}-(k-1)\log p_{\theta}({\bm{x}})+\prod_{i=1}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\Big{]}\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=-(k-1)\log p_{\theta}({\bm{x}})+\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\prod_{i=j}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{j})\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{split}$ (34)
Then,
$\begin{split}\log p_{\theta}({\bm{x}})&\geq\frac{1}{k}\sum_{i=1}^{k}\Bigg{[}\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg{]}\\ &-\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})}{\prod_{i=1}^{k}p(\hat{{\bm{z}}}_{i})}\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\\ &-\sum_{i=1}^{k}\bigg{[}D_{\mathrm{KL}}(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})||p(\hat{{\bm{z}}}_{i}))\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\bigg{]}\\ &=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})-\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})||p({\bm{z}}_{i}))\\ &=\frac{1}{k}\Big{[}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})-kD_{\mathrm{KL}}(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})||p(\hat{{\bm{z}}}_{i}))\Big{]}.\end{split}$ (35)
Therefore, we define ELBO as:
$\begin{split}\mathcal{L}^{\prime}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})&=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi,\psi_{i}}({\bm{z}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}}))-\\ &\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})||p_{\psi_{i}}({\bm{z}})).\end{split}$ (36)
However, following Eq. 36, k samples are generated, and each sample is disentangled for different factors. We implement the output as the average of the sum of the k samples to obtain a single sample with a superposition effect from k samples. Moreover, the KL divergence term in Eq. 36 represents that increasing number of MIPE-transformation is equal to increasing $\beta$ hyper-parameter in $\beta$-VAE Higgins et al. (2017).
The VAEs equipped with MIPE-transformation (MIPET-VAEs) can be trained with the following loss:
$\begin{split}&\mathcal{L}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})=\mathcal{L}_{rec}-\mathcal{L}_{kl}-\mathcal{L}_{el}-\mathcal{L}_{cali}.\end{split}$ (37)
Appendix D Details of Experimental Environment
D.1 Model Architecture
Table 7: VAE architecture for dSprites dataset.
Encoder Decoder
Input $64\times 64$ binary image input $\in\mathbb{R}^{10}$
$4\times 4$ conv. 32 ReLU. stride 2 FC. 128 ReLU.
$4\times 4$ conv. 32 ReLU. stride 2 FC. $4\times 4\times 64$ ReLU.
$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 64 ReLU. stride 2.
$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 32 ReLU. stride 2.
FC. 128. FC. $2\times 10$ $4\times 4$ upconv. 32 ReLU. stride 2.
$4\times 4$ upconv. 1. stride 2
Table 8: VAE architecture for 3D Shapes, and 3D Cars datasets. For exceptional case, CLG-VAE, we ues ten dimension size on 3D Shapes dataset Zhu et al. (2021).
Encoder Decoder
Input $64\times 64\times 3$ RGB image input $\in\mathbb{R}^{6}$ (3D Shapes), $\mathbb{R}^{10}$ (3D Cars)
$4\times 4$ conv. 32 ReLU. stride 2 FC. 256 ReLU.
$4\times 4$ conv. 32 ReLU. stride 2 FC. $4\times 4\times 64$ ReLU.
$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 64 ReLU. stride 2.
$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 32 ReLU. stride 2.
FC. 256. FC. $2\times 10$ $4\times 4$ upconv. 32 ReLU. stride 2.
$4\times 4$ upconv. 3. stride 2
D.2 Details of Setting
Table 9: Hyper-parameters for dSprites, 3D Cars, and 3D Shapes. The epochs for dSprites and 3D cars are 30 and 200, respectively. lr is learning rate, latent dim is latent vector size, group reconst is group reconstruction, and forward group is forward group pass.
(a) Hyper-parameters for dSprites and 3D Cars
models hyper-parameters values
common batch size 256
epoch $\{30,200\}$
optim Adam
lr 4e-4
lr for MIPET 4e-4
weight decay 1e-4
latent dim 10
$\beta$-VAE ##\## of IE and EF $\{1,2,4,10\}$
$\beta$-TCVAE $\beta$$\{4,6\}$
##\## of IE and EF $\{1,3\}$
$\alpha$, $\gamma$ 1.0
CLG-VAE $\lambda_{\text{decomp}}$ 40
$\lambda_{\text{hessian}}$ 40
forward group 0.2
group reconst $\{0.2,0.5,0.7\}$
(b) Hyper-parameters for 3D Shapes
models hyper-parameters values
common batch size 256
epoch 67
optim Adam
lr 4e-4
lr for MIPET 4e-4
$\beta$-VAE ##\## of IE and EF $\{1,2,4,10\}$
weight decay 0.0
latent dim 6
$\beta$-TCVAE $\beta$$\{4,6\}$
##\## of IE and EF $\{1,3\}$
$\alpha$, $\gamma$ 1.0
weight decay 1e-4
latent dim 6
CLG-VAE $\lambda_{\text{decomp}}$ 40
$\lambda_{\text{hessian}}$ 40
forward group 0.2
group reconst $\{0.2,0.5,0.7\}$
weight decay 0.0
latent dim 10
D.3 Impact of Symmetric Matrix Exponential
Table 10: The ratio of seeds to show better performance with symmetric matrix
dSprites 3D Shapes 3D Cars
0.58 0.56 0.67
We empirically show the benefit of using a symmetric matrix for $\psi$. Table 10 shows the ratio of runs with a symmetric matrix, which shows better performance than unrestricted matrices, to the total 240 (60 models $\times$ 4 metrics) runs for each dataset. All results are higher than 0.5, which implies that the constraint enhances I2L equivariance even with uncertain factors.
D.4 Additional Experiment of Computing Complexity
Table 11: Training complexity.
##\## of IE Complexity
0 $\times$ 1.00
1 $\times$ 0.75
3 $\times$ 0.50
4 $\times$ 0.33
We additionally estimate the computing complexity depending on the number of IPE-transformation. The results are in Table 11 and represent the training time complexity compare to baselines (when the number of IE is equal to 0).
D.5 Qualitative Analysis
[…]
In Figure 8, $\beta$-TCVAE struggles with y-position and rotation, as shown on the 6th row, and with scale and shape represented on the 7th row. On the contrary, MIPET-$\beta$-TCVAE separates y-position and rotation factor (10th, and 7th rows), also the activated dimensions of MIPET-$\beta$-TCVAE are not overlapped with each factor. Applied our method on $\beta$-TCVAE shows better disentangled representation on dSprites dataset. These results also show that our proposed method improves disentangled representation learning. As shown in the Figure 8, $\beta$-VAE struggles with rotation and scale factors in 4thdimension. Also, it struggles with x-position and scale factors in 8th dimension, and x-position and rotation factors in 9th dimension. However, MIPET-$\beta$-VAE only struggles with rotation and shape factors in 5th dimension. As shown in the Figure 8, CLG-VAE struggles with rotation and shape factors in 2nd dimension, and shape and scale factors in 7th dimension. However, MIPET-CLG-VAE separates rotation and shape factors in 10th, and 1st dimensions respectively.
The qualitative analysis with 3D Shapes dataset, as shown in the Figure 11, $\beta$-TCVAE struggles with shape and scale in $5^{th}$ dimension. However MIPET-$\beta$-TCVAE As shown in the Figure 11, CLG-VAE struggles with shape and wall color factors in 4th dimension, and shape and object color factors in 7th dimension. In particular, it struggles with tree factors in 9th dimension. On the other hand, MIPET-CLG-VAE separates shape, wall, and object color factors.
The qualitative analysis with 3D Cars dataset, as show in Figure 11, the left side is the $\beta$-TCVAE result, and it struggles with body, and azimuth factors shown in the 7th row. However, MIPET-$\beta$-TCVAE separates azimuth (6th row) and body (1st row). In particular, MIPET-$\beta$-TCVAE learns color factor (3rd row) which does not exist on $\beta$-TCVAE.
Figure 6: Qualitative results on dSprites. The left-side grids are input images and their variants by changing activations of each dimension of latent vectors. The first row shows input images. The right-side table shows matching pre-defined factors of the dataset (red: MIPET, blue: no MIPET).
Figure 7: Qualitative analysis result of $\beta$-VAE and MIPET-$\beta$-VAE.
Figure 8: Qualitative analysis result of CLG-VAE (0.2) and MIPET-CLG-VAE (0.2) with dSprites.
Figure 9: The Shape is object shape, Orien is an orientation of object, Scale is a scale factor of object, Wall is wall color factor, Floor is floor color, and Object is object color factors. It represents the $\beta$-VAE ($\beta=2$) results.
Figure 10: Qualitative analysis result of CLG-VAE (0.2) and MIPET-CLG-VAE (0.2) with 3D Shapes.
Figure 11: Qualitative analysis result of $\beta$-VAE (4.0) with 3D Cars.

### v3

@@
 Abstract
-Disentanglement learning is a core issue for understanding and re-using trained information in Variational AutoEncoder (VAE), and effective inductive bias has been reported as a key factor.
-However, the actual implementation of such bias is still vague.
-In this paper, we propose a novel method, called Multiple Invertible and partial-equivariant transformation (MIPE-transformation), to inject inductive bias by 1) guaranteeing the invertibility of latent-to-latent vector transformation while preserving a certain portion of equivariance of input-to-latent vector transformation, called Invertible and partial-equivariant transformation (IPE-transformation), 2) extending the form of prior and posterior in VAE frameworks to an unrestricted form through a learnable conversion to an approximated exponential family, called Exponential Family conversion (EF-conversion), and 3) integrating multiple units of IPE-transformation and EF-conversion, and their training.
-In experiments on 3D Cars, 3D Shapes, and dSprites datasets, MIPE-transformation improves the disentanglement performance of state-of-the-art VAEs.
+Disentanglement learning is central to understanding and reusing learned representations in variational autoencoders (VAEs).
+Although equivariance has been explored in this context, effectively exploiting it for disentanglement remains challenging.
+In this paper, we propose a novel method, called Multiple Invertible and Partial-Equivariant Transformation (MIPE-Transformation), which integrates two main parts: (1) Invertible and Partial-Equivariant Transformation (IPE-Transformation), guaranteeing an invertible latent-to–transformed-latent mapping while preserving partial input-to-latent equivariance in the transformed latent space; and (2) Exponential-Family Conversion (EF-Conversion) to extend the standard Gaussian prior to an approximate exponential family via a learnable conversion.
+In experiments on the 3D Cars, 3D Shapes, and dSprites datasets, MIPE-Transformation improves the disentanglement performance of state-of-the-art VAEs.
 1 Introduction
-Disentanglement learning to learn more interpretable representations is broadly useful in artificial intelligence fields such as classification Singla et al.
-(2021), zero-shot learning Tenenbaum (2018), and domain adaptation Li et al.
-(2019); Zou et al.
-(2020).
-The disentangled representation is defined as a change in a single dimension, which corresponds to unique semantic information.
-Several works have been conducted based on this framework.
-A major model for enhancing the disentanglement learning is Variational AutoEncoder (VAE) Kingma & Welling (2013).
-Based on VAE, unsupervised disentangled representation learning has been elaborated Higgins et al.
-(2017); Chen et al.
-(2018); Kim & Mnih (2018); Jeong & Song (2019); Li et al.
-(2020) through the factorizable variations and control of uncorrelatedness of each dimension of representations.
-Moreover, VAE models that handle the shape of prior as a Gaussian mixture Dilokthanakul et al.
-(2016b) or von Mises-Fisher Davidson et al.
-(2018) were also developed, but the disentanglement is still incomplete.
-As a critical point, there is a report that unsupervised disentanglement learning is impossible without inductive bias Locatello et al.
-(2019).
-Recently, such inductive bias has been introduced in various perspectives on transformation of latent vector space.
-Intel-VAE Miao et al.
-(2022) proposed the benefit of invertible transformation of the space to another latent space to provide better data representation, which includes hierarchical representations.
-Group theory based bias also shows significant improvement on disentanglement Zhu et al.
-(2021); Yang et al.
-(2022), whose definition follows Higgins et al.
-(2018a), which is based on the group theory.
-The works show that equivariant transformation between input and latent vector space has a key role of disentanglement.
-Inspired by the above works, we propose a Multiple Invertible and partial-equivariant transformation (MIPE-transformation) method, which is simply insertable to VAEs.
-The partial-equivariance is defined by Romero & Lohit (2022) and we follow its definition: $f(g\cdot x)=g\cdot f(x)$ $\forall x\in\mathcal{X},\forall g\in G^{\prime},G^{\prime}\subset G$, where $G^{\prime}$ is a subset of group $G$ ($f$ is partially equivariant to group $G$).
-First, we assume that an encoder is partial-equivariant and we call it an encoder equivariance condition.
-The method adopts the matrix exponential to hold the invertible property of latent-to-latent (L2L) vector transformation.
-Then, we constrain the L2L transformation to a symmetric matrix exponential to be partial-equivariant to a subgroup between latent and transformed latent space.
-Because it extends the encoder to be partial-equivariant to a subgroup between input and transformed latent space.
-The IPE-transformation generates an uncertain form of latent vector distributions, so we provide a training procedure to force them to be close to an exponential family, called exponential family conversion (EF-conversion).
-This conversion enables the uncertain distribution to work in the typical training framework of VAEs.
-Then, we mathematically show that the multiple uses of IPE-transformation work as $\beta$ parameters Higgins et al.
-(2017) controlled for enhancing disentanglement learning.
-Also, we propose the implicit semantic mask to induce a semantic mask in the latent vector space, different to Yang et al.
-(2020).
-In experiments with quantitative and qualitative analysis, MIPE-transformation shows significant improvement in disentangled representation learning in 3D Cars, 3D Shapes, and dSprites tasks.
-Our main contributions are summarized as follows.
+Disentanglement learning aims to learn interpretable representations and has utility in tasks such as classification (Singla et al., 2021), zero-shot learning (Tenenbaum, 2018), and domain adaptation (Li et al., 2019; Zou et al., 2020).
+A disentangled representation varies along one latent coordinate in a way that corresponds to a single factor of variation while keeping other factors essentially fixed.
+Among existing approaches, the Variational Autoencoder (VAE) (Kingma and Welling, 2013) is a primary framework; unsupervised methods encourage factorized latents and reduce inter-dimension correlation (Higgins et al., 2017; Chen et al., 2018; Kim and Mnih, 2018; Jeong and Song, 2019; Li et al., 2020), but Locatello et al.
+(2019) shows that disentanglement is not identifiable without inductive bias.
+To inject such bias, prior work follows two lines: 1) statistics-driven methods that replace the standard Gaussian prior with richer, often non-Gaussian alternatives.
+For example, Kim et al.
+(2019); Mathieu et al.
+(2018) investigate Gaussian-mixture and Student-$t$ priors and demonstrate limitations of the standard Gaussian prior for disentanglement.
+Similarly, Locatello et al.
+(2020); Togo et al.
+(2025) consider priors beyond the Gaussian family and report improved disentanglement performance.
+However, these approaches are grounded in the original notion of disentangled representation (Bengio et al., 2013) rather than the group-theoretic formulation, and they do not explicitly encode equivariance.
+The second line comprises group-theoretic methods that recast disentanglement in terms of the direct product of groups defined on the latent space (Higgins et al., 2018) by equivariant modeling between input and latent vector space rather than 2D planar symmetries.
+Representative works instantiate the general linear group (GL(n)GL(n)(Zhu et al., 2021; Jung et al., 2024; Winter et al., 2022) or cyclic groups (Yang et al., 2022; Keurti et al., 2023) on latent vector space, and design equivariant VAE objectives that improve disentangled representation quality.
+However, these models typically retain a fixed standard Gaussian prior, which limits statistical flexibility.
+Motivated by these observations, we target two desiderata: 1) preserving partial equivariance between the input and a transformed latent space, and 2) enabling diverse priors.
+We introduce a learnable, invertible latent-to-latent transformation that preserves partial equivariance between inputs and latent space, and via an exponential-family conversion (EF-conversion), maps the transformed latent vectors to flexible (non-Gaussian) priors.
+We refer to the module as the Multiple Invertible and Partial-Equivariant (MIPE) transformation, which can be inserted into existing VAEs.
+Through quantitative and qualitative experiments on 3D Cars, 3D Shapes, and dSprites, MIPE yields consistent improvements in disentangled representation learning. Our contributions are:
   1. 1.
-    We propose to use a symmetric matrix exponential as a latent-to-latent vector transformation function for inducing inductive bias based on invertible and equivariant properties with mathematical analysis.
+    A symmetric matrix exponential latent-to-latent transformation with invertibility and analyzed partial equivariance properties.
   2. 2.
-    We provide a training procedure and losses for VAEs to learn unknown latent vector distribution as an approximated exponential family.
+    A training procedure and losses that allow VAEs to learn an unknown latent distribution approximated by an exponential family.
   3. 3.
-    We propose the novel MIPE-transformation architecture to integrate multiple IPE-transformation and EF-conversion, which is widely applicable to state-of-the-art VAEs.
+    The MIPE architecture that integrates multiple IPE transformations with EF-conversion and can be plugged into state-of-the-art VAEs.
   4. 4.
-    We empirically analyze the properties of MIPE-transformation and validate its effectiveness in disentanglement learning on benchmarks.
-Table 1: Terms and Notations
-${\bm{z}}$ Latent vector from encoder $\psi(\cdot)$ Invertible function
-$\hat{{\bm{z}}}_{m}$ Transformed latent vector by $\psi_{m}(\cdot)$ $\hat{\bm{\epsilon}}_{m}$ Transformed prior samples by $\psi_{m}(\cdot)$
-$\bm{\theta}_{\hat{{\bm{z}}}_{m}}$ Natural Parameter of posterior $\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}$ Natural Parameter of prior
-$T$ Sufficient Statistics $A$ Log-Normalizer
-$\bm{\nu}$ Evidence $D_{\mathrm{KL}}(\cdot||\cdot)$ Kullback-Leibler divergence
-$f_{{\bm{x}}}(\cdot)$ Power Density Function $M_{n}(\mathbb{R})$ A set of $n\times n$ real matrix
-$GL_{n}(\mathbb{R})$ General Linear Group $Sym_{n}(\mathbb{R})$ A set of $n\times n$ symmetric real matrix
-$E_{M}$$\{\mathbf{e}^{M}|M\in M_{n}(\mathbb{R})\}$ $E_{S}$$\{\mathbf{e}^{S}|S\in Sym_{n}(\mathbb{R})\}$
-$G_{S}$ $G_{S}:(\mathbf{e}^{S},\ast)$ $G_{I}$ Group of input space for symmetries
-$G_{L}$ Group of latent space for symmetries $J$ $G_{S}\cap G_{L}$
-$\psi_{M}(\cdot)$ $\psi_{M}(\cdot)\in M_{n}(\mathbb{R})$ $\psi_{E_{M}}(\cdot)$ $\psi_{E_{M}}(\cdot)\in E_{M}$
-$\psi_{E_{S}}(\cdot)$ $\psi_{E_{S}}(\cdot)\in E_{S}$ $\mathbf{0}$ zero vector
-$\mathbf{0}_{n,n}$ n by n zero matrix $\mathcal{X}$ Input space
-$\mathcal{Z}$ Latent vector space $\hat{\mathcal{Z}}$ Transformed latent vector space
-$\Xi$ $G_{I}\times G_{L}\rightarrow G_{L}$ $\Gamma$ $G_{L}\times G_{T}\rightarrow G_{T}$
-$\Xi^{J}$ $G_{I}^{J}\times G_{L}^{J}\rightarrow G_{L}^{J}$ $\Gamma^{J}$ $G_{L}^{J}\times G_{T}^{J}\rightarrow G_{T}^{J}$
-2 Related Work
-Recently, various works have focused on the unsupervised disentanglement learning.
-Previous works are based on Bengio et al.
-(2013) definition.
-One of the branches is InfoGAN Chen et al.
-(2016) based works such as IB-GAN Jeon et al.
-(2021) implement extra regularizer to improve informativeness Eastwood & Williams (2018).
-The other branch is based on the VAE.
-$\beta$-VAE Higgins et al.
-(2017) penalizes Kullback-Leibler divergence (KL divergence) with weighted hyper-parameters.
-Factor VAE Kim & Mnih (2018) and $\beta$-TCVAE Chen et al.
-(2018) are trained with total correlation (TC) to make independent dimensions on a latent vector with discriminator and divided components of KL divergence term.
-Differently, we consider the recent disentanglement definition based on group theory Higgins et al.
-(2018a).
-Following the definitions of disentangled representation learning by group theory, several works have emphasized equivariant and improved disentangled representation learning.
-Commutative Lie Group VAE (CLG-VAE) Zhu et al.
-(2021) proposed direct mapping of the latent vector into Lie algebra to obtain group structure (inductive bias) with constraints: commutative and hessian loss.
-Furthermore, Groupified VAE Yang et al.
-(2022) utilizes Spatial Broadcast Decoder Watters et al.
-(2019) to implement an equivariant function to the cyclic group with guaranteeing commutativity and invertibility of group actions.
-Topographic VAE Keller & Welling (2021) combines Student’s-t distributions and variational inference.
-It enforces rotated latent vectors to be equivariant.
-On the other hand, we apply unrestricted prior and posterior for disentanglement learning.
-There are several inductive biases to learning unsupervised disentanglement, such as group theory based and sequential order.
-In this section, we briefly discuss sequential order inductive bias even though its method is considered in different domains such as text and video frames.
-To individualize the static (time-invariant) and dynamic (time-variant), Li & Mandt (2018); Bai et al.
-(2021) proposed the latent variables one ($f$) is only dependent on the given times series datasets $x_{1:T}$, and the other ($\mathbf{z}_{1:T})$ is dependent on the $x_{1:T}$ and $f$.
-Moreover Bai et al.
-(2021) propose the novel ELBO with maximizing mutual information between the input and the latent vectors.
-These works empirically show that sequential order which includes separated latent vectors improves unsupervised disentanglement learning with diverse qualitative analysis.
-Differently in group theory based approaches, the proposed methods consider equivariant function between input and latent vector space.
-Other VAE approaches implement other prior from Gaussian distribution to transformed Gaussian distribution, Gaussian mixture distribution Dilokthanakul et al.
-(2016a) or von Mises-Fisher distribution Davidson et al.
-(2018).
-InteL-VAE Miao et al.
-(2022) shows that transformed Gaussian distribution by the invertible function trains hierarchical representation with manual function.
-We show more clear relation of invertibility to disentanglement and improve VAEs to use its unrestricted form of prior.
-Invertible and equivariant Deep Neural Networks have been investigated with normalizing flows.
-As proven by Xiao & Liu (2020), utilized matrix exponential on Neural networks is invertible, but it only provides mathematical foundations of the tra nsformation.
-Matrix exponential is utilized to implement an invertible and equivariant function to improve the generative flow compare to linear function Hoogeboom et al.
-(2020).
-To specify the exponential familyt, other works contribute uncertainty of exponential family distribution with Bayesian update Charpentier et al.
-(2020; 2022b).
-In addition, Ranganath et al.
-(2014) hierarchically controls the natural parameter across the layers and determines the exponential family distribution with the moment of sufficient statistic.
-In our work, we show how to use it for disentanglement learning.
+    Empirical analysis demonstrating the effectiveness of MIPE on standard disentanglement benchmarks.
+2 Preliminaries
+Group: A group is a set $G$ together with binary operation $\ast$, that combines any two elements $g_{a}$ and $g_{b}$ in $G$, such that the following properties:
+  • closure: $g_{a},g_{b}\in G\Rightarrow g_{a}\ast g_{b}\in G$.
+  • Associativity: $\forall g_{a},g_{b},g_{c}\in G$, s.t. $(g_{a}\ast g_{b})\ast g_{c}=g_{a}\ast(g_{b}\ast g_{c})$.
+  • Identity element: There exists an element $e\in G$, s.t. $\forall g\in G,e\ast g=g\ast e=g$.
+  • Inverse element: $\forall g\in G,\exists g^{-1}\in G$: $g\ast g^{-1}=g^{-1}\ast g=e$.
+Group action: Let set $X$, and $(G,\circ)$ be a group, binary operation $\cdot:G\times X\rightarrow X$, then group action $\alpha:\alpha(g,x)=g\cdot x$ following properties:
+  • Identity: $e\cdot x=x$, where $e\in G,\,\,x\in X$.
+  • Compatibility: $\forall g_{1},g_{2}\in G,\,\,x\in X,\,\,\alpha((g_{1}\circ g_{2}),x)=\alpha(g_{1},\alpha(g_{2},x))$.
+Equivariant map: Given $X$ and $Y$ are $G$-set, and group action $\rho:G\times Y\rightarrow Y$. Then a function $f:X\rightarrow Y$ is equivariant if
+$f(\alpha(g,x))=\rho(g,f(x)).$ (1)
+Partial-Equivariant map: Let subset of $G$ be $G^{\prime}\subset G$, then a function $f:X\rightarrow Y$ is a partial-equivariant if
+$f(\alpha(g^{\prime},x))=\rho(g^{\prime},f(x))\,\,\text{where}\,\,g^{\prime}\in G^{\prime}.$ (2)
+Homogeneous space: Let a group $G$ act (on the left) on a smooth manifold $X$ via $\cdot:G\times X\to X$.
+The $G$-space $X$ is called homogeneous if the action is transitive, i.e., for all $x,y\in X$ there exists $g\in G$ with $g\cdot x=y$.
+Exponential Family
+We introduce the conjugate prior and posterior distribution of the exponential family.
+Power Density Function:
+$\begin{split}P{{\bm{x}}}({\bm{x}}|\bm{\theta})&=\text{exp}(\bm{\theta}^{\intercal}T({\bm{x}})-A(\bm{\theta})+B({\bm{x}})),\end{split}$ (3)
+where sufficient statistics $T(\cdot)$, log-normalizer $A(\cdot)$, and carrier or base measure $B(\cdot)$ are known functions, samples $\bm{x}$} from distribution, and natural parameter $\bm{\theta}$.
+If observation is a set of independent identically distributed, then Eq. 3 is modified as
+$p(\mathbf{X}|\bm{\theta})=\prod_{n=1}^{N}h({\bm{x}}_{n})\text{exp}(\bm{\theta}^{\intercal}\sum_{n=1}^{N}T({\bm{x}}_{n})-A(\bm{\theta})),$ (4)
+where observation $\mathbf{X}=\{{\bm{x}}_{1},\cdots{\bm{x}}_{N}\}$ (Bishop, 2006).
+Conjugate Prior: Exponential family admits a conjugate prior:
+$q(\bm{\theta}|\mathcal{\xi},\bm{\nu})=\text{exp}(\bm{\nu}\bm{\theta}^{\intercal}\mathcal{\xi}-\bm{\nu}A(\bm{\theta})+B^{\prime}(\mathcal{\xi},\bm{\nu})),$ (5)
+where $B^{\prime}(\cdot)$ is a normalize coefficient and $\nu$ is evidence, and it is expressed by prior natural parameter $\mathcal{\xi}$.
+Posterior: Let multiply the prior Eq. 5 by the Eq. 4 to obtain the posterior distribution in the form
+$p(\bm{\theta}|\mathbf{X},\mathcal{\xi},\bm{\nu})\propto\text{exp}(\bm{\theta}^{\intercal}(\sum_{n=1}^{N}T({\bm{x}}_{n})+\bm{\nu}\mathcal{\xi})-A(\bm{\theta})),$ (6)
+where the evidence $\bm{\nu}$, sufficient statistics $T(\cdot)$ and log-normalizer, $A(\cdot)$ are known functions, samples $\mathbf{X}$ from distribution, and natural parameter of posterior $\bm{\theta}$ and of prior $\mathcal{\xi}$ (Bishop, 2006).
 Figure 1: The overall architecture of our proposed MIPET-VAE.
-The invertible and partial-equivariant function $\psi(\cdot)$ for L2L transformation consists of a symmetric matrix exponential to be 1) invertible and 2) partial-equivariant.
+The invertible and partial-equivariant function $\psi(\cdot)$ for latent-to-latent (L2L) transformation consists of a symmetric matrix exponential to be 1) invertible and 2) partial-equivariant.
 Then 3) EF conversion module converges the distribution of unrestricted $\hat{{\bm{z}}}$ to be EF with $\mathcal{L}_{el}$ loss.
@@
 In the last, EF conversion reduces the computational error ($\mathcal{L}_{cali}$) between approximated and true KL divergence.
-4) The reddish color represents the integration parts.
-The blue figures represent each property.
-The details of the gray box are in Figure 2.
+3 Related Work
+3.1 Statistical Approaches for Disentangled Representations
+Recently, various studies have focused on unsupervised disentanglement learning.
+$\beta$-VAE (Higgins et al., 2017) penalizes the Kullback-Leibler divergence (KL divergence) using weighted hyper-parameters.
+Factor VAE (Kim and Mnih, 2018) and $\beta$-TCVAE (Chen et al., 2018) are trained using total correlation (TC) to encourage independent dimensions in a latent vector, employing a discriminator and decomposed components of the KL divergence term.
+However, these works are based on the standard Gaussian prior.
+Other VAE approaches implement other prior from Gaussian distribution to transformed Gaussian distribution, Gaussian mixture distribution (Dilokthanakul et al., 2016) or von Mises-Fisher distribution (Davidson et al., 2018).
+Mathieu et al.
+(2018) shows that model regularization with a Gaussian prior is inappropriate to convey sufficient dataset information and for disentanglement learning.
+Rather than using a Gaussian distribution, we employed an exponential family with high flexibility that encompasses a range of different distributions.
+InteL-VAE (Miao et al., 2022) utilize the invertible function to transform a Gaussian distribution to a complex distribution as Mathieu et al.
+(2018).
+Differently, we show a clearer relation of invertibility to disentanglement and improve VAEs to use its unrestricted form of prior.
+Differently, we consider the recent disentanglement definition based on group theory (Higgins et al., 2018).
+3.2 Inductive Bias with Group Theory-Based Approaches
+Following the definitions of disentangled representation learning based on group theory, several works have emphasized equivariant and improved disentangled representation learning.
+Commutative Lie Group VAE (CLG-VAE) (Zhu et al., 2021) proposed a direct mapping of the latent vector into Lie algebra to obtain a group structure (inductive bias) with constraints: commutative and hessian loss.
+Furthermore, Groupified VAE (Yang et al., 2022) utilizes the Spatial Broadcast Decoder (Watters et al., 2019) to implement an equivariant function to the cyclic group while guaranteeing the commutativity and invertibility of group actions.
+Topographic VAE (Keller and Welling, 2021) combines Student’s-t distributions with variational inference and enforces rotated latent vectors to be equivariant.
+On the other hand, we apply an unrestricted prior and posterior for disentanglement learning.
 4 Method
-The overview of a VAE equipped with MIPE-transformation is shown in Figure 1.
-The MIPE-transformation has three main components: 1) IPE-transformation Unit to transform latent vectors with invertible and partial-equivariant properties, 2) EF-conversion Unit to extend VAEs to learn the exponential family distribution of latent vectors, and 3) integrated training and generation process for multiple uses of IPE-transformation and EF-conversion.
-3.1 Invertible and Partial-Equivariant Function for L2L Transformation
-3.1.1 Invertible Property by Using Matrix Exponential
-To guarantee the invertible property of IPE-transformation, we use a function $\psi(\cdot)=\mathbf{e}^{{\bm{M}}}\ast\cdot$ for the transformation, where ${\bm{M}}$ is in $n\times n$ real number matrix set $M_{n}(\mathbb{R})$ Xiao & Liu (2020).
-The operator $\ast$ is matrix multiplication, and $\mathbf{e}^{{\bm{M}}}=\sum_{k}^{\infty}\frac{{\bm{M}}^{k}}{k!}$.
-Our motivation is to use the benefits of injecting explicit inductive bias for disentanglement Locatello et al.
-(2019); Miao et al.
-(2022).
-InteL-VAE effectively extracts hierarchical representation, which includes low-level features (affect to a specific factor) and high-level features (affect to complex factors) with an invertible transformation function Miao et al.
-(2022).
-Figure 2: $G_{I}$, and $G_{L}$ are obtained through encoder $q_{\phi}$ (encoder equivariance condition).
-The left side figure shows the relation between each space and symmetries.
-If $\psi(\cdot)$ is equivariant function over all $G_{L}$, and $G_{T}$, then there exist $\Gamma$, where $\Gamma:G_{L}\rightarrow G_{T}$, and $\Xi\circ\Gamma:G_{I}\rightarrow G_{T}$.
-However, unrestricted $\psi(\cdot)$ has no guarantee to be partial- or full-equivariant.
-The red arrows represent our method: L2L transformation guarantees $\Gamma^{J}:G_{L}^{J}\rightarrow G_{T}^{J}$, and $\Xi^{J}\circ\Gamma^{J}:G_{I}^{J}\rightarrow G_{T}^{J}$, given the encoder equivariance condition $\Xi:G_{I}\rightarrow G_{L}$.
-3.1.2 Why Should L2L Transformation Be Equivariant?
-Let’s consider equivariant function between the input and transformed latent vector space, directly used for a decoder in the VAE frameworks.
-All L2L transformations do not extend the encoder equivariance condition to the relation between input and transformed latent space.
-This problem is more precisely shown in Figure 2, which illustrates partial equivariance condition over the input space $\mathcal{X}$, latent vector space $\mathcal{Z}$, and its transformed latent vector space $\hat{\mathcal{Z}}$ with a corresponding group of symmetries $G_{I}$, $G_{L}$, and $G_{T}$, respectively.
-In the VAEs literature, it has not been reported to restrict L2L transformation to guarantee equivariant function between two spaces, so we propose a solution to guarantee at least a part of symmetries to be equivariant.
-3.1.3 Equivariance Property with Symmetric Matrix Exponential
-To enhance the equivariance of L2L transformation, we set ${\bm{M}}$ of $\psi(\cdot)$ to a symmetric matrix.
-We show that 1) a group with the constraint guarantees equivariance of $\psi(\cdot)$ over the specific group, 2) $\psi(\cdot)$ being equivariant over subset of symmetries between the input space and transformed latent vector space, and 3) the constraint increases the probability of $\psi(\cdot)$ to be in the group (equal to be equivariant over the subset of symmetries).
-We particularly call the transformations as symmetries Higgins et al.
-(2022) to distinguish them from IPE- and I2L-transformations.
-For the generality of our method, we consider an arbitrary VAE model that has no restriction on creating intersections to any set as Figure 2.
-In the next, we show that matrix exponential with symmetric matrix partially preserves encoder equivariance condition better than other matrices.
+The overview of a VAE equipped with MIPE-transformation is shown in Fig.
+1.
+We aim to explain the three core components of the MIPE-transformation.
+In Section 4.1, we mathematically prove why the IPE-transformation, which transforms latent vectors, better preserves partial-equivariance when it is an invertible and symmetric matrix.
+In Section 4.2, we describe EF-conversion, which extends the Gaussian distribution to a diverse exponential family distribution.
+Finally, in Section 4.3, we illustrate how multiple IPE-transformations and EF-conversions can be integrated into a VAE-based model for disentangled representation.
+Notation We denote $g_{I}\in G_{I},g_{L}\in G_{L},\,\text{and}\,g_{T}\in G_{T}$.
+$M_{n}(\mathbb{R})$ is $n\times n$ real matrix, $E_{M}$ is $\{\mathbf{e}^{M}|M\in M_{n}(\mathbb{R})\}$, $E_{S}$ is a set of $\{\mathbf{e}^{S}|S\in Sym_{n}(\mathbb{R})\}$, and $G_{S}$ is a group $G_{S}:(\mathbf{e}^{S},\ast)$, where $\ast$ is a matrix multiplication (equal to commutative).
+Further details are in Appendix A.
+Assumption Let’s assume that 1) encoder $q_{\phi}(\cdot)$ is a partial-equivariant function, 2) $\mathcal{X},\mathcal{Z}^{\prime},\,\text{and}\,\hat{\mathcal{Z}}^{\prime}$ are homogeneous space of group $G_{I},G_{L},\,\text{and}\,G_{T}$, 3) $G_{L}\subset GL_{n}(\mathbb{R})$, where $GL_{n}(\mathbb{R})$ is the general linear group, and 4) matrices are restricted to a compact subset $K\subset M_{n}(\mathbb{R})$ and probabilities are defined with respect to the normalized uniform probability measure $\mu$ on $K$ (a modeling assumption motivated by finite-precision implementations).
+(a) Visualization of homogeneous space (red) as best and worst L2L transformation $\psi(\cdot)$
+(b) Venn diagram of groups
+Figure 2: The homogeneous space $\mathcal{Z}^{\prime}$ is induced by the encoder $q_{\phi}$, and the cardinality of the $\widehat{\mathcal{Z}}^{\,J}$ depends on the latent-to-latent (L2L) transformation.
+4.1 Partial-Equivariant Function and Invertible L2L Transformation
+Why Should L2L Transformation be Equivariant?
+We transform the latent vector $\bm{z}$} sampled from a Gaussian distribution to the transformed latent vector $\hat{{\bm{z}}}$ to be sampled from a different distribution.
+However, not all L2L transformations preserve the partial-equivariance of encoder condition.
+As illustrated in Fig.
+2(a), we choose $\psi$ to maximize the cardinality of the homogeneous $G_{T}^{J}$-space induced in latent space:
+$\psi^{\star}\;=\;\arg\max_{\psi\in M_{n}(\mathbb{R})}\;\big|\widehat{\mathcal{Z}}^{\,J}(\psi)\big|,\quad$ (7)
+where $G_{T}^{J}$ act on the latent space via $\alpha$ and define $\widehat{\mathcal{Z}}^{\,J}(\psi):=\{\alpha(g,\hat{{\bm{z}}})\mid g\in G_{T}^{J}\}$.
+As a homogeneous $G_{T}^{J}$–space, fixing a base point $\hat{{\bm{z}}}_{0}$ determines all elements $\alpha(g,\hat{{\bm{z}}}_{0})$.
+If the latent space factors as $\hat{\mathcal{Z}}=\hat{\mathcal{Z}}_{1}\times\cdots\times\hat{\mathcal{Z}}_{k}$ and the action is componentwise, then every $\alpha(g,\hat{{\bm{z}}}_{0})$ with $g\in G_{1}\times\cdots\times G_{k}$ is also disentangled (Higgins et al., 2018).
+Equivariance Property with Symmetric Matrix Exponential
+We prove that an invertible and symmetric matrix preserves partial-equivariance better than other matrices for Eq.
+7.
+We show that 1) a group with symmetric and invertible matrices guarantees equivariance of $\psi(\cdot)$ over the specific group $G_{S}$, 2) this $\psi(\cdot)$ being equivariant over subset of symmetries between the input space and transformed latent vector space (equal to be a partial-equivariant), and 3) the invertible and symmetric matrix (symmetric matrix exponential) increases the probability of $\psi(\cdot)$ to be in the group (equal to satisfy Eq.
+7).
+For the generality of our method, we consider an arbitrary VAE model that has no restriction on creating intersections to any set as Fig. 2(b).
 Proposition 4.1.
@@
 The group $G_{S}$ is closed to matrix multiplication, and its element is always a symmetric matrix by definition.
-Then, any two elements in $G_{S}$ are commutative because if matrix multiplication of two symmetric matrices is symmetric then both are commutative.
+Then, any two elements in $G_{S}$ are commutative because if matrix multiplication of two symmetric matrices is symmetric then both are commutative by Lemma B.1.
 As a result, $\psi_{G_{S}}(\cdot)$ and group elements of $G_{S}$ are commutative ($G_{S}$ is an abelian group).
 Because of the commutativity, $\psi_{G_{S}}(g_{s}\circ{\bm{z}})=\mathbf{e}^{{\bm{S}}}g_{s}{\bm{z}}=g_{s}\mathbf{e}^{{\bm{S}}}{\bm{z}}=g_{s}\circ\psi_{G_{S}}({\bm{z}})$ for $g_{s}\in G_{S}$ if the group action $\circ$ is set to matrix multiplication, where $\psi_{G_{S}}\in G_{S}$.
-This equation satisfies the general definition of an equivariant function that a function $f(\cdot)$ is equivariant if $f(g\circ{\bm{z}})=g\circ f({\bm{z}})$ for all $g$ in a group $G$ by matching $f$, $g$, and $G$ to $\psi_{G_{S}}$, $g_{s}$, and $G_{S}$, respectively.
+This equation satisfies the general definition of an equivariant function Eq.
+1 by matching $f$, $g$, and $G$ to $\psi_{G_{S}}$, $g_{s}$, and $G_{S}$, respectively.
 $\blacksquare$
 Proposition 4.2.
-If $q_{\phi}$ is equivariant over defined on group of symmetries $G_{I}^{a}$ and $G_{L}^{a}$, then $\psi_{G_{S}}(q_{\phi}(\cdot))$ is equivariant to symmetries in $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$ by the equivariance of $q_{\phi}$.
+If $q_{\phi}$ is equivariant over defined on group of symmetries $G_{I}^{J}$ and $G_{L}^{J}$, then $\psi_{G_{S}}(q_{\phi}(\cdot))$ is equivariant to symmetries in $G_{I}$ corresponding to $G_{S}\cap G_{L}$ ($G_{I}^{J}$) and $G_{T}$ corresponding to $G_{S}\cap G_{L}$ ($G_{T}^{J}$) by the equivariance of $q_{\phi}$.
 Proof.
 The function $\psi_{G_{S}}(\cdot)$ is an equivariant function over group elements in $G_{S}\cap G_{L}$ by Proposition 4.1.
-Then, the composite function, $\psi_{G_{S}}(\cdot)$ and $q_{\phi}$, is an equivariant function of $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$.
-Let $g_{L}^{a}$ be a group element in $G_{S}\cap G_{L}$, and $g_{I}^{a}$ is a group element in $G_{I}$ corresponding to $G_{S}\cap G_{L}$.
-More details are in Appendix C.1.
+Then, the composite function, $\psi_{G_{S}}(\cdot)$ and $q_{\phi}$, is an equivariant function of $G_{I}^{J}$ and $G_{T}^{J}$.
+Let $g_{L}^{J}\in G_{L}^{J}$, and $g_{T}^{J}\in G_{T}^{J}$.
+Then, $\psi_{G_{S}}(g_{L}^{J}{\bm{z}}_{1})=g_{L}^{J}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{J}\psi_{G_{S}}({\bm{z}}_{1})$.
 $\blacksquare$
-Figure 3: Equivariant map: $\mathcal{X}$, $\mathcal{Z}$, and $\hat{\mathcal{Z}}$ are input space, latent vector space, and transformed latent vector space by L2L transfomration function $\psi(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
-respectively.
-${\bm{x}}\in\mathcal{X}$, ${\bm{z}}\in\mathcal{Z}$, and $\hat{{\bm{z}}}\in\hat{\mathcal{Z}}$.
-Then,
+Therefore, the invertible and symmetric matrix is a partial-equivariant function between the data space and transformed latent vector space. More details are in Proposition B.2.
+Let $P(G):=\frac{\mu(G\cap K)}{\mu(K)}$ by the assumption 4 in section 4, and $Pr(\psi_{G^{\prime}}(\cdot)\in G)=P(\psi(\cdot)\in G|\psi(\cdot)\in G^{\prime})$ after VAE training.
 Proposition 4.3.
@@
 Therefore $E_{M}\not\subset E_{S}$.
-Therefore, the probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{S})}$ is greater than $Pr(\psi_{E_{M}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{M})}$.
+Therefore, the conditional probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{S})}$ is greater than $Pr(\psi_{E_{M}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{M})}$.
 In the same way, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(M_{n}(\mathbb{R}))}$ because $E_{M}\subset M_{n}(\mathbb{R})$ and non-invertible functions are only in $M_{n}(\mathbb{R})$.
+More details are in Proposition B.3.
 $\blacksquare$
-Therefore, $\psi_{E_{S}}$ clearly increases the probability of preserving a certain type of equivariance compared to unrestricted $\psi$ functions.
-The conditional probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})$, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})$, and $Pr(\psi_{M}(\cdot)\in G_{S})$ is changed by the distribution of the observation of $\psi(\cdot)$, which depends on the model parameters.
+Therefore, $\psi_{E_{S}}$ clearly increases the probability of Eq.
+7, because the group $G_{L}$ is obtained ($|\hat{\mathcal{Z}}^{J}(\psi)|$ is fixed) by our assumption that $q_{\phi}(\cdot)$ is a partial-equivariant.
+It implies that symmetric and invertible properties for a matrix are necessary to preserve partial equivariance.
+The probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})$, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})$, and $Pr(\psi_{M}(\cdot)\in G_{S})$ is changed by the distribution of the observation of $\psi(\cdot)$, which depends on the model parameters.
 However, the inequality $Pr(\psi_{E_{S}}(\cdot)\in G_{S})>Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})$ is not changed regardless of the distribution of observation of $\psi(\cdot)$.
 We empirically validate the impact of equivariance with the uncertain $P(\cdot)$ to disentanglement in Section D.5.
-3.1.4 Relation Between $\psi(\cdot)$ and Disentanglement
-In addition to invertible and partial-equivariant properties, our IPE-transformation also guarantees zero Hessian matrix, which enhances disentanglement without any additional loss of Peebles et al.
-(2020).
-Hessian matrix of the transformation $\nabla_{\bm{z}}^{2}\psi({\bm{z}})=\nabla_{\bm{z}}(\nabla_{\bm{z}}\mathbf{e}^{{\bm{M}}}{\bm{z}})=0$ because of the irrelevance of ${\bm{M}}$ to ${\bm{z}}$.
-By this property, $\psi(\cdot)$ leads that independently factorizes each dimension Peebles et al.
-(2020), and it injects group theory based inductive bias simultaneously.
-This is because the group decomposition of ${\bm{z}}$ space $G=G_{1}\times G_{2}\times\cdots\times G_{k}$ corresponds to group decomposition of the transformed latent vector $\hat{{\bm{z}}}$ space $G^{\prime}=G_{1}^{\prime}\times G_{2}^{\prime}\times\cdots\times G_{k}^{\prime}$ such that each $G_{i}^{\prime}$ is fixed by the action of all the $G_{j}$ for $j\neq i$ Yang et al.
-(2022); Higgins et al.
-(2018b).
-This correspondence of decomposition is expected to transfer the independence between dimensions of ${\bm{z}}$ to the space of $\hat{{\bm{z}}}$ Higgins et al.
-(2018a).
+Invertible Property by Using Matrix Exponential
+To guarantee the invertible property of IPE-transformation, we use a function $\psi(\cdot)=\mathbf{e}^{{\bm{M}}}\ast\cdot$ for the transformation, where $\bm{M}$} is in $n\times n$ real number matrix set $M_{n}(\mathbb{R})$ (Xiao and Liu, 2020).
+The operator $\ast$ is matrix multiplication, and $\mathbf{e}^{{\bm{M}}}=\sum_{k}^{\infty}\frac{{\bm{M}}^{k}}{k!}$.
+Our motivation is to use the benefits of injecting explicit inductive bias for disentanglement (Locatello et al., 2019; Miao et al., 2022).
+InteL-VAE effectively extracts hierarchical representation, which includes low-level features (affect to a specific factor) and high-level features (affect to complex factors) with an invertible transformation function (Miao et al., 2022).
 4.2 Exponential Family Conversion for Unknown Prior
 In VAE frameworks, the Gaussian normal distribution is applied as a prior.
-However, a prior from data is usually unknown and may not follow the Gaussian distribution Miao et al.
-(2022).
+However, a prior from data is usually unknown and may not follow the Gaussian distribution (Miao et al., 2022).
 As a solution, we present a training procedure for VAEs to build an exponential family distribution from a latent variable of an arbitrary distribution.
 Then, we introduce training losses obtained from the unit IPE-transformation function and EF-conversion.
-3.2.1 Elements of Exponential Family Distribution Settings
-First, we set sufficient statistics $T(\cdot)$, log-normalizer $A(\cdot)$, and carrier or base measure $B(\cdot)$ are deterministic functions by maximizing conjugate prior for parameter $\mathcal{\xi}$.
-To determine the natural parameter of posterior and prior $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$, and $\hat{\bm{\epsilon}}_{m}$, we use a natural parameter generator (NPG) designed by multi-layer perceptron Charpentier et al.
-(2022a).
+Elements of Exponential Family Distribution Settings
 As introduced in Bishop (2006); Charpentier et al.
-(2022a), we assume exponential family always admits a conjugate prior:
-$q(\bm{\theta}|\mathcal{\xi},\bm{\nu})=\text{exp}(\bm{\nu}\bm{\theta}^{\intercal}\mathcal{\xi}-\bm{\nu}A(\bm{\theta})+B^{\prime}(\mathcal{\xi},\bm{\nu})),$ (1)
-where $B^{\prime}(\cdot)$ is a normalize coefficient and $\nu$ is evidence, and it is expressed by prior natural parameter $\mathcal{\xi}$.
+(2022), we assume exponential family always admits a conjugate prior Eq.
+5.
+To determine the natural parameter of posterior and prior $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$, and $\hat{\bm{\epsilon}}_{m}$, we use a natural parameter generator (NPG) designed by multi-layer perceptron (Charpentier et al., 2022).
 However, generated natural parameter $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$ is not guaranteed as the appropriate parameter of the exponential family corresponds to conjugate prior.
-To satisfy this condition, we assume observation is a set of independent identically distributed, then Eq.
-13 is modified: $p(\mathbf{X}|\bm{\theta})=\prod_{n=1}^{N}h(\mathbf{x}_{n})\text{exp}(\bm{\theta}^{\intercal}\sum_{n=1}^{N}T(\mathbf{x}_{n})-A(\bm{\theta}))$ Bishop (2006), where observation $\mathbf{X}=\{\mathbf{x}_{1},\cdots\mathbf{x}_{N}\}$.
-In the next, we multiply the modified formation by the prior Eq.
-1 to obtain the posterior distribution Bishop (2006) as Eq.
-2.
-3.2.2 Distribution Approximation As an Exponential Family
-The procedure represents a posterior distribution in the exponential family by adopting the following form:
-$p(\bm{\theta}|\mathbf{X},\mathcal{\xi},\bm{\nu})\propto\text{exp}(\bm{\theta}^{\intercal}(\sum_{n=1}^{N}T(\mathbf{x}_{n})+\bm{\nu}\mathcal{\xi})-A(\bm{\theta})),$ (2)
-where sufficient statistics $T(\cdot)$ and log-normalizer, $A(\cdot)$ are known functions, samples $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\}$ from distribution, and natural parameter of posterior $\bm{\theta}$ and of prior $\mathcal{\xi}$ Bishop (2006).
-The functions $T(\cdot)$, and $A(\cdot)$ are deterministic functions to maximize posterior distribution.
-The evidence is implemented as learnable parameters $\bm{\nu}\in\mathbb{R}^{n\times n}$.
-The natural parameter is generated by a multi-layer perceptron as Charpentier et al.
-(2022a).
+To satisfy this condition, we assume observation is a set of independent identically distributed, then we follow Eq.
+4.
+Distribution Approximation As an Exponential Family
+By our exponential family settings, a posterior distribution in the exponential family by adopting the following Eq.
+6.
+The evidence is implemented as learnable parameters $\bm{\nu}\in\mathbb{R}^{n\times n}$, and natural parameter is generated by a multi-layer perceptron as Charpentier et al.
+(2022).
 This general form approximating an exponential family distribution with learnable parameters can extend VAEs to use a wider distribution for latent variables by simply matching $\mathbf{X}$ to generated latent variables.
-After IPE-transformation, we can apply the form by using the $\hat{{\bm{z}}}_{m}$, $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$, and $\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}$ for $\mathbf{X}$, $\bm{\theta}$, and $\mathcal{\xi}$, respectively.
-3.2.3 EF Similarity Loss
-We added a loss to converge the unrestricted distributions of $\hat{{\bm{z}}}$ to the power density function of the exponential family by constraining the posterior maximization as:
-$\begin{split}\text{maximize}&~{}\log p(\bm{\theta}_{\hat{{\bm{z}}}_{m}}|\hat{{\bm{z}}}_{m},\bm{\theta}_{\hat{\bm{\epsilon}}_{m}},\bm{\nu}_{m})~{}~{}~{}\text{s.t.}~{}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\geq 0\end{split}$ (3)
-$\begin{split}\Rightarrow\mathcal{L}_{s}(\hat{{\bm{z}}}_{m},\hat{\bm{\epsilon}}_{m})&=\log p(\bm{\theta}_{\hat{{\bm{z}}}_{m}}|\hat{{\bm{z}}}_{m},\bm{\theta}_{\hat{\bm{\epsilon}}_{m}},\bm{\nu}_{m})+\bm{\lambda}_{m}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{\mathbf{x}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\end{split}$ (4)
-$\begin{split}\Rightarrow\mathcal{L}_{el}\coloneqq||\nabla_{\hat{{\bm{z}}}_{m},\hat{\bm{\epsilon}}_{m},\bm{\lambda}_{m}}\mathcal{L}_{s}||_{2}^{2}=0.\end{split}$ (5)
+After IPE-transformation, we can apply the form by using the $\hat{{\bm{z}}}_{m}$, $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$, and $\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}$ for $\mathbf{X}$, $\bm{\theta}$, and $\mathcal{\xi}$, respectively as shown in Fig.
+1.
+EF Similarity Loss
+To maximize Eq. 6 to be an exponential family, we added a loss through a Lagrange multiplier as:
+$\begin{aligned} \text{maximize}&~\log p(\bm{\theta}_{\hat{{\bm{z}}}_{m}}|\hat{{\bm{z}}}_{m},\bm{\theta}_{\hat{\bm{\epsilon}}_{m}},\bm{\nu}_{m})~~~\text{s.t.}~D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\geq 0\end{aligned}$
+$\begin{split}\Rightarrow\mathcal{L}_{s}(\hat{{\bm{z}}}_{m},\hat{\bm{\epsilon}}_{m})&=\log p(\bm{\theta}_{\hat{{\bm{z}}}_{m}}|\hat{{\bm{z}}}_{m},\bm{\theta}_{\hat{\bm{\epsilon}}_{m}},\bm{\nu}_{m})\\ &+\bm{\lambda}_{m}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{\mathbf{x}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\end{split}$ (9)
+[…]
 The notation $\bm{\theta}_{k}$ is a generated natural parameter by a given $k\in\{\hat{{\bm{z}}},\hat{\bm{\theta}}\}$, and $f_{{\bm{x}}}({\bm{x}}|\bm{\theta})$ is a power density function of the exponential family.
 Moreover, $\bm{\lambda}_{m}$ is a trainable parameter for optimizing the Lagrange multiplier, and $D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))$ is a KL divergence of the exponential family.
-3.2.4 KL Divergence for Evidence of Lower Bound
-The KL divergence of Gaussian distribution Kingma & Welling (2013) is computed using mean and variance, which are the parameters of a Gaussian distribution.
-To introduce a loss as the KL divergence of Gaussian distribution, we compute KL divergence of the exponential family in Eq.
+KL Divergence for Evidence of Lower Bound
+We compute the KL divergence of the exponential family in Eq.
 9 using the learnable parameter $T(\cdot)$ and $A(\cdot)$ with given natural parameter $\bm{\theta}_{\hat{{\bm{z}}}}$ and $\bm{\theta}_{\hat{\bm{\epsilon}}}$, expressed as:
-$\begin{split}\mathcal{L}_{kl}&\coloneqq D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\\ &=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{{\bm{z}}}}}A(\bm{\theta}_{\hat{{\bm{z}}}})-\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{\bm{\epsilon}}}}A(\bm{\theta}_{\hat{\bm{\epsilon}}}).\end{split}$ (6)
-3.2.5 KL Divergence Calibration Loss
-To reduce the error between the approximation and true matrix for the matrix exponential Bader et al.
-(2019), we add a loss to minimize the difference of their KL divergence measured by mean squared error (MSE) as:
-$\begin{split}\mathcal{L}_{cali}=\text{MSE}(&D_{\mathrm{KL}}(q_{\phi}({\bm{z}}|{\bm{x}})||p_{\theta}({\bm{z}})),D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))),\end{split}$ (7)
+[…]
+Further details are in the Appendix B.1.
+KL Divergence Calibration Loss
+Although $\mathrm{KL}$ is invariant under a common invertible map $\psi$ (Miao et al., 2022), practical approximations (e.g., matrix-exponential) can break this.
+We therefore add a consistency regularizer that penalizes the gap between empirical KLs before and after applying $\psi$:
+$\begin{aligned} \mathcal{L}_{\text{cali}}=\text{MSE}(D_{\mathrm{KL}}(q_{\phi}({\bm{z}}|{\bm{x}})\,||\,p_{\theta}({\bm{z}})),D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})\,||\,f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}})))\end{aligned}$
 which is the KL divergence calibration loss ($\mathcal{L}_{cali}$).
-3.2.6 Implicit Semantic Mask
+Table 1: Performance (mean $\pm$ std) of four metrics on dSprites, 3D Shapes, and 3D Cars.
+dSprites FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
+original MIPET original MIPET original MIPET original MIPET
+$\beta$-VAE 69.15($\pm$5.88) 74.19($\pm$5.62) 9.49($\pm$8.30) 19.72($\pm$11.37) 2.43($\pm$2.07) 5.08($\pm$2.90) 18.57($\pm$12.41) 28.81($\pm$10.19)
+$\beta$-TCVAE 78.50($\pm$7.93) 79.87($\pm$5.80) 26.00($\pm$9.06) 35.04($\pm$4.07) 7.31($\pm$0.61) 7.70($\pm$1.63) 41.80($\pm$8.55) 47.83($\pm$5.01)
+Factor-VAE 67.78($\pm$7.48) 68.38($\pm$8.55) 14.67($\pm$10.40) 19.31($\pm$13.46) 2.35($\pm$2.32) 4.24($\pm$3.36) 22.58($\pm$8.50) 30.18($\pm$12.99)
+CLG-VAE 79.06($\pm$6.83) 81.80($\pm$3.17) 23.40($\pm$7.89) 36.34($\pm$5.55) 7.37($\pm$0.96) 8.03($\pm$0.83) 37.68($\pm$7.83) 44.73($\pm$5.11)
+Control-VAE 62.36($\pm$8.62) 67.71($\pm$6.41) 4.36($\pm$2.86) 7.34($\pm$4.10) 2.11($\pm$1.88) 1.93($\pm$1.63) 10.40($\pm$3.42) 15.18($\pm$4.61)
+3D Shapes FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
+original MIPET original MIPET original MIPET original MIPET
+$\beta$-VAE 71.76($\pm$12.26) 75.19($\pm$8.16) 37.33($\pm$22.34) 47.37($\pm$10.13) 7.48($\pm$4.12) 9.20($\pm$2.44) 52.07($\pm$17.92) 54.95($\pm$8.99)
+$\beta$-TCVAE 76.62($\pm$10.23) 80.59($\pm$8.57) 52.93($\pm$20.5) 54.49($\pm$9.44) 10.64($\pm$5.93) 11.58($\pm$3.32) 65.32($\pm$11.37) 66.22($\pm$7.32)
+Factor-VAE 77.43($\pm$10.71) 78.34($\pm$8.35) 24.23($\pm$26.13) 48.59($\pm$10.58) 4.99($\pm$4.46) 9.84($\pm$2.73) 68.68($\pm$12.92) 60.23($\pm$9.90)
+CLG-VAE 77.04($\pm$8.22) 80.17($\pm$8.43) 49.74($\pm$8.18) 53.87($\pm$7.41) 9.20($\pm$2.44) 12.83($\pm$3.01) 57.70($\pm$8.60) 60.74($\pm$7.77)
+Control-VAE 71.05($\pm$14.35) 71.89($\pm$8.33) 24.88($\pm$13.68) 32.28($\pm$10.74) 6.60($\pm$3.59) 7.14($\pm$2.09) 40.08($\pm$13.45) 43.06($\pm$8.68)
+3D Cars FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
+original MIPET original MIPET original MIPET original MIPET
+$\beta$-VAE 89.48($\pm$5.22) 88.95($\pm$5.94) 6.90($\pm$2.70) 7.27($\pm$1.99) 1.30($\pm$0.48) 1.88($\pm$1.12) 19.85($\pm$4.87) 18.90($\pm$4.49)
+$\beta$-TCVAE 95.84($\pm$3.40) 96.43($\pm$2.42) 11.87($\pm$2.90) 10.80($\pm$1.22) 1.55($\pm$0.38) 1.88($\pm$1.12) 27.91($\pm$4.31) 26.08($\pm$2.47)
+Factor-VAE 89.64($\pm$2.62) 93.66($\pm$3.83) 10.63($\pm$1.30) 9.43($\pm$1.12) 2.51($\pm$0.37) 3.28($\pm$0.83) 26.27($\pm$3.17) 23.13($\pm$0.72)
+CLG-VAE 86.11($\pm$7.12) 91.06($\pm$5.09) 6.19($\pm$2.42) 8.51($\pm$2.11) 2.06($\pm$0.60) 1.99($\pm$0.93) 16.91($\pm$4.01) 18.31($\pm$2.83)
+Control-VAE 88.76($\pm$7.66) 89.10($\pm$6.90) 4.68($\pm$2.67) 5.08($\pm$2.68) 1.16($\pm$0.74) 1.45($\pm$0.86) 14.70($\pm$3.84) 15.22($\pm$4.15)
+Implicit Semantic Mask
 We propose an implicit semantic mask to improve disentanglement learning.
@@
 where $\mathcal{W}$ is the weight of log-normalizer, $\lambda$ is a hyper-parameter, $\mathbf{\mu}_{|\mathcal{W}_{ij}|}$, and $\mathbf{\sigma}_{|\mathcal{W}_{ij}|}$ are the mean, and standard deviation of weight respectively.
-Previous work Yang et al.
-(2020) utilizes a semantic mask in input space directly, but we inject the semantic mask implicitly on the latent space.
+Previous work (Yang et al., 2020) utilizes a semantic mask in input space directly, but we inject the semantic mask implicitly on the latent space.
 4.3 Integration for Multiple IPE-Transformation and EF-Conversion
@@
 Each IPE-transformation function operates independently, then the reconstruction error for objective function is defined as:
-$\begin{split}\mathcal{L}_{rec}&\coloneqq\frac{1}{k}\sum_{i=1}^{k}\Bigg{[}\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg{]}=\frac{1}{k}\sum_{i=1}^{k}E_{q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}})),\end{split}$ (9)
+$\begin{aligned} \mathcal{L}_{\text{rec}}&=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}}))\end{aligned}$
 where $\hat{{\bm{z}}}_{i}=\psi_{i}({\bm{z}})$. Therefore, we define ELBO as:
-$\begin{split}\mathcal{L}^{\prime}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})&=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi,\psi_{i}}({\bm{z}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}}))-\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})||p_{\psi_{i}}({\bm{z}})).\end{split}$ (10)
+$\begin{split}\mathcal{L}^{\prime}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})&=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi,\psi_{i}}({\bm{z}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}}))\\ &-\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})||p_{\psi_{i}}({\bm{z}})).\end{split}$ (15)
 However, following Eq.
-10, k samples are generated, and each sample is disentangled for different factors.
+15, k samples are generated, and each sample is disentangled for different factors.
 We implement the output as the average of the sum of the k samples to obtain a single sample with a superposition effect from k samples.
 Moreover, the KL divergence term in Eq.
-10 represents that increasing number of MIPE-transformation is equal to increasing $\beta$ hyper-parameter in $\beta$-VAE Higgins et al.
-(2017).
+15 represents that increasing number of MIPE-transformation is equal to an increasing $\beta$ hyper-parameter in $\beta$-VAE (Higgins et al., 2017) and more details are in Appendix C.
 The VAEs equipped with MIPE-transformation (MIPET-VAEs) can be trained with the following loss:
 $\begin{split}&\mathcal{L}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})=\mathcal{L}_{rec}-\mathcal{L}_{kl}-\mathcal{L}_{el}-\mathcal{L}_{cali}.\end{split}$ (16)
-Table 2: Performance (mean $\pm$ std) of four metrics on dSprites, 3D Shapes, and 3D Cars. The $\alpha=1$ and $\gamma=1$ of $\beta$-TCVAE as Chen et al. (2018).
-dSprites Disentanglement Metric
-FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
-original MIPET original MIPET original MIPET original MIPET
-$\beta$-VAE 69.15($\pm$5.88) 74.19($\pm$5.62) 9.49($\pm$8.30) 19.72($\pm$11.37) 2.43($\pm$2.07) 5.08($\pm$2.90) 18.57($\pm$12.41) 28.81($\pm$10.19)
-$\beta$-TCVAE 78.50($\pm$7.93) 79.87($\pm$5.80) 26.00($\pm$9.06) 35.04($\pm$4.07) 7.31($\pm$0.61) 7.70($\pm$1.63) 41.80($\pm$8.55) 47.83($\pm$5.01)
-CLG-VAE 79.06($\pm$6.83) 81.80($\pm$3.17) 23.40($\pm$7.89) 36.34($\pm$5.55) 7.37($\pm$0.96) 8.03($\pm$0.83) 37.68($\pm$7.83) 44.73($\pm$5.11)
-Control-VAE 62.36($\pm$8.62) 67.71($\pm$6.41) 4.36($\pm$2.86) 7.34($\pm$4.10) 2.11($\pm$1.88) 1.93($\pm$1.63) 10.40($\pm$3.42) 15.18($\pm$4.61)
-3D Shapes Disentanglement Metric
-FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
-original MIPET original MIPET original MIPET original MIPET
-$\beta$-VAE 71.76($\pm$12.26) 75.19($\pm$8.16) 37.33($\pm$22.34) 47.37($\pm$10.13) 7.48($\pm$4.12) 9.20($\pm$2.44) 52.07($\pm$17.92) 54.95($\pm$8.99)
-$\beta$-TCVAE 76.62($\pm$10.23) 80.59($\pm$8.57) 52.93($\pm$20.5) 54.49($\pm$9.44) 10.64($\pm$5.93) 11.58($\pm$3.32) 65.32($\pm$11.37) 66.22($\pm$7.32)
-CLG-VAE 77.04($\pm$8.22) 80.17($\pm$8.43) 49.74($\pm$8.18) 53.87($\pm$7.41) 9.20($\pm$2.44) 12.83($\pm$3.01) 57.70($\pm$8.60) 60.74($\pm$7.77)
-Control-VAE 71.05($\pm$14.35) 71.89($\pm$8.33) 24.88($\pm$13.68) 32.28($\pm$10.74) 6.60($\pm$3.59) 7.14($\pm$2.09) 40.08($\pm$13.45) 43.06($\pm$8.68)
-3D Cars Disentanglement Metric
-FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
-original MIPET original MIPET original MIPET original MIPET
-$\beta$-VAE 89.48($\pm$5.22) 88.95($\pm$5.94) 6.90($\pm$2.70) 7.27($\pm$1.99) 1.30($\pm$0.48) 1.88($\pm$1.12) 19.85($\pm$4.87) 18.90($\pm$4.49)
-$\beta$-TCVAE 95.84($\pm$3.40) 96.43($\pm$2.42) 11.87($\pm$2.90) 10.80($\pm$1.22) 1.55($\pm$0.38) 1.88($\pm$1.12) 27.91($\pm$4.31) 26.08($\pm$2.47)
-CLG-VAE 86.11($\pm$7.12) 91.06($\pm$5.09) 6.19($\pm$2.42) 8.51($\pm$2.11) 2.06($\pm$0.60) 1.99($\pm$0.93) 16.91($\pm$4.01) 18.31($\pm$2.83)
-Control-VAE 88.76($\pm$7.66) 89.10($\pm$6.90) 4.68($\pm$2.67) 5.08($\pm$2.68) 1.16($\pm$0.74) 1.45($\pm$0.86) 14.70($\pm$3.84) 15.22($\pm$4.15)
-Table 3: p-value of t-test for original vs MIPET results of Table 2, which are averaged over models (bold: positive and significant, italic: positive but insignificant, normal: lower performance).
-p-value VAEs CLG-VAE $\beta$-TCVAEs
-FVM MIG SAP DCI FVM MIG SAP DCI FVM MIG SAP DCI
-dSprites 0.000 0.000 0.000 0.000 0.030 0.000 0.005 0.000 0.281 0.000 0.170 0.009
-3D Shapes 0.080 0.007 0.016 0.191 0.085 0.029 0.000 0.088 0.111 0.383 0.277 0.390
-3D Cars 0.659 0.250 0.003 0.583 0.003 0.000 0.630 0.071 0.278 0.923 0.119 0.933
+Table 2: Impact of the number of MIPE-transformation function on the $\beta$-TCVAE and $\beta$-VAE with dSprites, 3D Shapes, and 3D Cars datasets in terms of the four metrics.
+The blue and red box plots represent each model’s single and multiple IPE-transformation cases, respectively (A-$n$: MIPET-$\beta$-TCVAE (4), B-$n$: MIPET-$\beta$-TCVAE (6), C-$n$: MIPET-$\beta$-VAE, $n$: the number of MIPE-transformation).
+Further details are in Table 9.
+dSprites 3D Shapes
+FVM MIG SAP DCI FVM MIG SAP DCI
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+[…]
+ratio $\beta$-VAE (1) CLG-VAE (0.5)
+$\lambda$ FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$ FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
+0.0 90.46($\pm$6.50) 4.84($\pm$2.32) 1.29($\pm$0.81) 16.76($\pm$4.68) 90.06($\pm$4.44) 9.28($\pm$2.09) 1.82($\pm$0.82) 19.12($\pm$3.41)
+0.5 91.35($\pm$5.52) 5.37($\pm$2.74) 1.17($\pm$0.67) 16.65($\pm$3.76) 88.69($\pm$4.78) 6.90($\pm$1.96) 1.85($\pm$0.67) 17.52($\pm$3.16)
+1.0 91.78($\pm$6.20) 4.99($\pm$2.27) 1.36($\pm$0.81) 16.50($\pm$2.53) 83.60($\pm$11.48) 8.12($\pm$3.66) 2.37($\pm$1.50) 17.07($\pm$3.89)
+1.5 90.04($\pm$5.88) 7.22($\pm$2.87) 1.36($\pm$0.48) 18.23($\pm$2.84) 84.76($\pm$6.86) 7.70($\pm$2.11) 2.05($\pm$0.73) 17.06($\pm$2.77)
+2.0 87.79($\pm$8.88) 4.75($\pm$2.49) 1.01($\pm$0.99) 16.64($\pm$3.75) 85.78($\pm$4.18) 7.83($\pm$1.79) 1.91($\pm$0.96) 17.26($\pm$2.07)
+$\infty$ 89.43($\pm$11.72) 3.74($\pm$2.32) 0.77($\pm$0.39) 15.45($\pm$4.59) 82.96($\pm$11.84) 8.07($\pm$2.52) 2.32($\pm$1.02) 17.46($\pm$4.07)
 5 Experiment Settings
-4.0.1 Models
-As baseline models, we select VAE, $\beta$-VAE, $\beta$-TCVAE, and CLG-VAE.
+5.1 Models
+As baseline models, we select VAE (Kingma and Welling, 2013), $\beta$-VAE (Higgins et al., 2017), $\beta$-TCVAE (Chen et al., 2018), Factor-VAE (Kim and Mnih, 2018), Control-VAE (Shao et al., 2020),and CLG-VAE (Zhu et al., 2021).
 These models are compared to their extension to adopt MIPET, abbreviated by adding the MIPET prefix.
+Also, we follow the same model architecture which are introduced in previous works (Kim and Mnih, 2018) and model details are in Table 7-7.
+5.2 Datasets
+As in prior works (Higgins et al., 2017; Chen et al., 2018; Kim and Mnih, 2018; Zhu et al., 2021; Yang et al., 2022; Keller and Welling, 2021), we primarily evaluate our models on synthetic datasets, where learning disentangled representations remains a well-known challenge (Matthey et al., 2017; Reed et al., 2015; Burgess and Kim, 2018).
+We benchmark MIPET-VAEs against several standard VAEs on the dSprites (Matthey et al., 2017), 3D Shapes (Burgess and Kim, 2018), and 3D Cars (Reed et al., 2015) datasets.
+Further details are in the Appendix D.2
+5.3 Training
+We set 256 mini-batch size in the datasets (dSprites, 3D Shapes, and 3D Cars), Adam optimizer with learning rate $4\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and epochs from $\{30,67,200\}$ as a common setting for all the comparative methods.
+In addition, we set masking ratio $\lambda$ from $\{0.0,0.5,\ldots 2.0,\infty\}$.
+We then evaluate familiar benchmarks, FVM score (Kim and Mnih, 2018), MIG (Chen et al., 2018), SAP (Kumar et al., 2018), and DCI (Eastwood and Williams, 2018).
+Further details are in the Appendix D.3.
+Table 4: Ablation study for the equivariant property (w/o E) and EF-conversion (w/o EF). Each metric is averaged over 40 and 20 settings of $\beta$-VAE and $\beta$-TCVAE, respectively.
+3D Shapes $\beta$-VAE $\beta$-TCVAE Factor-VAE CLG-VAE
+MIPET w/o E w/o EF MIPET w/o E w/o EF MIPET w/o E w/o EF MIPET w/o E w/o EF
+FVM 75.19($\pm$8.16) 74.91($\pm$10.46) 22.27($\pm$1.29) 80.59($\pm$8.57) 77.90($\pm$8.66) 66.38($\pm$7.57) 78.34($\pm$8.35) 71.50($\pm$6.78) 59.28($\pm$5.15) 80.17($\pm$8.43) 69.75($\pm$6.91) 61.55($\pm$9.05)
+MIG 47.37($\pm$10.13) 47.45($\pm$8.98) 0.28($\pm$0.09) 54.49($\pm$9.44) 51.37($\pm$11.54) 36.08($\pm$17.42) 48.59($\pm$8.35) 42.25($\pm$11.28) 32.67($\pm$10.49) 53.87($\pm$7.41) 45.92($\pm$4.77) 33.89($\pm$3.46)
+SAP 9.20($\pm$2.44) 9.43($\pm$2.59) 0.26($\pm$0.07) 11.58($\pm$3.32) 10.23($\pm$3.13) 7.13($\pm$3.09) 9.84($\pm$2.73) 8.19($\pm$0.69) 7.04($\pm$0.87) 12.83($\pm$3.01) 50.14($\pm$5.43) 46.66($\pm$4.81)
+DCI 54.95($\pm$8.99) 54.23($\pm$9.05) 0.10($\pm$0.02) 66.22($\pm$7.32) 61.18($\pm$8.87) 56.85($\pm$11.72) 60.23($\pm$9.90) 50.65($\pm$6.85) 48.83($\pm$5.37) 60.74($\pm$7.77) 50.14($\pm$5.43) 46.66($\pm$4.81)
+Figure 3: Each square represents a value in the DCI matrix, which describes the relationship between the $i^{th}$ latent dimension and each factor.
+The size of each square is relative to the values within each row.
+The ideal case resembles a sparse matrix.
+The y-axis represents the factors of each dataset, while the x-axis corresponds to the latent vector dimensions.
+The number shown in each row of the matrix indicates the maximum value and standard deviation of that row.
+Higher maximum and standard deviation values suggest greater sparsity, indicating closer alignment with the ideal case.
+(a) Posterior visualization with Beta distribution
+(b) Posterior visualization of MIPET layer
+Figure 4: Non-Gaussian posterior learned by an IPE module without intended guidance to a specific distribution in a toy setting.
+We compare VAE, MIPET-VAE, and MIPET-VAE without a semantic mask to assess how well each model captures the underlying distribution.
+We construct the VAE with a 4-layer Multi-Layer Perceptron (MLP) as the encoder and a single linear layer as the decoder.
+Blue plots are randomly sampled from a two-dimensional beta distribution, red plots are the posterior, and black plots are the output results.
+6 Results and Discussion
+6.1 Quantitative Analysis
+Disentanglement Metrics
+We set the number of IPE-transformation functions to be equal to balancing hyper-parameter $\beta$ on $\beta$-VAE because of Eq.
+16.
+The number of IPE-transform functions of $\beta$-TCVAE is 3.
+However, in the case of CLG-VAE, we set it to 1 because its approach is based on the group theory, not directly controlling a KL divergence term such as $\beta$-VAE.
+We average each model performance value with 40, 20, 60, 10, and 30 cases in VAEs, $\beta$-TCVAEs, Control-VAE, Factor-VAE and CLG-VAEs, respectively.
+As shown in Table 1, MIPET-VAEs disentanglement performance is broadly improved with four metrics on each dataset.
+In particular, most FVM results significantly affect the model performance and stability on all datasets.
+Therefore, our proposed method obtains a specific dimension that corresponds to a specific single factor.
+These results imply that applied to MIPE-transformation functions on VAEs elaborate disentangled representation learning.
+Sensitivity to the Number of IPE-transformation and EF-conversion
+We analyze the impact of the MIPE-transformation function.
+As presented in Table 2, MIPE-transformation outperforms IPE-transformation in disentanglement learning across all datasets.
+Indeed, the results of MIPET-$\beta$-VAEs more generally and clearly demonstrate the impact of the MIPE-transfomation function.
+Our derivation in Section C clearly explains the impact of MIPE-transformation.
+This result shows the impact of the multiple uses of IPE-transformation and EF-conversion.
+Impact of Implicit Semantic Mask
+We set masking hyper-parameter $\lambda$ from $\{0.0,0.5,\cdots,2.0,\infty\}$, and each model has different $\lambda$ for best case.
+In Table 3, VAE and CLG-VAE with masked log-normalizer show better and well-balanced results than the models without masking, which implies improvement of disentanglement.
+Figure 5: Qualitative results on various datasets, which show the factors learned for each dimension of $z$.
+Ablation Study
+We conduct an ablation study to evaluate the separate impact of equivariant property and the EF-conversion.
+We have already presented the impact of the multiple uses of IPE-transform and EF-conversion in the previous paragraph.
+We evaluate the impact of the other properties by setting MIPE-transformation 1) without equivariant (w/o E), which is implemented as an asymmetric matrix, and 2) without EF-conversion (w/o EF).
+To exclude group theory interference with other methods, we select $\beta$-VAE and $\beta$-TCVAE.
+As the results are shown in Table 4, most of the results show that MIPET-VAEs performance is better than other cases.
+In particular, MIPET (w/o EF) results are lower than MIPET (w/o E) results and are clearly shown in all cases.
+6.2 Qualitative Analysis
+Relation Between Factors and Latent Vector Dimensions
+As shown in Fig.
+3, our model shows a close to sparse matrix compared to the $\beta$-VAE result in 3D Shapes, also the maximum value and standard deviation of each row are larger than $\beta$-VAE.
+This indicates that our model achieves better alignment of individual factors with specific latent dimensions compared to other methods.
+More details are in Appendix D.4.
+The Impact of IPE-transformation
+The IPE-transformation $\psi(\cdot)$ is designed to achieve two objectives: 1) to represent a non-Gaussian exponential family distribution as the posterior, and 2) to preserve the inherent structure of the dataset.
+To assess the first objective, we construct a toy dataset sampled from a beta distribution (additional cases are provided in Appendix D.7) and configure the decoder as a single linear layer, noting that the linear transformation of a Beta distribution remains within the same family.
+As shown in Fig.
+4(a), the posterior learned by MIPET-VAE (red dots) closely approximates the beta distribution, whereas the posteriors of VAE and MIPET (without the semantic mask) retain Gaussian distribution.
+Regarding the second objective, MIPET-VAE exhibits improved preservation of the dataset structure relative to VAE (black dots).
+Additional experimental details are reported in Appendix D.7.
+Does Each IPE-transformation Independently Operate?
+We demonstrate how each IPE-transformation represents the exponential family.
+As shown in Fig.
+4(b), each IPE-transformation represents a different distribution because the green and red plots are distributed in different posteriors, even though all IPE-transformations share the same decoder weights.
+Disentanglement of Factors Across Latent Dimensions
+We randomly sample an image for each dimension of the latent vector space and creates 5 variants of its generated latent vector by selecting values from $\{$-2, 2}\} with 5 intervals for the dimension, then generate their corresponding output images.
+For the generation, we select $\beta$-TCVAE (6), which shows the best FVM scores in dSprites dataset.
+Thereafter, we evaluate the semantic roles of each dimension before and after applying MIPE-transformation function.
+In Fig.
+5(a), $\beta$-TCVAE struggles with rotation and y-position, as shown on the 3rd row, and with scale and shape represented on the 4th row.
+On the contrary, MIPET-$\beta$-TCVAE separates rotation and y-position factor (4th, and 5th rows).
+Applied our method on $\beta$-TCVAE shows better disentangled representation on dSprites dataset.
+These results also show that our proposed method improves disentangled representation learning.
+As shown in the Fig.
+5(b), CLG-VAE struggles with shape and wall color factors in 2nd dimension, and shape and object color factors in 3rd dimension.
+In particular, it struggles with three factors in 5th dimension.
+On the other hand, MIPET-CLG-VAE separates shape, wall, and object color factors.
+The qualitative analysis with 3D Cars dataset, as shown in Fig.
+5(c), the left side is the $\beta$-TCVAE result, and it struggles with azimuth and body factors shown in the 2nd row.
+However, MIPET-$\beta$-TCVAE separates azimuth (4th row) and body (1st row).
+In particular, MIPET-$\beta$-TCVAE learns color factor (2rd row) which does not exist on $\beta$-TCVAE.
+Furthermore, our model shows improvement on the CelebA Liu et al.
+(2015) dataset, as shown in Fig.
+5(d).
+$\beta$-VAE struggles with most factors of the dataset, but ours separates all factors.
+7 Conclusion
+In this paper, we address the problem of injecting inductive bias for learning unsupervised disentangled representations.
+To build the bias in VAE frameworks, we propose MIPE-transformation composed of 1) IPE-transformation for the benefits of invertibility and partial-equivariance for disentanglement, 2) a training loss and module to adapt unrestricted prior and posterior to an approximated exponential family, and 3) integration of multiple units of IPE-transformation function and EF-conversion for more expressive bias.
+The method is easily equipped on state-of-the-art VAEs for disentanglement learning and shows significant improvement on dSprites, 3D Shapes, and 3D Cars datasets.
+We expect that our method can be applied to more VAEs, and extended to downstream applications.
+Acknowledgments
+This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.2022R1A2C2012054, Development of AI for Canonicalized Expression of Trained Hypotheses by Resolving Ambiguity in Various Relation Levels of Representation Learning).
+Checklist
+The checklist follows the references.
+For each question, choose your answer from the three possible options: Yes, No, Not Applicable.
+You are encouraged to include a justification to your answer, either by referencing the appropriate section of your paper or providing a brief inline description (1-2 sentences).
+Please do not modify the questions.
+Note that the Checklist section does not count towards the page limit.
+Not including the checklist in the first submission won’t result in desk rejection, although in such case we will ask you to upload it during the author response period and include it in camera ready (if accepted).
+In your paper, please delete this instructions block and only keep the Checklist section heading above along with the questions/answers below.
+  1. 1.
+    For all models and algorithms presented, check if you include:
+    1. (a)
+      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes - We present in Section 4, and Appendix B.]
+    2. (b)
+      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes - We present in the Appendix D.6.]
+    3. (c)
+      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes - We submit our code as supplement materials.]
+  2. 2.
+    For any theoretical claim, check if you include:
+    1. (a)
+      Statements of the full set of assumptions of all theoretical results. [Yes - We present in Section 4, and Appendix B.]
+    2. (b)
+      Complete proofs of all theoretical results. [Yes - We present in Section 4, and Appendix B.]
+    3. (c)
+      Clear explanations of any assumptions. [Yes - We present in Section 4, and Appendix B.]
+  3. 3.
+    For all figures and tables that present empirical results, check if you include:
+    1. (a)
+      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes - We submit our code as supplement materials.]
+    2. (b)
+      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes - We present in Section 5 and Appendix D.]
+    3. (c)
+      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes - We present in Section 6.]
+    4. (d)
+      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes - We present in Appendix D.]
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
+    1. (a)
+      Citations of the creator If your work uses existing assets. [Yes - We submit our code as supplement materials.]
+    2. (b)
+      The license information of the assets, if applicable. [Yes - We submit our code as supplement materials.]
+    3. (c)
+      New assets either in the supplemental material or as a URL, if applicable. [Yes - We submit our code as supplement materials.]
+    4. (d)
+      Information about consent from data providers/curators. [Yes - We present in Appendix D.]
+    5. (e)
+      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects, check if you include:
+    1. (a)
+      The full text of instructions given to participants and screenshots. [Not Applicable]
+    2. (b)
+      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
+    3. (c)
+      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
+MIPET Supplementary Materials
+Appendix A Notation and Groups
+Table 5: Terms and Notations
+$\psi(\cdot)$ Invertible function
+$\bm{z}$} Latent vector from encoder
+$\hat{{\bm{z}}}_{m}$ Transformed latent vector by $\psi_{m}(\cdot)$
+$\bm{\theta}_{\hat{{\bm{z}}}_{m}}$ Natural Parameter of posterior
+$\hat{\bm{\epsilon}}_{m}$ Transformed prior samples by $\psi_{m}(\cdot)$
+$\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}$ Natural Parameter of prior
+AA Log-Normalizer
+$T$ Sufficient Statistics
+$\bm{\nu}$ Evidence
+$f_{{\bm{x}}}(\cdot)$ Power Density Function
+$GL_{n}(\mathbb{R})$ General Linear Group
+$M_{n}(\mathbb{R})$ A set of $n\times n$ real matrix
+$Sym_{n}(\mathbb{R})$ A set of $n\times n$ symmetric real matrix
+$E_{M}$ $\{\mathbf{e}^{M}|M\in M_{n}(\mathbb{R})\}$
+$E_{S}$ $\{\mathbf{e}^{S}|S\in Sym_{n}(\mathbb{R})\}$
+$\psi_{M}(\cdot)$ $\psi_{M}(\cdot)\in M_{n}(\mathbb{R})$
+$\psi_{E_{S}}(\cdot)$ $\psi_{E_{S}}(\cdot)\in E_{S}$
+$\psi_{E_{M}}(\cdot)$ $\psi_{E_{M}}(\cdot)\in E_{M}$
+$G_{S}$ $G_{S}:(\mathbf{e}^{S},\ast)$
+$G_{I}$ Group of input space for symmetries
+$G_{L}$ Group of latent space for symmetries
+$G_{T}$ Group of transformed latent space for symmetries
+$J$ $G_{S}\cap G_{L}$
+$G_{I}^{J}$ Group of input space for symmetries corresponds to $J$
+$G_{L}^{J}$ Group of latent space for symmetries corresponds to $J$
+$G_{L}^{J}$ Group of transformed latent space for symmetries corresponds to $J$
+$\mathcal{X}$ Input space
+$\mathcal{Z}$ Latent vector space
+$\hat{\mathcal{Z}}$ Transformed latent vector space
+$\mathcal{Z}^{\prime}$ Homogeneous space of group $G_{L}$ (homogeneous $G_{L}$-space), $\mathcal{Z}^{\prime}\subseteq\mathcal{Z}$
+$\hat{\mathcal{Z}}^{\prime}$ Homogeneous space of group $G_{T}$ (homogeneous $G_{T}$-space), $\hat{\mathcal{Z}}^{\prime}\subseteq\hat{\mathcal{Z}}$
+$\hat{\mathcal{Z}}^{J}$ Homogeneous space of group $G_{T}^{J}$ (homogeneous $G_{T}^{J}$-space), $\hat{\mathcal{Z}}^{J}\subseteq\hat{\mathcal{Z}}$
+$D_{\mathrm{KL}}(\cdot||\cdot)$ Kullback-Leibler divergence
+$\mathbf{0}$ zero vector
+$\mathbf{0}_{n,n}$ n by n zero matrix
+Notation.
+We denote $g_{I}\in G_{I}$, $g_{L}\in G_{L}$, and $g_{T}\in G_{T}$.
+Let $M_{n}(\mathbb{R})$ be the set of $n\times n$ real matrices and $\mathrm{Sym}_{n}(\mathbb{R})$ the set of real symmetric matrices.
+For any set $\mathcal{A}\subseteq\mathrm{Sym}_{n}(\mathbb{R})$, define $\exp(\mathcal{A}):=\{\exp(S)\mid S\in\mathcal{A}\}$.
+We write $E_{M}:=\{\exp(M)\mid M\in M_{n}(\mathbb{R})\}$ and $E_{S}:=\exp(\mathrm{Sym}_{n}(\mathbb{R}))$.
+Standing assumption on $G_{S}$.
+We assume a pairwise-commuting symmetric set $\mathcal{S}\subseteq\mathrm{Sym}_{n}(\mathbb{R})$, i.e., $S_{1},S_{2}]=0$ for all $S_{1},S_{2}\in\mathcal{S}$.
+Then we define
+$G_{S}:=\exp(\mathcal{S})=\{\exp(S)\mid S\in\mathcal{S}\}\subset\mathrm{GL}_{n}(\mathbb{R}),$
+with the group operation given by the usual matrix multiplication $\ast$”.
+Appendix B Proof
+Lemma B.1 (Abelian Lie subgroup).
+If $\mathcal{S}\subseteq\mathrm{Sym}_{n}(\mathbb{R})$ is pairwise-commuting, then $(G_{S},\ast)$ is an abelian Lie subgroup of $\mathrm{GL}_{n}(\mathbb{R})$.
+In particular, $G_{S}$ is closed under multiplication and inversion, has identity $I=\exp(0)$, and is commutative.
+Proof sketch.
+Since all $S\in\mathcal{S}$ commute and are symmetric, they are simultaneously diagonalizable by some orthogonal matrix $Q$: $S_{i}=Q\Lambda_{i}Q^{\top}$ with diagonal $\Lambda_{i}$.
+Hence $\exp(S_{i})=Q\exp(\Lambda_{i})Q^{\top}$ are symmetric positive definite and commute.
+Moreover, for any $S_{1},S_{2}\in\mathcal{S}$, the Baker–Campbell–Hausdorff formula reduces to $\exp(S_{1})\exp(S_{2})=\exp(S_{1}+S_{2})$ because $S_{1},S_{2}]=0$.
+Since $S_{1}+S_{2}\in\mathcal{S}$, we have closure.
+The inverse is $\exp(S)^{-1}=\exp(-S)$ and $S\in\mathcal{S}$.
+Commutativity follows from $S_{1},S_{2}]=0$.
+Smoothness of $\exp$ implies $G_{S}$ is a (abelian) Lie subgroup.
+∎
+Proposition B.2.
+By the Lemma B.1, if $q_{\phi}$ is equivariant over defined on group of symmetries $G_{I}^{J}$ and $G_{L}^{J}$, then $\psi_{G_{S}}(q_{\phi}(\cdot))$ is equivariant to symmetries in $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$ by the equivariance of $q_{\phi}$.
+Proof.
+The function $\psi_{G_{S}}(\cdot)$ is an equivariant function over group elements in $G_{S}\cap G_{L}$ by Proposition 4.1.
+Then, the composite function, $\psi_{G_{S}}(\cdot)$ and $q_{\phi}$, is an equivariant function of $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$.
+Let $g_{L}^{J}$ be a group element in $G_{S}\cap G_{L}$, and $g_{I}^{J}$ is a group element in $G_{I}$ corresponding to $G_{S}\cap G_{L}$, and $g_{T}^{J}$ is a group element where corresponding to $G_{S}\cap G_{L}$ on the latent vector space transformed from the original latent vector space.
+Then, group element $g_{T}^{J}$ is equal to $g_{L}^{J}$:
+$\hat{{\bm{z}}}_{1}=\psi_{G_{S}}({\bm{z}}_{1}),\text{and}$ (17)
+[…]
+$\begin{split}&\text{then}~g_{L}^{J}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{J}\psi_{G_{S}}({\bm{z}}_{1})~(\because\hat{{\bm{z}}}_{2}=g_{T}^{J}\hat{{\bm{z}}}_{1})\\ &\Rightarrow(g_{L}^{J}-g_{T}^{J})\psi_{G_{S}}({\bm{z}}_{1})=\mathbf{0},\end{split}$ (19)
+[…]
+Then, $\psi_{G_{S}}(g_{L}^{J}{\bm{z}}_{1})=g_{L}^{J}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{J}\psi_{G_{S}}({\bm{z}}_{1})$.
+The encoder is an equivariant function over input space $\mathcal{X}$ as $q_{\phi}(g_{I}^{J}{\bm{x}}_{1})=g_{L}^{J}q_{\phi}({\bm{x}}_{1})$.
+Mixing two equivarience property, we can derive another equivariance relation $g_{T}^{J}\psi_{G_{S}}(q_{\phi}({\bm{x}}_{1}))=\psi_{G_{S}}(q_{\phi}(g_{I}^{J}{\bm{x}}_{1}))$ This result implies that the equivariance between input space and a latent space is preserved for $G_{S}\cap G_{L}$ if the latent vector $\bm{z}$} is transformed by $\psi_{G_{S}}.\blacksquare$
+We show that $\psi_{G_{S}}$ preserves equivariance between $G_{L}^{J}$ and $G_{I}^{J}$.
+If there exists equivariant function between input and latent vector space, there should be a group $G_{L}$ for a latent space and its corresponding group $G_{I}$ in an input space by definition of equivariance ($q_{\phi}(g_{I}x)=g_{L}q_{\phi}(x)$).
+In other words, $\psi_{G_{S}}(\cdot)$ guarantees to preserve the equivariance of I2L-transformation to certain symmetries in $G_{S}\cap G_{L}$ after IPE-transformation as shown in Fig.
+2.
+Let $P(G):=\frac{\mu(G\cap K)}{\mu(K)}$ by the assumption 4 in section 4, and $Pr(\psi_{G^{\prime}}(\cdot)\in G)=P(\psi(\cdot)\in G|\psi(\cdot)\in G^{\prime})=\frac{P(G\cap G^{\prime})}{P(G^{\prime})}$ after VAE training.
+Then,
+Proposition B.3.
+Under Assumption 4, the following probability inequality holds: $Pr(\psi_{E_{S}}(\cdot)\in G_{S})>Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})$, where probabilities are defined with respect to the normalized uniform measure on a compact subset $K\subset M_{n}(\mathbb{R})$.
+Proof. Since $G_{S}\subset E_{S}\subset E_{M}\subset M_{n}(\mathbb{R})$,
+$\Pr(\psi_{E_{S}}(\cdot)\in G_{S})=\frac{P(G_{S}\cap E_{S})}{P(E_{S})}=\frac{P(G_{S})}{P(E_{S})}=\frac{\mu(G_{S}\cap K)}{\mu(E_{S}\cap K)},\,\,\big(\because G_{S}\subset E_{S}\big)$
+$\Pr(\psi_{E_{M}}(\cdot)\in G_{S})=\frac{P(G_{S}\cap E_{M})}{P(E_{M})}=\frac{P(G_{S})}{P(E_{M})}=\frac{\mu(G_{S}\cap K)}{\mu(E_{M}\cap K)},\,\,\big(\because G_{S}\subset E_{M}\big)$
+$\Pr(\psi_{M}(\cdot)\in G_{S})=\frac{P(G_{S}\cap M_{n}(\mathbb{{R}}))}{P(M_{n}(\mathbb{{R}}))}=\frac{P(G_{S})}{P(M_{n}(\mathbb{R}))}=\frac{\mu(G_{S}\cap K)}{\mu(M_{n}(\mathbb{R})\cap K)},\,\,\big(\because G_{S}\subset M_{n}(\mathbb{R})\big)$
+where $\mu$ is the normalized uniform measure on $K$. Because $G_{S}\subseteq E_{S}\subseteq E_{M}\subseteq M_{n}(\mathbb{R})$ and all denominators are finite by compactness of $K$, the ratios satisfy
+$\frac{\mu(G_{S}\cap K)}{\mu(E_{S}\cap K)}>\frac{\mu(G_{S}\cap K)}{\mu(E_{M}\cap K)}>\frac{\mu(G_{S}\cap K)}{\mu(M_{n}(\mathbb{R})\cap K)}.$
+Thus the stated inequality follows. $\blacksquare$
+B.1 KL Divergence for Evidence of Lower Bound
+The KL divergence of Gaussian distribution (Kingma and Welling, 2013) is computed using mean and variance, which are the parameters of a Gaussian distribution.
+To introduce a loss as the KL divergence of Gaussian distribution, we compute KL divergence of the exponential family in Eq.
+6 using the learnable parameter $T(\cdot)$ and $A(\cdot)$ with given natural parameter $\bm{\theta}_{\hat{{\bm{z}}}}$ and $\bm{\theta}_{\hat{\bm{\epsilon}}}$, expressed as:
+[…]
+Because $)D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}$)) is followed as:
+$\begin{split}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\mathsf{d}{\bm{x}}\\ &-\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\mathsf{d}{\bm{x}}.\end{split}$ (21)
+We designed sufficient statistics as matrix multiplication (multi-layer perceptron). Then,
+$\begin{aligned} \int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\mathsf{d}{\bm{x}}&=\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\cdot\\ &~~~~~~~~~~~~~~[\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\mathbf{T}({\bm{x}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+B({\bm{x}})]\text{d}{\bm{x}}\end{aligned}$
+Expanding the Eq. 22, and $\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}=1$, then
+$\begin{aligned} \int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\mathsf{d}{\bm{x}}&=-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}B({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}.\end{aligned}$
+Expanding $\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\mathsf{d}{\bm{x}}$ using Eq.
+22 and 23 and plug into Eq.
+21 with Eq.
+23, we have
+$\begin{split}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})\\ &+\bm{\theta}_{\hat{{\bm{z}}}}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &-\bm{\theta}_{\hat{\bm{\epsilon}}}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}.\end{split}$ (24)
+The gradient of the log-partition function with respect to the natural parameters yields the expected value of the sufficient statistics: $\nabla_{\bm{\theta}}A(\bm{\theta})=\mathbb{E}_{p(x;\bm{\theta})}[T(x)]$ (Bishop, 2006).
+Then the mean of the sufficient statistic is followed as:
+$\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta})\text{d}{\bm{x}}=\frac{\partial A^{\ast}(\bm{\theta})}{\partial\bm{\theta}}\approx\frac{\partial A(\bm{\theta})}{\partial\bm{\theta}}~~~~~\because A^{\ast}(\bm{\theta})=\bm{\theta}^{\intercal}A^{\ast},$
+where $A^{\ast}(\cdot)$ is a true log-partition function of the exponential family (ideal case of $A(\cdot)$).
+However, estimating $A^{\ast}$ is difficult, and there is no direct method without random samplings, such as mini-batch weighted sampling or mini-batch stratified sampling (Chen et al., 2018).
+Then, we approximate $A^{\ast}$ to $A$, and train $A$ to be close to $A^{\ast}$.
+Plugging Eq.
+25 into Eq.
+23 then we have
+$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}(\mathbf{x}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})&=-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{{\bm{z}}}})}{\partial\bm{\theta}_{\hat{{\bm{z}}}}}\\ &+\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})B({\bm{x}})\text{d}{\bm{x}},\end{split}$ (26)
+Applying same work to $\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\mathsf{d}{\bm{x}}$ and plug into Eq.
+21 with Eq.
+26, then the final Kullback-Leibler divergence of exponential family is followed as:
+[…]
+Appendix C Objective Function
+We mathematically extend IPE-transformation to MIPE-transformation, which is the equivalent process of $\beta$-VAE to enhance disentanglement.
+Each IPE-transformation function operates independently, then the reconstruction error for objective function is defined as:
+$\begin{aligned} \mathcal{L}_{rec}&\coloneqq\frac{1}{k}\sum_{i=1}^{k}\Bigg[\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,\,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg]\\ &=\frac{1}{k}\sum_{i=1}^{k}E_{q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}})),\end{aligned}$
+where $\hat{{\bm{z}}}_{i}=\psi_{i}({\bm{z}})$. Becuase the log likelihood of $p({\bm{x}})$ can be derived as follows:
+$\log p_{\theta}({\bm{x}})=\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}})\mathsf{d}\hat{{\bm{z}}}^{\prime},$ (29)
+[…]
+$\begin{aligned} \log p_{\theta}({\bm{x}})\geqq\int\prod_{i=1}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{aligned}$
+Expanding Eq. 30 then
+$\begin{aligned} \log p_{\theta}({\bm{x}})&\geqq\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\Big[\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})\\ &~~~+\log\frac{p(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\Big]\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{aligned}$
+The first term of Eq. 31 is expanded as:
+$p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})=\frac{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k}|{\bm{x}})p_{\theta}({\bm{x}})}{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})}.$ (32)
+Each IPE-transformation function operates independently, then all $\hat{{\bm{z}}}_{i}$ are independent and conditionally independent given $\bm{x}$}. Then we have
+$p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})=\frac{p_{\theta}({\bm{x}})\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})}{\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i})},$ (33)
+where $\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})=p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k}|{\bm{x}})$ and $\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i})=p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})$.
+Let replace $p_{\theta}({\bm{x}})$ as $p_{\theta}({\bm{x}})^{-(k-1)}\prod_{i=1}^{k}p_{\theta}({\bm{x}})$.
+Then we have
+$p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})=p_{\theta}({\bm{x}})^{-(k-1)}\prod_{i=1}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i}),$ (34)
+Plugging Eq. 34 into the first term of Eq. 31, then we have
+$\begin{aligned} &\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=-(k-1)\log p_{\theta}({\bm{x}})+\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\Big[\sum_{j=1}^{k}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{j})\Big]\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{aligned}$
+Applying same process between the Eq. 32 and Eq. 33, and plugging Eq. 35 into Eq. 31, then we have
+$\begin{aligned} k\log p_{\theta}({\bm{x}})&\geqq\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\Big[\sum_{j=1}^{k}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{j})\Big]\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &~~~+\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{\prod_{i=1}^{k}p(\hat{{\bm{z}}}_{i})}{\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})}\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{aligned}$
+Expanding the first term and inverse the second term of Eq. 36, then we have
+$\begin{aligned} \log p_{\theta}({\bm{x}})&\geq\frac{1}{k}\sum_{i=1}^{k}\Bigg[\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,\,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg]\\ &\quad-\sum_{i=1}^{k}D_{\mathrm{KL}}\left(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})\,\|\,p({\bm{z}}_{i})\right)\end{aligned}$
+Replacing $\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}$ as $1$, then we have:
+$\begin{aligned} \log p_{\theta}({\bm{x}})\geq\frac{1}{k}\Bigg[\sum_{i=1}^{k}\mathbb{E}_{q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})-k\,D_{\mathrm{KL}}\left(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})\,\|\,p(\hat{{\bm{z}}}_{i})\right)\Bigg]\end{aligned}$
+Therefore, we define ELBO as:
+$\begin{split}\mathcal{L}^{\prime}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})&=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi,\psi_{i}}({\bm{z}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}}))-\\ &\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})||p_{\psi_{i}}({\bm{z}})).\end{split}$ (39)
+However, following Eq.
+39, k samples are generated, and each sample is disentangled for different factors.
+We implement the output as the average of the sum of the k samples to obtain a single sample with a superposition effect from k samples.
+Moreover, the KL divergence term in Eq.
+39 represents that increasing number of MIPE-transformation is equal to increasing $\beta$ hyper-parameter in $\beta$-VAE (Higgins et al., 2017).
+The VAEs equipped with MIPE-transformation (MIPET-VAEs) can be trained with the following loss:
+$\begin{split}&\mathcal{L}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})=\mathcal{L}_{rec}-\mathcal{L}_{kl}-\mathcal{L}_{el}-\mathcal{L}_{cali}.\end{split}$ (40)
+Appendix D Details of Experimental Environment
+D.1 Model Architecture and Hyper-Parameter Tuning
+The details of model architecture is demonstrated in Table 7and 7. The details of hyper-parameter settings are in Table 5(f).
+Table 6: VAE architecture for dSprites dataset.
+Encoder Decoder
+Input $64\times 64$ binary image input $\in\mathbb{R}^{10}$
+$4\times 4$ conv. 32 ReLU. stride 2 FC. 128 ReLU.
+$4\times 4$ conv. 32 ReLU. stride 2 FC. $4\times 4\times 64$ ReLU.
+$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 64 ReLU. stride 2.
+$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 32 ReLU. stride 2.
+FC. 128. FC. $2\times 10$ $4\times 4$ upconv. 32 ReLU. stride 2.
+$4\times 4$ upconv. 1. stride 2
+Table 7: VAE architecture for 3D Shapes, and 3D Cars datasets. For exceptional case, CLG-VAE, we use ten dimension size on 3D Shapes dataset (Zhu et al., 2021).
+Encoder Decoder
+Input $64\times 64\times 3$ RGB image input $\in\mathbb{R}^{6}$ (3D Shapes), $\mathbb{R}^{10}$ (3D Cars)
+$4\times 4$ conv. 32 ReLU. stride 2 FC. 256 ReLU.
+$4\times 4$ conv. 32 ReLU. stride 2 FC. $4\times 4\times 64$ ReLU.
+$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 64 ReLU. stride 2.
+$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 32 ReLU. stride 2.
+FC. 256. FC. $2\times 10$ $4\times 4$ upconv. 32 ReLU. stride 2.
+$4\times 4$ upconv. 3. stride 2
+Table 8: Hyper-parameters for dSprites, 3D Cars, and 3D Shapes.
+The epochs for dSprites and 3D cars are 30 and 200, respectively.
+lr is learning rate, latent dim is latent vector size, group reconst is group reconstruction, and forward group is forward group pass.
+models hyper-parameters values
+common batch size 256
+epoch $\{30,200\}$
+optim Adam
+lr 4e-4
+lr for MIPET 4e-4
+weight decay 1e-4
+latent dim 10
+$\beta$-VAE #\# of IE and EF $\{1,2,4,10\}$
+$\beta$-TCVAE $\beta$ $\{4,6\}$
+#\# of IE and EF $\{1,3\}$
+$\alpha$, $\gamma$ 1.0
+CLG-VAE $\lambda_{\text{decomp}}$ 40
+$\lambda_{\text{hessian}}$ 40
+forward group 0.2
+group reconst $\{0.2,0.5,0.7\}$
+models hyper-parameters values
+common batch size 256
+epoch 67
+optim Adam
+lr 4e-4
+lr for MIPET 4e-4
+$\beta$-VAE #\# of IE and EF $\{1,2,4,10\}$
+weight decay 0.0
+latent dim 6
+$\beta$-TCVAE $\beta$ $\{4,6\}$
+#\# of IE and EF $\{1,3\}$
+$\alpha$, $\gamma$ 1.0
+weight decay 1e-4
+latent dim 6
+CLG-VAE $\lambda_{\text{decomp}}$ 40
+$\lambda_{\text{hessian}}$ 40
+forward group 0.2
+group reconst $\{0.2,0.5,0.7\}$
+weight decay 0.0
+latent dim 10
+D.2 Details of Datasets
+We benchmark MIPET-VAEs against several standard VAEs on the following datasets: 1) dSprites (Matthey et al., 2017) which consists of 737,280 binary $64\times 64$ images of dSprites with five independent ground truth factors(number of values), $i.e.$ shape(3), orientation(40), scale(6), x-position(32), and y-position(32).
+2) 3D Shapes (Burgess and Kim, 2018) which consists of 480,000 RGB $64\times 64\times 3$ images of 3D Shapes with six independent ground truth factors: shape(4) orientation(15), scale(8), wall color(10), floor color(10), and object color(10).
+3) 3D Cars (Reed et al., 2015) which consists of 17,568 RGB $64\times 64\times 3$ images of 3D Shapes with three independent ground truth factors: car models(183), azimuth directions(24), and elevations(4).
+D.3 Details of Training and Evaluation
+Training
+We set 256 mini-batch size in the datasets (dSprites, 3D Shapes, and 3D Cars), Adam optimizer with learning rate $4\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and epochs from $\{30,67,200\}$ as a common setting for all the comparative methods.
+For the comparison, we follow training and inference on the whole dataset.
+We train each model for 30, 67, and 200 epochs on the dSprites, 3D Shapes, and 3D Cars, respectively, as introduced in Kim and Mnih (2018); Ren et al.
+(2022).
+We tune $\beta$ from $\{1,2,4,10\}$ and $\{4,6\}$ for $\beta$-VAE and $\beta$-TCVAE, respectively.
 We apply the proposed method to $\beta$-TCVAE only with the EF similarity loss term because $\beta$-TCVAE penalizes the divided KL divergence terms.
 We set the same encoder and decoder architecture in each model to exclude the overlapped effects.
-Also, we follow the same model architecture which are introduced in previous works Kim & Mnih (2018) and model details are in Table 8-8.
-4.0.2 Datasets
-We compare well-known VAEs to CHIC-VAEs: VAE, $\beta$-VAE, $\beta$-TCVAE, and CLG-VAE on the following data sets with 1) dSprites Matthey et al.
-(2017) which consists of 737,280 binary $64\times 64$ images of dSprites with five independent ground truth factors(number of values), $i.e.$ shape(3), orientation(40), scale(6), x-position(32), and y-position(32).
-2) 3D Shapes Burgess & Kim (2018) which consists of 480,000 RGB $64\times 64\times 3$ images of 3D Shapes with six independent ground truth factors: shape(4) orientation(15), scale(8), wall color(10), floor color(10), and object color(10).
-3) 3D Cars Reed et al.
-(2015) which consists of 17,568 RGB $64\times 64\times 3$ images of 3D Shapes with three independent ground truth factors: car models(183), azimuth directions(24), and elevations(4).
-4.0.3 Training
-We set 256 mini-batch size in the datasets (dSprites, 3D Shapes, and 3D Cars), Adam optimizer with learning rate $4\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and epochs from $\{30,67,200\}$ as a common setting for all the comparative methods.
-For the comparison, we follow training and inference on the whole dataset.
-We train each model for 30, 67, and 200 epochs on the dSprites, 3D Shapes, and 3D Cars, respectively, as introduced in Kim & Mnih (2018); Ren et al.
+The $\alpha=1$ and $\gamma=1$ of $\beta$-TCVAE as Chen et al.
 (2018).
-We tune $\beta$ from $\{1,2,4,10\}$ and $\{4,6\}$ for $\beta$-VAE and $\beta$-TCVAE, respectively.
 We set the dimension size of the latent vectors from $\{6,10\}$ for 10 on dSprites and 3D Cars datasets and 6 for 3D Shapes, but we set 10 for CLG-VAE because it sets 10 dimensions size on 3D Shapes in Zhu et al.
@@
 To check the impact of MIPE-transformation, we do not consider the Groupified VAE because the latter is implemented with an extended decoder (different capacity).
-4.0.4 Evaluation
+Evaluation
 We conduct experiments on NVIDIA A100, RTX 2080 Ti, and RTX 3090.
-We set 100 samples to evaluate global empirical variance in each dimension and run it a total of 800 times to estimate the FVM score introduced in Kim & Mnih (2018).
-For the MIG Chen et al.
-(2018), SAP Kumar et al.
-(2018), and DCI Eastwood & Williams (2018), we follow default values introduced in Michlo (2021), training and evaluation 100 and 50 times with 100 mini-batches, respectively.
+We set 100 samples to evaluate global empirical variance in each dimension and run it a total of 800 times to estimate the FVM score introduced in Kim and Mnih (2018).
+For the MIG (Chen et al., 2018), SAP (Kumar et al., 2018), and DCI (Eastwood and Williams, 2018), we follow default values introduced in Michlo (2021), training and evaluation 100 and 50 times with 100 mini-batches, respectively.
 We evaluate four disentanglement metrics for a less biased understanding of the actual states of disentanglement.
@@
 dataset Metrics
-FVM IMG SAP DCI
-dSprites [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+FVM MIG SAP DCI
 3D Shapes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 3D Cars [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 5: Impact of the mask (mean±$\pm$±std.) and its ratio $\lambda$ in Eq.
-8 on 3D Cars.
-($\infty$: no masking case, gray box: the best setting over all metrics, bold text: the best in each metric.) Each model runs with ten random seeds.
-ratio $\beta$-VAE (1) CLG-VAE (0.5)
-$\lambda$ FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$ FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DIC $\uparrow$
-0.0 90.46($\pm$6.50) 4.84($\pm$2.32) 1.29($\pm$0.81) 16.76($\pm$4.68) 90.06($\pm$4.44) 9.28($\pm$2.09) 1.82($\pm$0.82) 19.12($\pm$3.41)
-0.5 91.35($\pm$5.52) 5.37($\pm$2.74) 1.17($\pm$0.67) 16.65($\pm$3.76) 88.69($\pm$4.78) 6.90($\pm$1.96) 1.85($\pm$0.67) 17.52($\pm$3.16)
-1.0 91.78($\pm$6.20) 4.99($\pm$2.27) 1.36($\pm$0.81) 16.50($\pm$2.53) 83.60($\pm$11.48) 8.12($\pm$3.66) 2.37($\pm$1.50) 17.07($\pm$3.89)
-1.5 90.04($\pm$5.88) 7.22($\pm$2.87) 1.36($\pm$0.48) 18.23($\pm$2.84) 84.76($\pm$6.86) 7.70($\pm$2.11) 2.05($\pm$0.73) 17.06($\pm$2.77)
-2.0 87.79($\pm$8.88) 4.75($\pm$2.49) 1.01($\pm$0.99) 16.64($\pm$3.75) 85.78($\pm$4.18) 7.83($\pm$1.79) 1.91($\pm$0.96) 17.26($\pm$2.07)
-$\infty$ 89.43($\pm$11.72) 3.74($\pm$2.32) 0.77($\pm$0.39) 15.45($\pm$4.59) 82.96($\pm$11.84) 8.07($\pm$2.52) 2.32($\pm$1.02) 17.46($\pm$4.07)
-5 Results and Discussion
-5.1 Quantitative Analysis
-5.1.1 Disentanglement Metrics
-We set the number of IPE-transformation functions to be equal to balancing hyper-parameter $\beta$ on $\beta$-VAE because of Eq.
-11.
-The number of IPE-transform functions of $\beta$-TCVAE is 3.
-However, in the case of CLG-VAE, we set it to 1 because its approach is based on the group theory, not directly controlling a KL divergence term such as $\beta$-VAE.
-We average each model performance value with 40, 20, 60, and 30 cases in VAEs, $\beta$-TCVAEs, Control-VAE and CLG-VAEs, respectively.
-As shown in Table 2, MIPET-VAEs disentanglement performance is broadly improved with four metrics on each dataset.
-In particular, most FVM results significantly affect the model performance and stability on all datasets.
-Therefore, our proposed method obtains a specific dimension that corresponds to a specific single factor.
-These results imply that applied to MIPE-transformation functions on VAEs elaborate disentangled representation learning.
-We additionally estimate the p-value of each metrics over models in Table 3. Previous work shows the average case of each models Yang et al. (2022).
-[…]
-Table 6: Ablation study for the equivariant property (w/o E), and EF-conversion (w/o EF). Each metric is averaged over 40 and 20 settings of $\beta$-VAE and $\beta$-TCVAE, respectively.
-3D Shapes $\beta$-VAE $\beta$-TCVAE
-MIPET MIPET (w/o E) MIPET (w/o EF) MIPET MIPET (w/o E) MIPET (w/o EF)
-FVM 75.19($\pm$8.16) 74.91($\pm$10.46) 22.27($\pm$1.29) 80.59($\pm$8.57) 77.90($\pm$8.66) 66.38($\pm$7.57)
-MIG 47.37($\pm$10.13) 47.45($\pm$8.98) 0.28($\pm$0.09) 54.49($\pm$9.44) 51.37($\pm$11.54) 36.08($\pm$17.42)
-SAP 9.20($\pm$2.44) 9.43($\pm$2.59) 0.26($\pm$0.07) 11.58($\pm$3.32) 10.23($\pm$3.13) 7.13($\pm$3.09)
-DCI 54.95($\pm$8.99) 54.23($\pm$9.05) 0.10($\pm$0.02) 66.22($\pm$7.32) 61.18($\pm$8.87) 56.85($\pm$11.72)
-5.1.2 Sensitivity to the Number of IPE-transformation and EF-conversion
-We investigate the impact of the MIPE-transformation function.
-As presented in Table.
-4, MIPE-transformation is better than IPE-transformation for disentanglement learning on each dataset.
-Indeed, MIPET-$\beta$-VAEs results more generally and clearly show the impact of the MIPE-transfomation function.
-Our derivation in Section C.1 clearly explains MIPE-transformation impact.
-This result shows the impact of the multiple uses of IPE-transformation and EF-conversion.
-5.1.3 Impact of Implicit Semantic Mask
-We set masking hyper-parameter $\lambda$ from $\{0.0,0.5,\cdots,2.0,\infty\}$, and each model has different $\lambda$ for best case.
-In Table 5, VAE and CLG-VAE with masked log-normalizer show better and well-balanced results than the models without masking, which implies improvement of disentanglement.
-5.1.4 Ablation Study
-We conduct an ablation study to evaluate the separate impact of equivariant property and the EF-conversion.
-We have already presented the impact of the multiple uses of IPE-transform and EF-conversion in the previous paragraph.
-We evaluate the impact of the other properties by setting MIPE-transformation 1) without equivariant (w/o E), which is implemented as an asymmetric matrix, and 2) without EF-conversion (w/o EF).
-To exclude group theory interference with other methods, we select $\beta$-VAE and $\beta$-TCVAE.
-As the results are shown in Table 6, most of the results show that MIPET-VAEs performance is better than other cases.
-In particular, MIPET (w/o EF) results are lower than MIPET (w/o E) results and are clearly shown in all cases.
-Figure 4: Qualitative results on dSprites.
-The left-side grids are input images and their variants by changing activations of each dimension of latent vectors.
-The first row shows input images.
-The right-side table shows matching pre-defined factors of the dataset (red: MIPET, blue: no MIPET).
-Figure 5: Qualitative analysis result of $\beta$-VAE (4.0) with 3D Cars.
-5.2 Qualitative Analysis
-[…]
-In Figure 4, $\beta$-TCVAE struggles with y-position and rotation, as shown on the 6th row, and with scale and shape represented on the 7th row.
-On the contrary, MIPET-$\beta$-TCVAE separates y-position and rotation factor (10th, and 7th rows), also the activated dimensions of MIPET-$\beta$-TCVAE are not overlapped with each factor.
-Applied our method on $\beta$-TCVAE shows better disentangled representation on dSprites dataset.
-These results also show that our proposed method improves disentangled representation learning.
-The qualitative analysis with 3D Cars dataset, as show in Figure 5, the left side is the $\beta$-TCVAE result, and it struggles with body, and azimuth factors shown in the 7th row.
-However, MIPET-$\beta$-TCVAE separates azimuth (6th row) and body (1st row).
-In particular, MIPET-$\beta$-TCVAE learns color factor (3rd row) which does not exist on $\beta$-TCVAE.
-6 Conclusion
-In this paper, we address the problem of injecting inductive bias for learning unsupervised disentangled representations.
-To build the bias in VAE frameworks, we propose MIPE-transformation composed of 1) IPE-transformation for the benefits of invertibility and partial-equivariant for disentanglement, 2) a training loss and module to adapt unrestricted prior and posterior to an approximated exponential family, and 3) integration of multiple units of IPE-transformation function and EF-conversion for more expressive bias.
-The method is easily equipped on state-of-the-art VAEs for disentanglement learning and shows significant improvement on dSprites, 3D Shapes, and 3D Cars datasets.
-We expect that our method can be applied to more VAEs, and extended to downstream applications.
-Our work is limited to holding partial equivariance of I2L transformation, so more direct methods to induce it can be integrated in the future.
-Acknowledgments
-This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (2022R1A2C2012054).
-Appendix A Appendix
-Appendix B Preliminaries
-B.1 Group Theory
-Binary operation: Binary operation on a set $S$ is a function that $\ast:S\times S\rightarrow S$, where $\times$ is a cartesian product.
-Group: A group is a set $G$ together with binary operation $\ast$, that combines any two elements $g_{a}$ and $g_{b}$ in $G$, such that the following properties:
-  • closure: $g_{a},g_{b}\in G\Rightarrow g_{a}\ast g_{b}\in G$.
-  • Associativity: $\forall g_{a},g_{b},g_{c}\in G$, s.t. $(g_{a}\ast g_{b})\ast g_{c}=g_{a}\ast(g_{b}\ast g_{c})$.
-  • Identity element: There exists an element $e\in G$, s.t. $\forall g\in G,e\ast g=g\ast e=g$.
-  • Inverse element: $\forall g\in G,\exists g^{-1}\in G$: $g\ast g^{-1}=g^{-1}\ast g=e$.
-Group action: Let $(G,\ast)$ be a group and set $X$, binary operation $\cdot:G\times X\rightarrow X$, such that following properties:
-  • Identity: $e\cdot x=x$, where $e\in G,x\in X$.
-  • Compatibility: $\forall g_{a},g_{b}\in G,x\in X,(g_{a}\ast g_{b})\cdot x=g_{a}\cdot(g_{b}\cdot x)$.
-Equivariant map: Let $G$ be a group and $X_{1},X_{2}$ be two sets with corresponding group action of $G$ in each sets: $T^{X_{1}}_{g},T^{X_{2}}_{g}$, where $g\in G$.
-Then a function $f:X_{1}\rightarrow X_{2}$ is equivariant if $f(T^{X_{1}}_{g}\cdot X_{1})=T^{X_{2}}_{g}\cdot f(X_{1})$.
-Partial Equivariance:Romero & Lohit (2022): Let subset of $G$ be $\Upsilon\subset G$, then $f$ is a partially equivariant map to $G$:
-$f(T^{X_{1}}_{\upsilon}\cdot X_{1})=T^{X_{2}}_{\upsilon}\cdot f(X_{1}),~{}\text{where}~{}\forall\upsilon\in\Upsilon.$ (12)
-Homomorphsim: Let $(G,\cdot),(H,\circ)$ be two groups. If mapping function $h:G\rightarrow H$, s.t. $h(g_{i}\cdot g_{j})=f(g_{i})\circ f(g_{j})$, then $f$ is called homomorphism.
-B.2 Exponential Family
-Power density function of the exponential family (PDF) generalized formulation:
-$\begin{split}f_{{\bm{x}}}({\bm{x}}|\bm{\theta})&=h({\bm{x}})\text{exp}(\bm{\theta}^{\intercal}T({\bm{x}})-A(\bm{\theta}))\\ &=\text{exp}(\bm{\theta}^{\intercal}T({\bm{x}})-A(\bm{\theta})+B({\bm{x}})),\end{split}$ (13)
-where sufficient statistics $T(\cdot)$, log-normalizer $A(\cdot)$, and carrier or base measure $B(\cdot)$ are known functions, samples ${\bm{x}}$ from distribution, and natural parameter $\bm{\theta}$.
-Appendix C Proof
-Proposition C.1.
-If $q_{\phi}$ is equivariant over defined on group of symmetries $G_{I}^{a}$ and $G_{L}^{a}$, then $\psi_{G_{S}}(q_{\phi}(\cdot))$ is equivariant to symmetries in $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$ by the equivariance of $q_{\phi}$.
-Proof.
-The function $\psi_{G_{S}}(\cdot)$ is an equivariant function over group elements in $G_{S}\cap G_{L}$ by Proposition 3.1.
-Then, the composite function, $\psi_{G_{S}}(\cdot)$ and $q_{\phi}$, is an equivariant function of $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$.
-Let $g_{L}^{a}$ be a group element in $G_{S}\cap G_{L}$, and $g_{I}^{a}$ is a group element in $G_{I}$ corresponding to $G_{S}\cap G_{L}$, and $g_{T}^{a}$ is a group element where corresponding to $G_{S}\cap G_{L}$ on the latent vector space transformed from the original latent vector space.
-Then, group element $g_{T}^{a}$ is equal to $g_{L}^{a}$:
-$\hat{{\bm{z}}}_{1}=\psi_{G_{S}}({\bm{z}}_{1}),\text{and}$ (14)
-[…]
-$\begin{split}&\text{then}~{}g_{L}^{a}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{a}\psi_{G_{S}}({\bm{z}}_{1})~{}(\because\hat{{\bm{z}}}_{2}=g_{T}^{a}\hat{{\bm{z}}}_{1})\\ &\Rightarrow(g_{L}^{a}-g_{T}^{a})\psi_{G_{S}}({\bm{z}}_{1})=\mathbf{0},\end{split}$ (16)
-where $\mathbf{0}$ is a zero vector.
-Eq.
-16 is defined when $\forall{\bm{z}}\in\mathcal{Z}$ by the equivariance definition.
-In other words, Eq.
-16 is satisfied only if the kernel (linear algebra) of $g_{L}^{a}-g_{T}^{a}$, notated as ker$(g_{L}^{a}-g_{T}^{a})$, includes the basis of $\mathbb{R}^{n}$ vector space.
-If the standard basis of $\mathbb{R}^{n}$ vector space is in ker$(g_{L}^{a}-g_{T}^{a})$, then $(g_{L}^{a}-g_{T}^{a})=\mathbf{0}_{n,n}$, where $\mathbf{0}_{n,n}$ is an n by n zero matrix.
-Other bases of $\mathbb{R}^{n}$ vector space are expressed by the standard basis.
-Therefore $g_{L}^{a}-g_{T}^{a}=\mathbf{0}_{n,n}$.
-Then, $\psi_{G_{S}}(g_{L}^{a}{\bm{z}}_{1})=g_{L}^{a}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{a}\psi_{G_{S}}({\bm{z}}_{1})$.
-The encoder is an equivariant function over input space $\mathcal{X}$ as $q_{\phi}(g_{I}^{a}{\bm{x}}_{1})=g_{L}^{a}q_{\phi}({\bm{x}}_{1})$.
-Mixing two equivarience property, we can derive another equivariance relation $g_{T}^{a}\psi_{G_{S}}(q_{\phi}({\bm{x}}_{1}))=\psi_{G_{S}}(q_{\phi}(g_{I}^{a}{\bm{x}}_{1}))$ This result implies that the equivariance between input space and a latent space is preserved for $G_{S}\cap G_{L}$ if the latent vector ${\bm{z}}$ is transformed by $\psi_{G_{S}}.$$\blacksquare$
-We show that $\psi_{G_{S}}$ preserves equivariance between $G_{L}^{a}$ and $G_{I}^{a}$.
-If there exists equivariant function between input and latent vector space, there should be a group $G_{L}$ for a latent space and its corresponding group $G_{I}$ in an input space by definition of equivariance ($q_{\phi}(g_{I}x)=g_{L}q_{\phi}(x)$).
-In other words, $\psi_{G_{S}}(\cdot)$ guarantees to preserve the equivariance of I2L-transformation to certain symmetries in $G_{S}\cap G_{L}$ after IPE-transformation as shown in Figure 2.
-Let $P(B)$ be the probability of $\psi(\cdot)\in B$ for a subset $B\subset M_{n}(\mathbb{R})$ after VAE training, and $Pr(\psi_{B}\in B^{\prime})$ be the conditional probability of $\psi(\cdot)\in B^{\prime}$ given $\psi(\cdot)\in B$.
-Then,
-Proposition C.2.
-$Pr(\psi_{E_{S}}(\cdot)\in G_{S})>Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})$.
-Proof.
-All $\mathbf{e}^{{\bm{S}}}\in E_{S}$ are in $E_{M}$ since $Sym_{n}(\mathbb{R})\subset M_{n}(\mathbb{R})$.
-However, $E_{M}\not\subset E_{S}$ because $\mathbf{e}^{{\bm{S}}}$ is always symmetric, but $\mathbf{e}^{{\bm{M}}}$ can be an asymmetric matrix.
-All elements of $E_{S}$ are symmetric because of the matrix exponential property that $\mathbf{e}^{{\bm{M}}^{\intercal}}=(\mathbf{e}^{{\bm{M}}})^{\intercal}$.
-If ${\bm{M}}$ is a symmetric matrix then $\mathbf{e}^{{\bm{M}}^{\intercal}}=\mathbf{e}^{{\bm{M}}}=(\mathbf{e}^{{\bm{M}}})^{\intercal}$.
-Therefore, if ${\bm{M}}$ is symmetric then the exponential of ${\bm{M}}$ is also symmetric.
-We show a counter example to $E_{M}\subset E_{S}$.
-When ${\bm{M}}=\begin{bmatrix}1&1\\ 0&1\end{bmatrix}$,
-$\begin{split}\mathbf{e}^{{\bm{M}}}&=\sum_{k=0}^{\infty}\frac{1}{k!}{\bm{M}}^{k}\\ &=I+\begin{bmatrix}1&1\\ 0&1\end{bmatrix}+\frac{1}{2!}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}^{2}+\cdots+\frac{1}{(n-1)!}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}^{(n-1)}+\cdots\\ &=I+\begin{bmatrix}\sum_{n=0}^{\infty}\frac{1}{n!}&1+\sum_{n=0}^{\infty}\frac{1}{(n-1)!}\\ 0&\sum_{n=0}^{\infty}\frac{1}{n!}\end{bmatrix}\\ &=\begin{bmatrix}1+e&1+e\\ 0&1+e\end{bmatrix}.\end{split}$ (17)
-The matrix $\mathbf{e}^{{\bm{M}}}$ is asymmetric and not in $E_{S}$.
-Therefore $E_{M}\not\subset E_{S}$.
-Therefore, the probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{S})}$ is greater than $Pr(\psi_{E_{M}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{M})}$.
-In the same way, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(M_{n}(\mathbb{R}))}$ because $E_{M}\subset M_{n}(\mathbb{R})$ and non-invertible functions are only in $M_{n}(\mathbb{R})$.
-$\blacksquare$
-C.0.1 KL Divergence for Evidence of Lower Bound
-The KL divergence of Gaussian distribution Kingma & Welling (2013) is computed using mean and variance, which are the parameters of a Gaussian distribution.
-To introduce a loss as the KL divergence of Gaussian distribution, we compute KL divergence of the exponential family in Eq.
-2 using the learnable parameter $T(\cdot)$ and $A(\cdot)$ with given natural parameter $\bm{\theta}_{\hat{{\bm{z}}}}$ and $\bm{\theta}_{\hat{\bm{\epsilon}}}$, expressed as:
-$\begin{split}\mathcal{L}_{kl}&\coloneqq D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\\ &=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{{\bm{z}}}}}A(\bm{\theta}_{\hat{{\bm{z}}}})-\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{\bm{\epsilon}}}}A(\bm{\theta}_{\hat{\bm{\epsilon}}}).\end{split}$ (18)
-Because $D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))$ is followed as:
-$\begin{split}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\mathsf{d}{\bm{x}}\\ &-\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\mathsf{d}{\bm{x}}.\end{split}$ (19)
-We designed sufficient statistics as matrix multiplication (multi-layer perceptron). Then,
-$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\mathsf{d}{\bm{x}}&=\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\cdot\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}[\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\mathbf{T}({\bm{x}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+B({\bm{x}})]\text{d}{\bm{x}}\\ &=-A(\bm{\theta}_{\hat{{\bm{z}}}})\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})[\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\mathbf{T}({\bm{x}})+B({\bm{x}})]\text{d}{\bm{x}}\\ &=-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}B({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}},\end{split}$ (20)
-and
-$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\mathsf{d}{\bm{x}}&=-A(\bm{\theta}_{\hat{\bm{\epsilon}}})\\ &+\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}B({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}.\end{split}$ (21)
-$\begin{split}\therefore D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})\\ &+\bm{\theta}_{\hat{{\bm{z}}}}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &-\bm{\theta}_{\hat{\bm{\epsilon}}}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}.\end{split}$ (22)
-The mean of the sufficient statistic is followed as:
-$\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta})\text{d}{\bm{x}}=\frac{\partial A^{\ast}(\bm{\theta})}{\partial\bm{\theta}}\approx\frac{\partial A(\bm{\theta})}{\partial\bm{\theta}}~{}~{}~{}~{}~{}\because A^{\ast}(\bm{\theta})=\bm{\theta}^{\intercal}A^{\ast},$ (23)
-where $A^{\ast}(\cdot)$ is a true log-partition function of the exponential family (ideal case of $A(\cdot)$).
-However, estimating $A^{\ast}$ is difficult, and there is no direct method without random samplings, such as mini-batch weighted sampling or mini-batch stratified sampling Chen et al.
-(2018).
-Then, we approximate $A^{\ast}$ to $A$, and train $A$ to be close to $A^{\ast}$.
-Consequently, we obtain KL divergence of the exponential family as:
-$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}(\mathbf{x}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})&=-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{{\bm{z}}}})}{\partial\bm{\theta}_{\hat{{\bm{z}}}}}\\ &+\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})B({\bm{x}})\text{d}{\bm{x}},\end{split}$ (24)
-$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}(\mathbf{x}|\bm{\theta}_{\hat{\bm{\epsilon}}})&=-Z(\bm{\theta}_{\hat{\bm{\epsilon}}})+\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{\bm{\epsilon}}})}{\partial\bm{\theta}_{\hat{\bm{\epsilon}}}}\\ &+\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})B({\bm{x}})\text{d}{\bm{x}}.\end{split}$ (25)
-Therefore, the final Kullback-Leibler divergence of exponential family is followed as:
-$\begin{split}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{{\bm{z}}}})}{\partial\bm{\theta}_{\hat{{\bm{z}}}}}-\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{\bm{\epsilon}}})}{\partial\bm{\theta}_{\hat{\bm{\epsilon}}}}.\end{split}$ (26)
-C.1 Integration for Multiple IPE-Transformation and EF-Conversion
-We mathematically extend IPE-transformation to MIPE-transformation, which is the equivalent process of $\beta$-VAE to enhance disentanglement.
-Each IPE-transformation function operates independently, then the reconstruction error for objective function is defined as:
-$\begin{split}\mathcal{L}_{rec}&\coloneqq\frac{1}{k}\sum_{i=1}^{k}\Bigg{[}\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg{]}\\ &=\frac{1}{k}\sum_{i=1}^{k}E_{q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}})),\end{split}$ (27)
-where $\hat{{\bm{z}}}_{i}=\psi_{i}({\bm{z}})$. Becuase the log likelihood of $p({\bm{x}})$ can be derived as follows:
-$\log p_{\theta}({\bm{x}})$ $=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}})\mathsf{d}\hat{{\bm{z}}}^{\prime}$ (28)
-$=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\mathsf{d}\hat{{\bm{z}}}^{\prime}$ (29)
-$\begin{split}&=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\cdot\\ &~{}~{}~{}~{}\Big{[}\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}-\log\frac{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\Big{]}\mathsf{d}\hat{{\bm{z}}}^{\prime}\end{split}$ (30)
-$\geqq\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\mathsf{d}\hat{{\bm{z}}}^{\prime}$ (31)
-$\begin{split}&=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\cdot\\ &~{}~{}~{}~{}\Big{[}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})+\log\frac{p(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\Big{]}\mathsf{d}\hat{{\bm{z}}}^{\prime},\end{split}$ (32)
-[…]
-$\begin{split}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})&=\frac{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k}|{\bm{x}})p_{\theta}({\bm{x}})}{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})}\\ &=\frac{p_{\theta}({\bm{x}})\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})}{\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i})}(\because(\hat{{\bm{z}}}_{i}\mathrel{\perp\!\!\!\perp}\hat{{\bm{z}}}_{j}s|{\bm{x}}))\\ &=\prod_{i=1}^{k}\frac{p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})p_{\theta}({\bm{x}}^{\frac{1}{k}})}{p_{\theta}(\hat{{\bm{z}}}_{i})}\\ &=p_{\theta}({\bm{x}})^{-(k-1)}\prod_{i=1}^{k}\frac{p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})p_{\theta}({\bm{x}})}{p_{\theta}(\hat{{\bm{z}}}_{i})}\\ &=p_{\theta}({\bm{x}})^{-(k-1)}\prod_{i=1}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i}),\end{split}$ (33)
-where $\hat{{\bm{z}}}_{j}s=\cap_{j=1,j\neq i}^{k}\hat{{\bm{z}}}_{j}$. Therefore,
-$\small\begin{split}&\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\Big{[}-(k-1)\log p_{\theta}({\bm{x}})+\prod_{i=1}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\Big{]}\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=-(k-1)\log p_{\theta}({\bm{x}})+\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\prod_{i=j}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{j})\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{split}$ (34)
-Then,
-$\begin{split}\log p_{\theta}({\bm{x}})&\geq\frac{1}{k}\sum_{i=1}^{k}\Bigg{[}\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg{]}\\ &-\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})}{\prod_{i=1}^{k}p(\hat{{\bm{z}}}_{i})}\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\\ &-\sum_{i=1}^{k}\bigg{[}D_{\mathrm{KL}}(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})||p(\hat{{\bm{z}}}_{i}))\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\bigg{]}\\ &=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})-\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})||p({\bm{z}}_{i}))\\ &=\frac{1}{k}\Big{[}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})-kD_{\mathrm{KL}}(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})||p(\hat{{\bm{z}}}_{i}))\Big{]}.\end{split}$ (35)
-Therefore, we define ELBO as:
-$\begin{split}\mathcal{L}^{\prime}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})&=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi,\psi_{i}}({\bm{z}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}}))-\\ &\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})||p_{\psi_{i}}({\bm{z}})).\end{split}$ (36)
-However, following Eq.
-36, k samples are generated, and each sample is disentangled for different factors.
-We implement the output as the average of the sum of the k samples to obtain a single sample with a superposition effect from k samples.
-Moreover, the KL divergence term in Eq.
-36 represents that increasing number of MIPE-transformation is equal to increasing $\beta$ hyper-parameter in $\beta$-VAE Higgins et al.
-(2017).
-The VAEs equipped with MIPE-transformation (MIPET-VAEs) can be trained with the following loss:
-$\begin{split}&\mathcal{L}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})=\mathcal{L}_{rec}-\mathcal{L}_{kl}-\mathcal{L}_{el}-\mathcal{L}_{cali}.\end{split}$ (37)
-Appendix D Details of Experimental Environment
-D.1 Model Architecture
-Table 7: VAE architecture for dSprites dataset.
-Encoder Decoder
-Input $64\times 64$ binary image input $\in\mathbb{R}^{10}$
-$4\times 4$ conv. 32 ReLU. stride 2 FC. 128 ReLU.
-$4\times 4$ conv. 32 ReLU. stride 2 FC. $4\times 4\times 64$ ReLU.
-$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 64 ReLU. stride 2.
-$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 32 ReLU. stride 2.
-FC. 128. FC. $2\times 10$ $4\times 4$ upconv. 32 ReLU. stride 2.
-$4\times 4$ upconv. 1. stride 2
-Table 8: VAE architecture for 3D Shapes, and 3D Cars datasets. For exceptional case, CLG-VAE, we ues ten dimension size on 3D Shapes dataset Zhu et al. (2021).
-Encoder Decoder
-Input $64\times 64\times 3$ RGB image input $\in\mathbb{R}^{6}$ (3D Shapes), $\mathbb{R}^{10}$ (3D Cars)
-$4\times 4$ conv. 32 ReLU. stride 2 FC. 256 ReLU.
-$4\times 4$ conv. 32 ReLU. stride 2 FC. $4\times 4\times 64$ ReLU.
-$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 64 ReLU. stride 2.
-$4\times 4$ conv. 64 ReLU. stride 2 $4\times 4$ upconv. 32 ReLU. stride 2.
-FC. 256. FC. $2\times 10$ $4\times 4$ upconv. 32 ReLU. stride 2.
-$4\times 4$ upconv. 3. stride 2
-D.2 Details of Setting
-Table 9: Hyper-parameters for dSprites, 3D Cars, and 3D Shapes.
-The epochs for dSprites and 3D cars are 30 and 200, respectively.
-lr is learning rate, latent dim is latent vector size, group reconst is group reconstruction, and forward group is forward group pass.
-(a) Hyper-parameters for dSprites and 3D Cars
-models hyper-parameters values
-common batch size 256
-epoch $\{30,200\}$
-optim Adam
-lr 4e-4
-lr for MIPET 4e-4
-weight decay 1e-4
-latent dim 10
-$\beta$-VAE ##\## of IE and EF $\{1,2,4,10\}$
-$\beta$-TCVAE $\beta$$\{4,6\}$
-##\## of IE and EF $\{1,3\}$
-$\alpha$, $\gamma$ 1.0
-CLG-VAE $\lambda_{\text{decomp}}$ 40
-$\lambda_{\text{hessian}}$ 40
-forward group 0.2
-group reconst $\{0.2,0.5,0.7\}$
-(b) Hyper-parameters for 3D Shapes
-models hyper-parameters values
-common batch size 256
-epoch 67
-optim Adam
-lr 4e-4
-lr for MIPET 4e-4
-$\beta$-VAE ##\## of IE and EF $\{1,2,4,10\}$
-weight decay 0.0
-latent dim 6
-$\beta$-TCVAE $\beta$$\{4,6\}$
-##\## of IE and EF $\{1,3\}$
-$\alpha$, $\gamma$ 1.0
-weight decay 1e-4
-latent dim 6
-CLG-VAE $\lambda_{\text{decomp}}$ 40
-$\lambda_{\text{hessian}}$ 40
-forward group 0.2
-group reconst $\{0.2,0.5,0.7\}$
-weight decay 0.0
-latent dim 10
-D.3 Impact of Symmetric Matrix Exponential
+D.4 Additional Result of Relation Between Factors and Latent Vector Dimensions
+Figure 6: 3D Car Dataset: Azim. refers to the azimuth factor, and Elev. refers to the elevation factor.
+In the 3D Car dataset, a trade-off was observed.
+While the maximum value across models and evaluation factors showed a slight decrease, the azimuth factor exhibited a notable improvement compared to the declines seen in other factors.
+D.5 Impact of Symmetric Matrix Exponential
 Table 10: The ratio of seeds to show better performance with symmetric matrix
@@
 All results are higher than 0.5, which implies that the constraint enhances I2L equivariance even with uncertain factors.
-D.4 Additional Experiment of Computing Complexity
+D.6 Additional Experiment of Computing Complexity
 Table 11: Training complexity.
-##\## of IE Complexity
+#\# of IE Complexity
 0 $\times$ 1.00
@@
 The results are in Table 11 and represent the training time complexity compare to baselines (when the number of IE is equal to 0).
-D.5 Qualitative Analysis
-[…]
-In Figure 8, $\beta$-TCVAE struggles with y-position and rotation, as shown on the 6th row, and with scale and shape represented on the 7th row.
-On the contrary, MIPET-$\beta$-TCVAE separates y-position and rotation factor (10th, and 7th rows), also the activated dimensions of MIPET-$\beta$-TCVAE are not overlapped with each factor.
-Applied our method on $\beta$-TCVAE shows better disentangled representation on dSprites dataset.
-These results also show that our proposed method improves disentangled representation learning.
-As shown in the Figure 8, $\beta$-VAE struggles with rotation and scale factors in 4thdimension.
-Also, it struggles with x-position and scale factors in 8th dimension, and x-position and rotation factors in 9th dimension.
-However, MIPET-$\beta$-VAE only struggles with rotation and shape factors in 5th dimension.
-As shown in the Figure 8, CLG-VAE struggles with rotation and shape factors in 2nd dimension, and shape and scale factors in 7th dimension.
-However, MIPET-CLG-VAE separates rotation and shape factors in 10th, and 1st dimensions respectively.
-The qualitative analysis with 3D Shapes dataset, as shown in the Figure 11, $\beta$-TCVAE struggles with shape and scale in $5^{th}$ dimension.
-However MIPET-$\beta$-TCVAE As shown in the Figure 11, CLG-VAE struggles with shape and wall color factors in 4th dimension, and shape and object color factors in 7th dimension.
-In particular, it struggles with tree factors in 9th dimension.
-On the other hand, MIPET-CLG-VAE separates shape, wall, and object color factors.
-The qualitative analysis with 3D Cars dataset, as show in Figure 11, the left side is the $\beta$-TCVAE result, and it struggles with body, and azimuth factors shown in the 7th row.
-However, MIPET-$\beta$-TCVAE separates azimuth (6th row) and body (1st row).
-In particular, MIPET-$\beta$-TCVAE learns color factor (3rd row) which does not exist on $\beta$-TCVAE.
-Figure 6: Qualitative results on dSprites.
-The left-side grids are input images and their variants by changing activations of each dimension of latent vectors.
-The first row shows input images.
-The right-side table shows matching pre-defined factors of the dataset (red: MIPET, blue: no MIPET).
-Figure 7: Qualitative analysis result of $\beta$-VAE and MIPET-$\beta$-VAE.
-Figure 8: Qualitative analysis result of CLG-VAE (0.2) and MIPET-CLG-VAE (0.2) with dSprites.
-Figure 9: The Shape is object shape, Orien is an orientation of object, Scale is a scale factor of object, Wall is wall color factor, Floor is floor color, and Object is object color factors.
-It represents the $\beta$-VAE ($\beta=2$) results.
-Figure 10: Qualitative analysis result of CLG-VAE (0.2) and MIPET-CLG-VAE (0.2) with 3D Shapes.
-Figure 11: Qualitative analysis result of $\beta$-VAE (4.0) with 3D Cars.
+D.7 Semantic Mask: Mapping Gaussian Distribution to Diverse Exponential Family
+As shown in Fig.
+7, MIPET-VAE with semantic mask preserves the dataset structure compared to VAE.
+Also, EF conversion with semantic mask maps the Gaussian distribution to the exponential family compared to MIPET without the semantic mask and VAE cases.
+Figure 7: 2D Exponential Family Distribution Datasets